A Complete Bibliography of Publications in Fish Physiology and Biochemistry: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

30 August 2024 Version 1.07

Title word cross-reference

2n − 6 [287]. 3β [206]. 3n − 3 [287, 551]. ^{+} [367, 533, 384, 569, 522, 270, 409, 536]. ^{10} [54]. ^{125} [425, 313]. ^{14} [683, 348, 303, 163, 238, 551]. ^{2+} [412]. ^{3} [179]. ^{35} [220]. ^{6} [54]. _{177} [521]. _{188} [521]. _{2} [410, 390, 484, 50, 170, 550, 269]. _{3} [128, 32, 271, 578, 409, 585]. _{3}^{−} [269]. _{4} [128]. A_{2} [109]. α [190, 287, 541, 384, 567, 375, 544, 569, 208, 502, 77, 78, 573, 572, 509, 396, 492, 613, 500, 356, 372, 521, 536, 355, 179, 510]. β [686, 541, 84, 373, 382, 367, 533, 567, 63, 560, 375, 502, 45, 685, 141, 524, 77, 78, 573, 499, 319, 509, 396, 555, 603, 175, 314, 359, 298, 629, 500, 377, 455, 498, 167, 668, 356, 501, 355, 600, 156, 225, 179, 670, 510]. δ [544]. Δ^{5−4} [206]. γ [641, 362, 544]. μ [578]. n − 3 [553, 265]. ′ [686, 442, 142, 32, 1, 651, 114, 162]. × [370, 342, 353].

-adrenergic [175, 167]. -adrenoceptor [567]. -adrenoceptors [377]. -ATPase [367, 533, 384, 569, 409, 536]. -cells [141]. -dependent [52].

-dihydroprogesterone [355]. -dihydroxy-4-pregnen-3-one [77, 78, 573, 396, 356, 600, 510, 179]. -dihydroxy-4-pregnen-3-one-induced [375]. -dihydroxy-4-pregnen-3-ones [502]. -Estradiol [685, 686, 560, 524, 455]. -ethynylestradiol [541]. -HCH [641]. -hydroxy-20 [355]. -hydroxysteroid [603, 206]. -induced [314]. -iodomelatonin [425]. -isomerase [206]. -K [569]. -like [219, 109]. -linolenic [287, 362]. -methylparatyrosine [372]. -methyltestosterone [521]. -monodeiodinase [442, 142, 1, 114]. -MSH [190]. -naphthoflavone [84, 319, 555, 668]. -naphthoflavone-inducible [63]. -receptor [410]. -reduced [208]. -S [501]. -stimulated [45]. -subunit [500, 536]. -subunits [492]. -sulfate [220]. -tension [50]. -tocopherol [572, 613]. -triiodo-L-thyronine [686, 32, 162].

/H [270]. /K [384, 536].

1- [551]. 11- [208]. 11-ketotestosterone [686, 357, 432, 619, 502, 189]. 123 [216]. 14^{o} [677]. 17 [373, 314]. 17-hydroxy-4-pregnene-3 [600]. 17-hydroxyprogesterone [457, 308]. 17ß [458]. 17ß-estradiol [458]. 1998 [547]. 1A1 [310, 98].

2 [425, 313]. 2- [425, 313]. 2-methylisoborneol [667]. 20-dione [77, 600]. 20-sulphate [396]. 21-dimethyl-19-nor-pregn-4 [77]. 21-oxygenated [208]. 21-trihydroxy-4-pregnen-3-one [573, 501, 510].

3-hydroxy-acyl [358].

4 [480]. 4-nonylphenol [524]. 4’-tetrachlorobiphenyl [480]. 450 [84].

5-hydroxytryptamine [71, 55]. 5-triiodo-l-thyronine [651].

6-bisphosphatase [138].

8^{o} [253].

9-diene-3 [77].

A. [606, 607]. A1 [441]. abduction [422]. ability [651]. abnormal [565]. Absence [6, 579, 283]. absorption [595, 526, 113, 162]. abundance [274]. acanthias [209]. Acanthopagrus [373, 382, 685, 510]. accelerated [625]. acceleration [305]. access [214]. acclimated [402, 436, 254]. acclimation [259, 51, 129, 318, 320, 195, 277, 333, 677, 132]. acclimatization [469, 487]. accommodations [193]. accompanied [49, 274]. accompanying [144]. Accumulation [459, 640, 103, 266, 123, 132]. acetate [161]. acetate-induced [161]. acetylcholine [370, 73]. acetylcholinesterase [160, 486]. acid [139, 287, 362, 340, 636, 410, 562, 236, 106, 128, 376, 341, 381, 681, 577, 86, 639, 473, 142, 462, 485, 558, 271, 252, 119, 14, 154, 284, 453, 107, 248, 155, 583, 450, 428, 11, 12, 466, 30, 80, 105, 613, 576, 359, 664, 553, 612, 667, 126, 354, 161, 35, 339, 675, 265, 65]. acid-base [236, 376, 639, 252, 453, 576, 354, 65]. acid-deficient [636]. acidic [416, 141, 335]. acids [127, 287, 610, 348, 392, 622, 250, 163, 238, 405, 394, 30, 105, 613, 623, 288, 579, 344, 28, 29, 35, 282, 85, 383, 265, 395]. Acipenser [470, 357, 256, 360, 428, 443, 88]. acquisition [7]. action [38, 529, 476, 195, 183, 194]. Actions [72, 92, 194, 525, 575, 523]. activate [282]. activated [543, 69, 684]. active [440]. Activin [505]. activities [471, 118, 251, 358, 587, 147, 548, 559, 319, 586, 363, 254, 589]. Activity [634, 322, 236, 442, 84, 160, 262, 549, 142, 138, 448, 203, 49, 273, 33, 1, 627, 385, 107, 152, 141, 682, 360, 572, 428, 393, 503, 114, 192, 306, 372, 609, 395]. aculeatus [477, 318, 432, 619]. acute [291, 230, 376, 26, 351, 123, 564]. acyl [358]. adaptation [312, 342, 389, 226]. Adaptations [193]. adapted [331, 309, 563, 278, 612]. adaptive [153]. adduct [83]. adenosine [69, 441]. administered [546]. Administration [86, 281, 508, 54, 409, 156]. adrenaline [279, 582]. adrenergic [280, 175, 167, 269]. adrenoceptor [567]. adrenoceptors [377]. Adriatic [360]. adult [45]. adults [668]. Advanced [547]. aerobic [234]. aestivate [17]. affect [567, 289]. affected [428]. affects [308, 612]. affinis [20]. affinity [421, 169, 313, 274]. African [506, 451, 465, 577, 639, 493, 679, 568, 500, 189, 407, 363, 467, 227, 204]. after [76, 155, 31, 644, 565, 660]. Age [425, 318, 154, 418, 363]. Age-related [425]. agents [391, 280]. agglutinin [581]. Ah [542]. air [404, 680]. air-breathing [404, 680]. Al [132]. Ala [54, 216]. Albula [251, 419]. albumin [579]. Alburnus [398]. alevins [298]. alimentary [118]. alkaline [103, 273]. All-fish [215]. Allozyme [588]. Alosa [627]. alpha [300, 492]. alpha-tocopherol [300]. alpinus [253, 686, 94, 557, 80, 105, 613, 652, 226, 265]. alterations [306]. altered [259]. altering [617]. alternative [572]. alters [669]. aluminium [335, 132]. aluminum [32]. amago [518]. ambient [32, 125]. American [627]. americana [374]. americanus [158, 534, 257, 446]. Amia [17]. Amino [577, 341, 119, 392, 14, 155, 450, 466, 394, 344, 395]. aminotransferase [634]. ammocoetes [19]. Ammonia [554, 338, 103, 337, 170, 611, 467]. ammonia-exposed [337]. ammonium [680]. among [318]. AMP [317]. AMP-deaminase [317]. Amylase [406, 586, 20, 66]. Anabas [315]. anadromous [557, 627]. anaemia [390]. anaesthesia [337]. analog [74, 158]. analogue [608]. analyses [656]. analysis [244, 176, 563, 668]. Anatomic [51]. Androgen [511, 207, 432]. Androgen-induced [511]. androgens [506, 449, 79, 226]. angiotensin [388, 133, 528]. Angiotensins [145]. Anguilla [433, 537, 456, 367, 533, 140, 188, 34, 168, 447, 606, 607, 463, 71, 125, 509, 507, 441, 508, 18]. anguillicaudatus [413]. anhydrase [262, 358]. animal [568, 56]. anion [639, 69]. Announcement [102, 678]. Annual [112, 161, 147, 323]. Anoplopoma [54]. anserine [304]. Antarctic [8, 351, 272, 582, 334]. Antarctica [571]. antibiotics [546]. Antidiuretic [529]. antifreeze [549, 22]. antioxidant [418, 480]. antioxidantive [681]. antiport [474]. antiporters [270]. Antisomatostatin [305]. Antisomatostatin-induced [305]. antithyroid [113]. aorta [268]. Apomorphine [568]. Apomorphine-induced [568]. appearance [683]. Application [547, 508]. April [173]. Aquaculture [547, 546]. aquatic [404]. Arachidonic [139, 340]. arches [220]. Arctic [253, 686, 94, 40, 557, 457, 80, 105, 613, 652, 226, 265]. Arg [152]. Arginine [198, 663]. Argulus [616]. Aristichthys [589]. aromatase [633, 329, 477]. aromatizable [79]. arterial [244]. articles [21, 171, 276, 325, 378, 438]. Artificial [507]. Ascorbate [47, 46, 149]. ascorbate-2-monophosphate [149]. Ascorbate-2-sulfate [47, 46]. Ascorbic [428, 381, 86]. aspartate [476, 634]. aspects [193]. assay [645, 219]. assays [587]. assimilation [246]. associated [549, 75]. association [661]. astaxanthin [683]. Atlantic [683, 329, 322, 588, 255, 294, 384, 539, 584, 424, 429, 386, 279, 247, 388, 98, 266, 544, 273, 337, 552, 414, 147, 131, 638, 117, 250, 300, 155, 47, 524, 239, 644, 628, 548, 295, 583, 328, 60, 620, 664, 221, 598, 672, 46, 149, 637, 435, 59, 110, 554, 275, 35, 321, 339, 245, 580, 231, 512, 677, 343, 520]. ATP [647, 669]. ATPase [471, 367, 533, 384, 569, 409, 536, 254]. ATPases [125]. atrial [92, 36, 279, 370]. aurata [599, 658, 596, 425, 517, 246, 154, 617, 11, 12, 163, 238, 551, 393, 296, 553]. auratus [180, 349, 445, 320, 434, 659, 208, 220, 282, 401, 85, 186, 670, 514, 397]. aureus [370, 342, 353, 297, 150]. Australian [461]. australis [49, 19, 45, 606, 607]. authors [61, 174, 327, 380]. autumn [107]. availability [548, 371]. avoidance [334]. AVT [531]. axis [241].

B [593, 615]. bacteria [615]. bacterial [376]. baeri [357, 428, 88]. Baikal [594]. balance [468, 376, 639, 309, 439, 131, 576]. barbel [130]. Barbus [130]. Basal [410, 562, 98]. base [236, 376, 639, 271, 252, 453, 576, 354, 65]. based [587]. basic [416]. bass [68, 346, 368, 403, 448, 462, 399, 75, 387, 264, 374, 364, 277, 458, 56, 406, 483, 586, 228, 254]. batrachus [116, 375, 408, 680, 634]. before [273, 606, 607, 660]. behavior [511]. behavioral [591]. behaviour [133, 52]. benthic [594]. benthic-pelagic [594]. benzo [83, 286, 479]. bernacchii [351, 272]. Bester [649]. beta [83, 286, 89]. between [378, 438, 506, 260, 384, 365, 595, 445, 50, 42, 644, 219, 135, 576, 151, 126, 528]. bifunctional [39]. big [589]. big-head [589]. bile [543]. biliary [123]. billabongs [461]. Binding [222, 501, 6, 425, 99, 182, 432, 622, 219, 313, 296, 221, 274, 624, 579, 311, 581, 179, 397]. bioactivity [46, 222]. Biochemical [547, 596, 441, 556, 660, 260, 381, 646, 224, 617]. Biochemistry [197, 446]. Biological [92, 216, 503, 366]. biologically [440]. biology [489, 603]. biosynthesis [357]. Biotransformation [319]. bisphosphatase [138, 448]. bitterling [316]. black [373, 382, 235, 685, 418, 510]. bladder [311]. Bleeker [685]. blends [584]. blennies [590]. Bloch [160, 33, 293, 478, 161, 306]. blockers [58]. blocks [152]. Blood [683, 376, 235, 461, 244, 582, 10, 94, 421, 50, 281, 219, 278, 377, 167, 82, 269, 65]. blue [297]. bluegills [659]. boddaerti [69, 70, 578]. bodies [435]. body [332, 674, 229, 597, 135, 126, 85, 132]. Bohr [421]. Boleophthalmus [69, 70, 578]. bonariensis [661]. bonefish [251, 419]. bonnethead [591]. borage [362]. borchgrevinki [8, 582]. borne [104]. both [120]. bound [317]. brain [641, 201, 310, 454, 160, 188, 425, 610, 119, 187, 494, 385, 551, 496, 441, 55, 112, 23, 192, 435, 564]. brain-pituitary [494]. brains [659]. Branchial [538, 199, 582, 69, 447, 436]. Brandt [470, 357, 428]. bream [596, 517, 484, 543, 426, 11, 12, 163, 551, 550, 296, 109, 156]. breams [425]. breathing [404, 680]. brevirostris [24]. broodstock [11, 12, 553]. broodstocks [94]. brook [291, 145, 141, 77, 78, 3, 371, 422]. brown [459, 511, 195, 643, 324, 466, 444]. buffering [304]. building [619]. bullhead [418]. Burchell [639, 467].

C [683, 253, 136, 365, 348, 181, 303, 47, 163, 238, 551, 621, 46, 149, 231, 282, 677]. C-astaxanthin [683]. C-labelled [348, 303, 163, 238]. Ca [412, 149]. cadmium [481, 539, 475]. cadmium-pretreated [475]. caeca [484, 59, 110]. calcitropic [272]. Calcium [309, 530, 58, 481, 538, 399, 420, 475, 151, 314, 657, 199]. calculations [96]. calva [17]. cAMP [167]. Can [26, 578]. canal [302, 118]. canaliculatus [347]. canicula [82]. cannulation [268]. canola [584]. capacities [318, 628]. capacity [304, 234, 662, 672]. captive [144]. Carassius [471, 180, 349, 320, 434, 659, 208, 220, 282, 85, 186, 670, 514, 397]. carbamate [512]. Carbohydrate [247, 164, 107, 406, 435, 564, 257]. carbohydrates [342]. carbon [244]. carbonic [262, 358]. Carcharhinus [235]. Cardiac [236, 95]. Cardiovascular [351, 666, 528]. carotenoid [664]. carotenoids [686]. carp [471, 658, 538, 242, 81, 153, 216, 203, 421, 50, 308, 439, 185, 389, 107, 676, 555, 222, 579, 565, 640, 675, 15, 13, 292, 589, 545]. carpio [658, 566, 538, 81, 153, 216, 50, 308, 555, 222, 579, 565, 675, 15, 13, 292]. carried [27]. cartilage [350]. castration [161]. catabolism [402]. Catalytic [84, 147]. Catecholamine [192, 618]. catecholamines [351, 237, 175, 528, 65]. Catfish [657, 136, 506, 451, 465, 577, 639, 493, 375, 408, 679, 157, 568, 655, 268, 293, 500, 680, 667, 189, 407, 363, 161, 306, 372, 609, 467, 227, 204]. Cathepsin [654]. cathepsins [593, 615]. cation [639, 69]. cation-anion [639]. caught [238]. caused [572]. CCK [656]. Cd [285]. cDNA [658, 176, 644, 206, 656, 670, 514]. Cell [210, 669, 562, 456, 8, 283, 50, 151, 377, 407, 675, 556, 292]. Cell-cell [210]. cells [139, 127, 537, 27, 539, 485, 558, 165, 181, 286, 312, 436, 563, 141, 499, 202, 290, 519, 167, 602, 648, 530, 28, 29, 35, 321, 339, 409, 413, 581, 412, 661, 575, 269, 43, 204]. Cellular [632]. centrifugation [93]. cephalus [146]. chain [673, 30]. Chalcalburnus [103]. change [230, 264, 234]. Changes [201, 68, 259, 473, 75, 273, 336, 19, 155, 518, 405, 509, 540, 245, 585, 291, 153, 425, 142, 108, 361, 203, 69, 663, 45, 141, 323, 144, 628, 548, 169, 170, 561, 423, 80, 652, 296, 274, 363, 161, 372, 609, 210, 57, 536, 446, 114]. Channa [33]. channel [136, 58, 655, 268, 667]. channelblockers [538]. Chanos [106, 121, 144]. char [613, 652, 226]. characteristics [669, 381, 202, 23, 501]. Characterization [599, 382, 403, 486, 655, 427, 393, 240, 444, 478, 637, 110, 413, 397, 566, 138, 601, 116, 408, 100, 645, 484, 543, 243, 97, 53, 263, 328, 550, 441, 228, 13, 556, 660]. charr [253, 686, 94, 557, 457, 80, 105, 265]. chemistry [376]. Chinook [647, 305, 148, 290]. chloride [539, 399, 151, 680, 409]. chloromethyl [152]. chlorophenylalanine [372]. cholecystokinin [656]. cholesterol [303]. cholinergic [447]. chondroitin [240]. choriogenins [504]. Chromatographic [258]. chromatography [23, 227]. Chromis [445]. chronic [641, 142, 548, 123]. Chrysophrys [156]. chrysops [374]. chrysospilos [26]. chum [127, 336, 299]. chymotrypsin [429]. cichlid [440]. Circadian [128]. circasemilunar [143]. Circulating [573, 506, 296, 65]. Circulatory [36, 351]. Clarias [506, 451, 465, 577, 639, 116, 375, 408, 568, 500, 680, 189, 407, 363, 634, 467, 227, 204]. Clariidae [679]. clarki [423, 423]. class [287, 473, 557, 464, 284, 594, 80]. classes [595, 11, 405, 28, 29, 339]. clearance [298, 631]. clearnose [350]. clofibric [667]. Cloning [176, 658, 498, 670, 514]. Close [661]. closely [590]. Clupea [345]. clupeid [627]. CO [390, 170, 269]. CoA [39, 358]. coastal [423]. coated [545]. cobalt [538]. cod [410, 164, 285, 249, 638, 155, 628, 522, 59, 110]. Coho [633, 218, 604, 241, 169, 217, 79, 298, 196, 479]. cohort [291]. cold [332, 278, 612, 254]. cold-acclimated [254]. cold-adapted [278, 612]. combination [330]. combined [236]. common [599, 658, 81, 185, 427, 579, 640]. communicate [602, 648]. communication [54]. Comparative [559, 73, 313]. Comparison [384, 590, 162, 322, 216, 126]. compensation [143]. complement [572]. complex [494]. component [433]. composition [362, 214, 332, 674, 473, 584, 448, 462, 485, 557, 464, 284, 594, 398, 453, 250, 583, 11, 12, 80, 613, 664, 553, 612, 126, 321, 254, 265]. compositions [287, 340, 636, 339]. compound [666]. compounds [272, 55, 112]. computerized [7]. concentrating [176]. concentration [266, 49, 50, 308, 428, 522, 289]. concentrations [322, 128, 341, 365, 142, 399, 19, 685, 155, 125, 371, 281, 578, 571, 680, 299]. conditions [344]. confinement [616]. conjugates [502]. consequence [576]. constant [320]. constituents [260]. Constitutive [589]. constructs [215]. consumption [94, 360, 472, 423, 65]. contaminated [101]. content [128, 142, 531, 50, 334, 241, 405, 522, 289, 161, 306]. Contents [62, 173, 326, 379, 587, 295]. contractility [259, 230, 72, 460, 677]. contraction [145]. contractions [58]. Control [580, 184, 70, 618, 191, 41, 490, 199]. controlled [7]. conversion [638, 30, 598, 156, 269]. converting [388]. corn [635]. coronary [301]. correlated [625]. correlating [372]. Correlations [614]. Corrigendum [648]. cortices [179]. corticosteroids [208, 280, 123]. Cortisol [291, 294, 567, 5, 423, 641, 128, 457, 31, 651, 431, 169, 450, 281, 91, 605, 175, 274, 662, 616, 257]. cortisol-binding [274]. Cortisol-induced [423]. costs [465]. Cottocomephorus [594]. Cottoidei [594]. Cottus [540]. coupling [602, 648]. creosote [101]. creosote-contaminated [101]. croaker [520]. cross [63]. cross-reactivity [63]. crowding [617]. crucian [471, 107]. Ctenopharyngodon [389]. culture [485, 558, 297, 602, 648, 321, 339, 675, 227]. cultured [433, 127, 279, 122, 563, 370, 88, 664, 28, 29, 35]. cultures [222]. current [511]. currents [530]. cutthroat [423]. cycle [291, 625, 108, 448, 203, 147, 323, 458, 22, 634]. cyclic [161]. cyclostomes [44]. Cynoscion [501]. CYP1A1 [668]. Cyprinformes [413]. Cyprinidae [103, 146, 398]. Cyprinus [658, 566, 538, 81, 153, 216, 50, 308, 555, 222, 579, 565, 675, 15, 13, 292]. cyproterone [161]. Cys [216]. Cys-123 [216]. cysteine [395]. cytochrome [310, 84, 415, 98, 63, 147, 319]. cytokeratin [413]. cytoplasmic [389]. cytoskeleton [539]. cytosol [77, 444]. cytosolic [48, 274].

D [54, 272, 256, 476]. D-Ala [54]. D-glucose [256]. dab [427]. daily [37, 573]. daily-spawning [37]. Danio [542]. data [7]. day [306]. day-night [306]. daylight [489]. days [657]. deaminase [317]. decarboxylase [322]. Decreased [266, 124]. decreases [333]. defense [480]. deficiencies [572]. deficiency [553, 675]. deficient [636]. degradation [251]. dehydrogenase [358, 283, 603, 206]. dehydrogenase/ [206]. dehydrogenases [389]. deiodinase [211, 559, 585]. deiodinases [630]. deionized [293]. delivery [217]. demand [128, 214]. demand-feeders [128, 214]. demonstrated [602, 648]. Demonstration [77]. density [433, 111, 165, 93, 617, 579]. dentex [599, 599, 653]. deoxyribonuclease [59]. depend [406]. Dependence [229, 411, 468]. dependent [410, 535, 79, 564, 52]. Depletion [572]. depots [557]. deprivation [597, 435]. desaturation [348, 105]. Detection [542, 23]. Determination [32, 587, 646]. determined [620]. determines [614]. determining [318]. developing [481, 679, 405, 542]. Development [40, 264, 219, 508, 68, 588, 451, 373, 456, 473, 666, 130, 241, 5, 458, 682, 431, 353, 499, 393, 324, 509, 298, 113, 114, 407, 565, 57, 159, 581, 15, 292]. Developmental [561, 459, 654, 406, 600]. devriesi [334]. Dexamethasone [454]. diadromous [57]. diandric [478]. Dicentrarchus [541, 368, 403, 448, 462, 75, 264, 364, 458, 406, 483, 586, 254]. dichlorvos [558]. dieffenbachii [606, 607]. Diel [663, 322, 291]. diene [77]. diet [411, 468, 140, 4, 448, 264, 360, 555, 126, 640, 66]. Dietary [136, 164, 392, 213, 583, 483, 586, 231, 362, 340, 346, 381, 332, 674, 81, 639, 584, 36, 462, 610, 595, 266, 246, 154, 117, 250, 342, 47, 148, 472, 217, 572, 261, 428, 394, 105, 613, 406, 635, 288, 46, 149, 104, 521, 150, 344, 275, 265, 545, 80]. dietary-protein [545]. diets [287, 636, 673, 610, 11, 12]. Differences [464, 562, 596, 639, 94, 79]. different [340, 669, 128, 368, 610, 63, 345, 285, 284, 303, 342, 148, 657, 590, 598, 672, 680, 624, 490, 87, 600, 85]. Differential [188, 33]. Differentiation [456]. digesta [284, 303]. digestibility [652]. digestion [284, 303, 156]. Digestive [118, 682, 599, 368, 361, 264, 246, 284, 303, 638, 393]. dihydroprogesterone [355]. dihydroxy [375, 502, 77, 78, 573, 396, 356, 600, 179, 510]. dimethyl [77]. dinitrate [666]. dione [77, 600]. dioxide [244]. dioxygenase [395]. diploid [291, 217]. diploids [297]. Direct [575]. directions [511]. discontinuous [93]. disease [376, 8, 231]. dismutase [49]. dispersed [181]. dispilus [445]. distance [627]. Distinct [492, 496, 202]. Distribution [310, 55, 401, 411, 686, 118, 454, 513, 218, 188, 544, 249, 241, 317, 632, 16, 240, 20]. disturbances [439]. Diurnal [37, 214, 289, 330]. DNA [209]. DNase [110]. Do [638, 396, 282]. docosahexaenoic [340]. Does [17, 234, 289, 86, 567]. dogfish [210, 82]. domesticated [112]. dominance [489]. domperidone [478]. donor [388]. Dopamine [186, 385, 23, 575]. Dopaminergic [443]. dorsal [268]. Dosage [79]. Dosage-dependent [79]. Dose [564, 620]. Dose-dependent [564]. down [629]. down-regulate [629]. Drinking [388, 133, 166, 561]. drug [113]. drugs [582]. drum [332, 674, 673, 248, 213, 261, 134, 520]. due [304]. duration [24]. during [633, 253, 2, 68, 291, 562, 259, 456, 235, 473, 142, 108, 130, 361, 49, 273, 336, 229, 146, 42, 147, 312, 19, 300, 643, 323, 239, 682, 295, 561, 12, 653, 393, 509, 396, 657, 209, 298, 123, 196, 114, 407, 479, 634, 18, 57, 299, 536, 245, 132, 225, 204, 585]. dye [602, 648]. dye-coupling [602, 648]. Dynamic [23]. Dynamics [643, 344, 560, 557, 554, 25].

early [459, 541, 107, 561, 324, 298]. Eastern [103]. eating [440]. ectoparasitic [616]. eel [433, 482, 592, 593, 615, 537, 456, 367, 533, 188, 34, 168, 447, 526, 125, 509, 507, 519, 449, 441, 270, 508, 18, 197]. eels [140, 606, 607, 463]. Effect [214, 346, 341, 538, 140, 539, 666, 386, 462, 387, 420, 119, 457, 131, 11, 12, 555, 635, 624, 667, 226, 600, 139, 368, 260, 262, 429, 272, 32, 485, 246, 154, 19, 627, 281, 280, 267, 418, 79, 455, 474, 104, 275, 556, 65, 64]. Effects [633, 641, 329, 287, 362, 340, 636, 625, 477, 58, 391, 481, 381, 469, 332, 674, 9, 95, 279, 582, 10, 256, 558, 421, 626, 320, 249, 597, 337, 117, 71, 524, 431, 125, 268, 170, 360, 630, 261, 466, 574, 522, 662, 288, 598, 354, 306, 372, 521, 205, 321, 339, 307, 460, 412, 257, 677, 265, 66, 136, 541, 230, 577, 218, 604, 36, 247, 390, 285, 83, 286, 33, 185, 370, 463, 213, 458, 353, 611, 96, 519, 480, 605, 56, 175, 359, 377, 435, 564, 512, 338, 397, 47, 113]. efficiency [255, 638, 598, 422]. egg [625, 646, 504, 553, 57]. eggs [473, 349, 654, 11, 405, 664, 298, 57]. eglanteria [350]. eicosanoid [287]. eicosapentaenoic [362]. Eilat [547]. eitherwith [545]. Elasmobranch [302, 618, 603]. elasmobranchs [528]. electric [511]. electric-social [511]. electrochemical [23]. electrolytes [139]. Electromechanical [370]. electron [204]. electroreceptor [657]. elements [655, 47]. elevate [442]. Elevated [571, 611]. elicited [424]. elimination [667]. ELISA [645]. elongation [348, 105]. elopomorph [240]. embryonic [241]. embryos [309, 416, 457, 665, 548, 631, 668, 565, 467, 542, 383]. emphasis [404]. emulsion [508]. enalapril [388]. encoding [176, 504, 206, 670, 514]. endemic [103]. Endocrine [323, 565, 515, 224, 512, 43]. endocrinology [178]. endocytic [288]. Endocytosis [27, 76]. Endogenous [418, 358, 152]. endolymph [576]. endoplasmic [412]. endothelins [529]. endothelium [6]. Energetics [419, 345]. energy [597, 369, 146, 150]. enhance [395]. Enhanced [659, 31]. enhancement [217]. enkephalin [74, 71]. enoyl [39]. enoyl-CoA [39]. enriched [362]. enterocytes [652, 623]. envelope [504]. Environment [547, 141, 101]. Environmental [490, 592, 166, 26, 332, 420, 252, 453, 19, 548, 578]. environments [436, 657]. Enzyme [549, 39, 566, 118, 388, 645, 627, 342, 682, 548, 319, 393, 363, 52]. enzyme-linked [645]. enzymes [658, 346, 95, 264, 246, 638, 107, 364, 317]. enzymology [603]. Epidermal [482, 592, 593, 615, 413]. Epinephrine [442, 72]. epithelial [312, 602, 648]. epithelium [537, 436]. epoxide [427]. equilibria [94]. equivocal [605]. ergic [191]. Errata [233]. Erratum [172, 400]. Erythrinidae [229]. Erythrocyte [93, 153, 67]. erythrocytes [614, 3, 270, 474]. erythrophthalmus [115]. Esox [473, 348, 126]. Essential [248, 636, 106, 392, 675]. ester [303]. esterification [623]. esterified [128]. Estradiol [685, 686, 541, 382, 560, 243, 45, 463, 524, 458, 88, 359, 298, 455, 156, 225, 373, 314]. estradiol-17 [541, 382, 45, 359, 298, 156, 225, 373, 314]. Estrogen [494, 14, 503]. ethanol [667]. ethylamide [54]. ethynylestradiol [541]. European [367, 533, 140, 188, 462, 168, 447, 157, 463, 519, 449]. euryhaline [5, 651, 561]. evaluation [364, 575]. evidence [605]. evoke [86]. evolution [527]. evolutionary [224]. excretion [411, 103, 390, 467, 338]. exercise [244, 95, 317, 450]. exercise-training [95]. exhaustive [244, 450]. exocytosis [530]. exogenous [246, 480, 298]. experimental [188, 390, 630, 267]. experimentally [376]. experimentally-induced [376]. experiments [7]. exposed [236, 477, 252, 337, 453, 141, 480, 335]. exposure [641, 647, 541, 26, 657, 565]. expressing [661]. Expression [533, 569, 416, 552, 459, 367, 384, 218, 130, 659, 644, 504, 496, 503, 632, 498, 598, 206, 634, 656, 184, 66, 670]. extant [118]. external [399, 42]. Extracellular [96, 334, 354]. extract [243, 292]. extragonadal [42]. Extrapituitary [182]. extremely [103]. eye [610]. eyes [551].

factor [92, 218, 375, 408, 421, 416, 552, 643, 651, 518, 574, 219, 296, 632, 311, 186]. factor-I [552, 643, 518, 574, 296]. factors [42]. facultative [404]. Fadrozole [633]. families [118]. family [218, 197]. farmed [596, 414, 250]. Fast [465, 72, 556]. fasted [124, 162]. fasting [142, 75, 164, 119, 120, 466, 585]. fasting/refeeding [585]. Fatty [664, 612, 127, 287, 340, 636, 106, 128, 473, 142, 462, 610, 485, 558, 348, 154, 284, 622, 250, 107, 248, 583, 11, 12, 163, 238, 405, 30, 80, 105, 613, 623, 553, 288, 126, 579, 28, 29, 35, 339, 675, 282, 85, 383, 265]. fed [128, 36, 124, 610, 266, 652]. Feed [253, 364, 598, 90]. feedback [189]. feeders [128, 214]. feeding [341, 429, 142, 256, 164, 345, 119, 155, 120, 466, 289, 330, 616, 90]. female [633, 625, 365, 645, 239, 316, 297, 306, 422]. females [357, 606]. ferritin [459]. ferritins [6]. ferrugineus [608]. fertility [260]. fertilization [507]. fibers [661]. fibres [72]. fibrinogen [684]. fibroblast [416, 311]. fibroblasts [122, 91]. fimbria [54]. final [374, 396, 574]. fingerling [136]. Fish [650, 527, 198, 362, 566, 215, 415, 260, 681, 242, 81, 86, 351, 160, 9, 516, 193, 582, 7, 491, 98, 63, 434, 417, 416, 187, 502, 644, 618, 504, 178, 418, 480, 234, 296, 404, 175, 532, 478, 474, 612, 634, 642, 490, 311, 35, 675, 461]. fishes [445, 594, 559, 240, 571]. five [486, 559, 258, 657]. flatfish [92, 122]. flesus [531]. flooding [620]. Flounder [212, 608, 531, 158, 22, 534, 656, 525]. fluctuations [371]. fluid [294, 53, 96, 245]. Fluorescence [593, 587]. fluorescence-based [587]. fluviatilis [654, 665]. flux [255, 4, 399]. foliaceus [616]. follicle [625]. follicles [649, 349, 560, 145, 78, 605, 662, 671, 356, 205, 25, 307]. following [647, 244, 124, 217, 450, 54]. fontinalis [291, 145, 141, 77, 78, 3, 371, 422]. Food [435, 639, 597, 638, 141, 371, 612, 156]. forebrain [497]. form [346]. formalin [558]. formation [350, 83]. forms [188, 486, 493, 243]. Forsskål [106, 121, 144]. fossilis [160, 293, 161, 306, 372, 609]. found [433]. fractional [545]. fractionation [93]. fractions [544, 317, 444]. fragments [649]. francisci [302]. free [341, 604, 142, 317, 155, 394, 344, 545]. Freerunning [143]. freezing [334]. fresh [315, 125, 561, 267, 133, 634, 580]. freshwater [253, 331, 562, 81, 486, 168, 309, 312, 563, 606, 607, 323, 570, 618, 293, 151, 612, 73, 57, 675, 461]. freshwater-adapted [309]. Fructose [448, 138]. fructose-1 [138]. fry [388, 298, 149]. function [302, 331, 166, 674, 352, 242, 256, 229, 154, 45, 512, 461]. Functional [537, 599, 669, 541, 513]. functioning [657]. functions [368, 224, 186]. Fundulus [143, 651, 574, 101]. Further [605, 575]. future [511].

GABA [191]. GABA-ergic [191]. gabaergic [641]. Gadus [410, 562, 164, 249, 638, 71, 628, 59, 110]. Gag [645]. gairdneri [2, 27, 36, 32, 5, 31, 28, 25, 65, 64]. gairneri [4]. Gambusia [20]. gap [602, 648]. gariepinus [506, 451, 465, 577, 639, 568, 500, 189, 407, 363, 467, 227, 204]. gas [227]. Gasterosteus [477, 318, 432, 619]. gastrointestinal [552, 364]. gelatin [508]. gene [215, 242, 384, 516, 503, 498, 642, 184, 670]. general [519]. genes [504, 496]. genetically [672]. genotypes [624]. Geotria [49, 19, 45]. germ [581, 292]. get [440]. GH [519, 200, 499, 219, 296, 521]. GH-binding [296]. ghost [511]. gill [139, 6, 376, 537, 384, 8, 539, 271, 286, 563, 263, 220, 125, 611, 151, 532, 602, 648, 409, 536, 412, 132, 312, 404, 134]. gills [535, 569, 252, 502, 152, 581, 254]. gilthead [599, 658, 596, 517, 246, 426, 572, 617, 11, 12, 163, 551, 393, 296, 553]. gland [241, 500]. gland-thyroid [241]. glands [38]. glanis [157]. glucagon [34, 75, 224, 90, 87]. glucagon-like [224, 90, 87]. glucocorticoid [169]. glucokinase [658]. glucokinase-like [658]. Gluconeogenesis [138, 283]. Glucose [137, 291, 128, 256, 369, 148, 289, 90]. Glucose-stimulated [137]. glue [619]. glutamine [401]. Glutathione [101, 2, 48, 3, 555, 480]. glutathione-S-transferases [48]. glutathione-treated [3]. gluten [635]. glutinosa [247]. Gly [54]. glycerol [549]. glycerophospholipid [339]. glycogen [128, 34, 295, 650, 289]. glycogenolysis [34]. glycoprotein [509, 492, 500]. glycosaminoglycan [419]. Glycosidase [251]. GnRH [201, 495, 188, 182, 158, 183, 496, 497, 478]. GnRH-A [158, 478]. golden [238]. goldfish [180, 513, 349, 505, 182, 320, 434, 181, 659, 208, 488, 622, 220, 282, 85, 383, 186, 575, 670, 514, 397]. Gonad [606, 607]. gonadal [495, 68, 445, 273, 374, 42, 323, 458, 353, 91, 292, 204]. gonadotropes [202]. gonadotropic [158, 407]. Gonadotropin [608, 493, 488, 38, 180, 188, 182, 181, 494, 685, 509, 258, 443, 500, 372, 205, 661, 600, 575]. gonadotropin-II [575]. gonadotropin-induced [205]. Gonadotropin-releasing [493, 488, 188, 182, 258, 661, 575]. Gonadotropins [396, 625, 159]. gonads [541, 470, 448, 396, 299]. gonochoristic [541]. graded [229]. gradient [93]. grandis [143]. grass [389]. grey [238]. grouper [645]. growing [297, 556]. Growth [200, 217, 18, 322, 253, 287, 362, 636, 489, 625, 128, 255, 215, 74, 180, 465, 332, 352, 242, 140, 639, 515, 584, 218, 604, 124, 142, 516, 386, 216, 462, 318, 476, 345, 387, 416, 181, 246, 552, 195, 638, 568, 117, 185, 31, 643, 651, 431, 305, 353, 518, 178, 574, 219, 522, 519, 576, 234, 296, 632, 635, 113, 598, 672, 194, 354, 521, 150, 642, 311, 321, 307, 556, 661, 186, 575, 156, 523, 660, 397, 46]. GTH [201, 625, 183, 202, 498, 189, 204, 38, 499]. GTH-cells [204]. GTH-II [189]. gueldenstaedti [470]. gulf [83, 286]. gulonolactone [86, 428]. Gut [44, 166, 587].

H [270, 179, 410, 474]. habitats [590]. haematology [46, 265]. haemoglobin [50, 237]. hagfish [247]. halibut [328, 554]. handling [351, 124]. hangiongensis [540]. harengus [345]. hatchery [94]. hatching [201, 232, 668, 554]. HCH [641]. HCl [271]. HCO [269]. head [589]. heart [471, 235, 10, 333]. hearts [313]. heat [392, 637]. held [128]. hematocrit [291]. Hematological [434, 417, 67]. hemisphere [45, 19]. Hemitripterus [257]. Hemoglobin [422, 94, 266, 320, 624, 461]. heparin [311]. heparin-binding [311]. Hepatic [1, 342, 546, 84, 223, 352, 567, 142, 247, 601, 544, 100, 49, 147, 427, 555, 296, 449, 289, 101, 257, 395, 585]. hepatocyte [650]. hepatocytes [2, 34, 283, 89, 221, 288, 222]. hepatopancreas [543, 550]. hepatosomatic [128]. herring [345, 52]. Heterobranchus [679]. heteroclitus [651, 574, 101]. Heterodontus [302]. Heteropneustes [160, 293, 161, 306, 372, 609]. heterozygosity [588]. hexokinase [658]. hexokinase- [658]. High [617, 433, 36, 256, 304, 579]. highly [553]. Hippoglossus [328, 554]. histamine [562]. histamine-stimulated [562]. histochemical [361]. Histological [361, 141]. histology [606, 607]. histopathology [362, 636]. holbrooki [20]. homeostasis [580]. homogenate [139, 509]. homologous [159, 292]. Hoplias [229]. Hormonal [242, 220, 144, 297, 89, 534, 312, 70, 650, 632]. hormone [315, 489, 128, 214, 215, 180, 608, 352, 242, 218, 188, 211, 604, 124, 142, 516, 386, 216, 200, 176, 182, 158, 476, 181, 488, 19, 241, 195, 568, 185, 31, 651, 37, 353, 509, 219, 522, 258, 56, 296, 631, 500, 289, 330, 194, 508, 407, 354, 521, 205, 642, 18, 299, 227, 661, 186, 575, 184, 523, 660, 397, 519]. hormone- [519]. hormone-analogue [608]. hormone-release [568]. hormone-releasing [158, 186]. hormone/prolactin [218]. hormones [527, 68, 537, 247, 493, 626, 213, 573, 261, 324, 492, 44, 629, 113, 114, 57, 257, 670, 212]. horn [302]. HUFA [572]. human [242]. humoral [46]. hybrid [649, 353]. hydratase [39]. hydrolase [427]. hydroperoxide [2]. hydroxy [358, 355, 600]. hydroxyapatite [350]. hydroxylase [385, 192]. hydroxyprogesterone [457, 308]. hydroxysteroid [603, 206]. hydroxytryptamine [71, 55]. hypercapnia [236, 252, 453]. hyperoxia [271]. hyperthyroidism [630]. hypo [630]. hypo- [630]. Hypoosmoregulation [166]. hypoosmoregulatory [5, 31, 651]. hypophyseal [609]. hypophysectomized [242]. hypophysectomy [397]. hypophysis [40]. Hypostomus [404]. hypothalamic [33, 306, 372, 609]. hypothalamo [609]. hypothalamo-hypophyseal-ovarian [609]. hypothalamus [176, 385]. hypothyroidism [604]. hypotonic [436]. hypoxia [236, 10, 229, 591, 377]. hypoxic [26].

I- [202]. Ictalurus [136, 655, 268, 667]. idella [389]. Identification [408, 311, 39, 258, 227]. IGF [218, 203, 643, 219, 632]. IGF-I [218, 643, 203, 632]. IGF-II [632]. IGF-like [219]. IgM [626, 281]. II [625, 349, 147, 70, 685, 607, 499, 509, 396, 202, 105, 632, 500, 114, 189, 159, 600, 575, 585]. II-cells [202]. III. [116]. immature [98, 31, 355]. immobility [235, 24]. immune [46, 401]. immunity [231]. immunization [47]. immuno [204]. immuno-electron [204]. immunoassay [228]. Immunochemical [415, 63, 39, 84]. immunocompetence [275]. immunocytochemical [109, 319]. immunoglobulin [626, 281]. Immunohistochemical [536]. immunolocalization [668]. immunological [258]. Immunoreactive [290, 141]. immunoreactivity [249, 296]. immunosorbent [645]. Impact [547, 67]. impermeability [6]. implications [401]. imposed [8]. improves [5]. including [298]. Incorporation [610, 163, 238, 28, 29, 35, 85, 127, 154, 551]. increase [49]. increased [49]. increases [608]. increment [392]. incubation [18]. Index [61, 174, 327, 380, 128]. Indian [680]. indicate [544]. indicator [650]. indices [613]. individual [255, 339]. induce [153]. Induced [54, 58, 376, 382, 511, 604, 375, 402, 619, 447, 237, 439, 568, 305, 423, 509, 80, 314, 274, 161, 609, 205]. induces [369, 483]. inducible [63]. inducing [116, 573, 356, 311]. Induction [243, 24, 310, 456, 48, 122, 98, 319, 328, 507, 449, 478, 668, 150]. infestation [616]. Influence [686, 352, 584, 158, 237, 453, 250, 317, 651, 91, 330, 553, 156, 545, 639, 4, 399, 476, 14, 613, 443, 88, 150, 269, 448, 105]. influenced [472]. influences [220, 583, 671]. infused [680]. infusions [271]. ingested [88, 612]. ingestion [246]. Inhibition [350, 526, 609, 269]. inhibitor [633, 262, 388, 152]. inhibitors [329, 477]. inhibitory [209]. inhibits [152]. initiation [540]. injection [538, 395, 397]. injections [509]. inner [559]. inorganic [237, 534]. inplasma [576]. insubcellular [544]. Insulin [223, 218, 99, 203, 221, 140, 34, 75, 552, 141, 643, 651, 295, 518, 574, 219, 296, 632, 672, 222, 90, 87, 43]. insulin-immunoreactive [141]. Insulin-like [218, 552, 643, 651, 518, 574, 219, 296, 632]. Insulin-receptor [99]. insulin-secreting [43]. intake [253, 639, 141, 406]. integumental [566]. intensive [297]. Interaction [165, 81, 318]. Interactions [528, 210, 575]. interactive [175]. intermedia [530]. internal [42]. interplay [42]. Interrelationships [50, 151]. interrenal [331, 451, 42]. intestinal [391, 346, 368, 513, 294, 526, 71, 483, 347, 109, 460, 245, 677, 162]. intestine [58, 259, 230, 101]. Intracellular [180, 76, 183, 676, 421, 96]. Intrahepatic [504]. Intraperitoneal [395]. Investigation [358, 449]. Investigations [241, 4]. Involvement [181, 374, 211, 349]. involvements [434]. involves [609]. iodomelatonin [425, 313]. iodothyronine [559, 630, 585]. ion [535, 252, 312, 125, 561, 96]. ion-poor [312]. ionic [453, 423, 354, 65]. Ionoregulatory [335, 36]. ions [132]. iron [459, 49, 414]. ishikawai [518]. Island [22]. isoenzyme [273]. isoenzymes [555]. isoforms [533, 550]. isolated [139, 2, 410, 34, 560, 100, 286, 623, 221, 288, 530]. Isolation [39, 566, 48, 601, 263, 206, 328, 97]. isomerase [206]. isomorph [320]. isosafrol [84]. isosorbide [666]. isotocin [663]. isotonic [421]. isotope [620]. isozyme [147]. isozymes [598, 672]. Israel [547].

Japanese [433, 482, 592, 593, 615, 537, 456, 526, 37, 509, 507, 508, 18, 656]. japonica [433, 537, 456, 40, 37, 509, 507, 508, 18]. japonicus [573]. José [437]. Journal [21, 171, 276, 325, 378, 438]. junctions [602, 648]. juvenile [287, 362, 340, 636, 332, 674, 673, 639, 388, 256, 399, 117, 342, 364, 248, 148, 371, 163, 613, 497, 56, 314, 479, 132, 292, 523, 265]. juveniles [121, 462, 617, 238, 551].

kallikrein [152]. kallikrein-like [152]. Kaup [405]. keratan [251, 419]. keta [127, 336]. ketone [152, 435]. ketotestosterone [686, 357, 432, 619, 502, 189]. kidney [529, 432, 619, 570, 532, 16]. kidney-secreted [619]. kidneys [254]. kinase [203, 181, 621, 282]. kinetic [317, 52]. kinetics [481, 283, 168, 312, 88]. kisutch [633, 218, 604, 241, 217, 79, 298, 196, 479]. knifefish [511].

l [651, 593, 615, 686, 429, 32, 388, 476, 654, 14, 428, 662, 162]. L-aspartate [476]. L-gulonolactone [428]. L-thyroxine [162]. L.

[683, 471, 253, 287, 686, 566, 596, 255, 81, 140, 473, 153, 424, 403, 425, 108, 130, 462, 601, 544, 348, 434, 557, 115, 337, 552, 665, 131, 232, 220, 239, 628, 548, 295, 11, 12, 163, 551, 653, 60, 80, 105, 613, 652, 620, 621, 553, 616, 598, 672, 126, 59, 110, 554, 275, 245, 254, 512, 85, 292, 265]. labelled [348, 303, 163, 238]. labelling [77]. labrax [541, 368, 403, 448, 462, 75, 264, 364, 458, 406, 483, 586, 254]. lactate [358, 283]. lacustris [646]. lagged [371]. lake [646, 103, 594, 241]. Lampetra [40]. lamprey [118, 108, 49, 40, 19, 45, 43]. lampreys [464, 42, 41]. lapites [504]. large [119]. larvae [166, 346, 118, 481, 465, 473, 361, 309, 420, 475, 264, 464, 654, 246, 154, 431, 12, 405, 653, 406, 483, 586, 240, 113, 114, 363, 565, 467, 85, 383]. larval [577, 345, 387, 682, 393, 56, 292]. lateral [596, 130]. latipes [205]. lecithin [154, 652]. lecitotrophia [653]. lemon [24]. Lepomis [659]. leptocephali [251, 419]. leptocephalous [240]. lesions [101]. lethal [611, 335]. leucine [4]. Leuciscus [146]. leukocytes [169]. level [424, 256, 304, 629, 330, 372, 521, 422]. levels [641, 201, 647, 669, 625, 214, 451, 549, 36, 124, 32, 445, 420, 336, 47, 458, 37, 573, 295, 499, 518, 466, 56, 88, 586, 196, 500, 521, 150, 90, 87]. Leydig [456, 407]. LH [54]. LH-RH [54]. life [108, 561, 90]. like [658, 218, 249, 552, 152, 655, 643, 651, 224, 518, 574, 219, 296, 632, 90, 87, 109, 415]. Limanda [427]. limit [638]. Limitations [8]. limited [136]. Lindane [641]. linked [645]. Linnaeus [246]. linoleic [287, 30, 105]. linolenate [551]. linolenic [287, 362, 30, 105]. linseed [652]. lipase [100, 543, 587, 154]. Lipid [430, 557, 284, 303, 594, 652, 254, 287, 362, 128, 332, 473, 584, 142, 610, 595, 485, 266, 464, 369, 250, 11, 405, 80, 553, 321, 339, 265, 284]. Lipid-stimulated [430]. Lipids [80, 105, 674, 154, 472, 12, 653, 85, 343]. lipogenesis [223]. Lipolysis [303, 137]. lipophilized [508]. lipoprotein [433, 544, 579]. lipoproteins [111, 165, 250, 463]. liquid [232, 23]. liver [362, 658, 128, 214, 442, 674, 549, 403, 247, 256, 448, 48, 122, 165, 98, 63, 137, 402, 76, 414, 295, 120, 611, 551, 274, 612, 680, 435, 101]. Liza [238]. loach [416, 413]. localization [593, 566, 493, 319]. locomotor [360]. Long [519, 22, 477, 90]. Long-term [519]. longer [30]. longifilis [679]. longterm [18]. loricariid [404]. Low [333, 433, 32, 111, 165, 266, 131, 611]. lucius [473, 348, 126]. lumen [162]. lysine [545]. Lysis [615]. Lysozyme [366].

M [626, 281, 578]. M. [374, 232]. macrochirus [659]. Macrozoarces [446]. maculosus [338]. Magnesium [81, 411, 468, 104]. main [433]. mainly [585]. maintained [141, 267]. Maintenance [669, 672]. major [484, 543, 135, 550, 311, 159, 109, 156]. malabaricus [229]. malate [389]. male [201, 329, 477, 451, 158, 432, 619, 353, 499, 297, 500, 189, 407, 512]. males [14, 355, 607]. malic [52]. mammalian [152]. manipulations [458]. mannose [27]. mannose-receptor [27]. marine [362, 562, 6, 272, 10, 122, 445, 618, 135, 258, 684, 296, 278, 57]. marinus [108, 464]. marker [359]. marmoratus [478]. masou [518, 304, 201, 281, 304, 536, 225]. mass [135, 227]. masses [318]. masu [201, 495, 488, 281, 304, 536, 225]. matrix [334]. Maturation [368, 633, 201, 495, 329, 188, 116, 375, 408, 273, 374, 643, 144, 239, 573, 178, 396, 574, 507, 483, 196, 662, 356, 299, 204]. maturation-inducing [116, 573, 356]. maturation-promoting [375, 408]. maturational [38]. mature [98, 500, 512]. maturing [357, 518]. maturity [645, 492]. maxima [629]. maximus [287, 362, 340, 636, 166, 122, 462, 595, 587, 284, 303, 621, 624, 29, 321, 339]. McMurdo [571]. meal [635]. measurable [274]. Measurement [4, 267]. measurements [96]. mechanisms [180, 5, 199, 338]. medaka [504, 205]. mediated [391, 351, 512]. mediating [180]. Mediation [447, 183]. Mediterranean [684, 82]. medium [673]. melanin [176]. melanin-concentrating [176]. melanoptera [235]. Melatonin [671, 663, 426, 372, 564]. membrane [97, 77, 444]. membranes [153, 612, 501]. menhaden [584]. mesoderm [311]. mesoderm-inducing [311]. Mestres [437]. met-enkephalin [71]. Metabolic [548, 224, 363, 683, 51, 129, 318, 164, 627, 628, 135, 422, 452]. Metabolism [111, 139, 2, 127, 686, 465, 469, 403, 247, 485, 558, 83, 286, 457, 369, 115, 308, 14, 398, 207, 107, 47, 239, 450, 89, 163, 238, 551, 16, 635, 23, 435, 564, 28, 29, 35, 675, 257, 156]. metabolites [214]. Metallothionein [122, 601, 285]. metamorphosing [251, 464]. metamorphosis [49, 19, 419, 212]. methaemoglobinaemia [439]. Methallibure [609]. Methallibure-induced [609]. method [410, 364]. Methodologies [547]. methyl [476]. methylisoborneol [667]. methylparatyrosine [372]. methyltestosterone [521]. metomidate [337]. mevalonate [403]. microlepis [645]. microscopical [204]. microsomal [427, 16, 254]. microsomes [63, 402, 427, 16]. Midas [440]. midbrain [497]. migrating [290]. migration [336, 627, 606, 607]. milkfish [106, 121, 144]. milt [158]. Mineral [468]. minimum [149]. Misgurnus [413]. Mismatch [506]. Mitchill [291]. Mitochondria [436, 487, 581, 412]. Mitochondria-rich [436, 581]. Mitochondrial [52]. mixed [352]. mobilisation [428]. mobilization [665, 450]. mode [38, 194]. model [675]. modification [129, 301]. modified [410]. modulate [622]. modulating [455]. Modulation [475, 500, 474, 616, 181, 650, 192]. modulatory [391]. Molecular [527, 547, 178, 603, 670, 514, 346, 9, 486, 493, 402, 478]. monoamine [306, 609]. monoaminergic [33]. monoamines [372]. monodeiodinase [442, 142, 1, 114]. monooxygenase [546]. monophosphate [149]. Monosaccharides [146]. mordax [549]. morhua [410, 562, 164, 249, 638, 71, 628, 59, 110]. Morone [68, 399, 387, 374, 277, 56, 228]. Morphogenesis [43]. Morphological [271]. morphology [539, 485, 151]. morphometry [404, 134]. mortalities [362]. mortality [636]. mosquitofish [20, 66]. mossambicus [190, 331, 411, 468, 481, 309, 420, 475, 569, 243, 436, 431, 561, 267, 113, 114, 474, 530, 521, 366, 581]. Motility [398, 608, 146, 446]. mRNA [367, 218, 517, 569, 414, 644, 499, 586, 632, 629, 22, 500]. mRNAs [542]. MSH [190]. mucin [263]. mucosa [410]. mucus [440]. mudskipper [26, 69, 70]. mudskippers [578]. mullet [238]. mummichog [101]. muscarinic [447]. muscle [322, 259, 230, 391, 596, 469, 584, 95, 386, 138, 130, 610, 129, 318, 487, 99, 266, 679, 72, 597, 369, 71, 317, 304, 234, 452, 460, 677, 545]. mutant [216]. Mycteroperca [645]. mykiss [450, 641, 139, 669, 658, 562, 128, 214, 58, 259, 230, 376, 430, 341, 84, 381, 469, 223, 244, 365, 358, 567, 666, 124, 142, 95, 386, 138, 485, 558, 390, 271, 129, 560, 100, 137, 402, 476, 626, 252, 119, 392, 597, 663, 453, 207, 312, 195, 385, 53, 263, 319, 611, 551, 546, 96, 301, 202, 280, 219, 605, 576, 67, 623, 632, 314, 359, 123, 274, 191, 635, 377, 455, 288, 55, 112, 616, 206, 167, 192, 354, 333, 104, 344, 452, 307, 460, 355, 162, 132, 660, 395, 179]. myocardium [279]. myofibril [130]. Myofibrillar [679]. myosepta [343]. Myosin [130]. Myxine [247].

n [362, 154, 572, 288, 28, 29, 35, 476]. n-3 [362, 572, 288, 28, 29, 35]. n-6 [362, 288, 28, 29, 35]. n-9 [154]. N-methyl-D [476]. Na/H [474]. naccarii [360]. NaCl [409]. NAD [52]. NADH [237]. NaHCO [271]. naphthoflavone [84, 63, 319, 555, 668]. Natriuretic [526, 618, 92, 36, 197]. natural [606, 607]. nature [417]. nebulosus [501]. necessary [195]. Negaprion [24]. nerves [249]. nest [619]. neural [511]. neurally [391]. neurally-mediated [391]. Neuroendocrine [41, 193, 494]. neurohypophysial [527, 525]. Neuroimmunological [190]. neurons [497]. neuropeptide [38]. neurotensin [71]. neutral [610, 595, 119]. Newfoundland [628]. newly [595]. newly-weaned [595]. NH [578]. night [306]. Nile [630]. nilotica [370, 30]. niloticos [150]. niloticus [342, 353, 630, 523, 585]. ninespine [51]. nitric [666, 388, 589]. nitrite [237, 439]. nitrite-induced [237, 439]. nitrogen [346, 255]. Nitrogenous [103]. nitroprusside [388]. nobilis [589]. non [128]. non-esterified [128]. nonanadromous [664]. nonaromatizable [79]. nonautonomous [655]. nonnutrients [440]. nonphagocytic [27]. nonylphenol [524]. nor [77]. noradrenaline [301]. norepinephrine [442]. normal [36]. northern [126, 461]. Nos [173, 326, 62, 379, 589]. NPY [38]. nuclear [614, 274, 444, 542]. nuclei [315]. nucleoside [50, 237]. number [169]. nutrient [639]. nutrition [136, 363]. nutritional [568].

O [370, 50]. O. [342, 353, 150]. Occurrence [357]. ocean [446]. ocellatus [316, 332, 674, 673, 248, 213, 316, 261, 134]. October [378, 438]. Odontesthes [661]. oil [362, 584, 652]. old [295]. olfactory [319, 444, 668]. Oligocottus [338]. olivaceus [656]. Oncorhynchus [633, 641, 201, 139, 127, 647, 669, 658, 562, 128, 214, 58, 259, 230, 376, 430, 341, 84, 381, 469, 223, 244, 365, 218, 567, 604, 666, 124, 142, 95, 386, 138, 485, 558, 390, 271, 129, 560, 100, 137, 402, 476, 626, 252, 119, 392, 336, 597, 663, 453, 207, 312, 241, 195, 385, 53, 263, 305, 148, 217, 450, 319, 423, 611, 518, 551, 546, 96, 301, 281, 202, 280, 219, 304, 290, 605, 576, 67, 623, 632, 314, 359, 79, 298, 123, 274, 191, 635, 196, 377, 455, 288, 55, 112, 616, 206, 167, 192]. Oncorhynchus [354, 333, 479, 104, 344, 452, 307, 460, 536, 355, 162, 132, 225, 660, 395, 179]. Oncorhyncus [358]. one [562, 375, 77, 78, 573, 295, 396, 356, 501, 600, 179, 510]. one- [295]. ones [502]. only [612]. onset [606, 607]. Ontario [241]. Ontogeny [517, 631, 577, 361, 109]. onwards [462]. Oocyte [116, 633, 68, 374, 243, 178, 574, 662, 179]. oocytes [375, 77, 542]. opercular [422]. Opsanus [83, 286, 89]. Optical [60]. options [527]. oral [54, 156]. Oreochromis [30, 190, 331, 411, 468, 481, 309, 420, 475, 569, 243, 436, 370, 342, 431, 353, 561, 297, 630, 267, 113, 114, 474, 530, 521, 150, 366, 581, 523, 585]. organ [358, 432, 319, 657, 668]. organization [129, 452]. organs [486, 73]. origin [664]. ornithine [322]. Oryzias [504, 205]. Osmerus [549]. osmolality [139, 522]. Osmoregulation [69, 70, 293, 527, 253, 198, 590]. Osmoregulatory [194, 523]. osmotic [4, 540]. Osteichthyes [653]. other [582]. otolith [576]. outer [559]. ovalbumin [76]. ovarian [649, 315, 625, 491, 560, 445, 643, 78, 509, 605, 662, 671, 206, 356, 372, 609, 205, 25]. ovariectomy [307]. ovaries [203, 115, 157, 308]. ovary [455, 366]. Overripening [349]. ovulated [349]. ovulation [491, 145, 54]. oxidase [352, 428, 306]. Oxidation [2, 266, 237]. oxidative [613]. oxide [666, 388, 278, 571, 589]. Oxygen [8, 472, 624, 647, 669, 94, 421, 237, 614, 548, 360, 423, 461, 65]. oxygenated [208]. oxygenation [269]. oxynticopeptic [562].

P [71, 52, 372, 84, 249]. P-450 [84]. p-chlorophenylalanine [372]. P-like [249]. P450 [310, 98, 63, 147, 546]. P4501A1 [319]. P4503A [415]. P4503A-like [415]. P450IA [63]. Pagothenia [8, 272, 582]. Pagrus [445, 484, 543, 135, 550, 311, 159, 109, 401]. pancreas [43]. pancreatic [346, 368, 141, 110, 109]. Paralichthys [656]. Paraliparis [334]. parameters [337, 646, 617]. paraquat [9]. parents [440]. Park [347]. parr [329, 384, 539, 295, 423, 245, 512, 225]. parr-smolt [295, 245, 225]. pars [530]. Partial [53, 658, 566, 97, 263, 328, 684, 13]. Participation [189]. partition [537]. parvalbumin [130]. pathogenic [615]. pathway [572]. pathways [571]. pattern [128, 214, 413]. patterns [506, 416, 147]. PCB [150, 565]. pejerrey [661]. pelagic [594]. pepsin [429]. peptidases [346]. peptide [36, 526, 90, 87, 311]. peptides [618, 224, 197, 525]. Perca [654, 665]. perch [315, 654, 374, 665]. performance [95, 583, 132, 64]. perfused [529, 680]. pericardial [302]. pericardioperitoneal [302]. perifused [38]. period [654, 295]. Periophthalmodon [578]. Periophthalmus [26]. permeability [561]. permeabilized [412]. peroxidase [566]. peroxisomal [39]. perspective [490]. perspectives [224]. pesticide [512]. Petromyzon [108, 464]. pH [410, 608, 32, 50, 170, 611, 96, 676, 624, 460, 64]. pH-dependent [410]. pH-static [410]. phagocytosis [280]. pharmacological [441]. pharmacologically [58]. phase [42]. phases [562]. Phe [152]. Phe-Phe-Arg-chloromethyl [152]. phenotype [79]. pheromonal [512]. pHi [563]. phosphatase [273, 359]. phosphate [534]. phosphatidylcholine [610, 303]. phosphatidylcholines [402]. phospholipase [484, 587, 550, 109]. phospholipid [340, 462, 595, 28, 29]. phospholipids [127, 612]. phosphorus [580]. phosphorylase [34, 650]. photoaffinity [77]. photoperiod [625, 128, 33, 306]. Photoperiodic [458]. photoperiods [477]. phylogenetic [658]. physical [153]. physicochemical [612, 59]. Physiological [646, 591, 638, 526, 617]. Physiology [277, 666, 197, 520]. phytase [635]. picomolar [313]. pike [473, 348, 126]. pineal [426, 60]. pituitaries [191]. Pituitary [531, 201, 38, 181, 494, 241, 185, 499, 509, 202, 492, 290, 519, 500, 189, 18, 292, 575, 184, 204]. plaice [48]. Planas [437]. Plasma [124, 97, 463, 295, 88, 196, 90, 87, 683, 633, 433, 625, 128, 214, 451, 681, 608, 365, 153, 262, 604, 142, 4, 75, 531, 158, 445, 119, 273, 336, 243, 337, 663, 398, 152, 606, 607, 155, 458, 170, 518, 281, 466, 522, 267, 296, 289, 612, 672, 354, 372, 579, 344, 501, 446, 269, 585]. Plasticity [487, 511]. Platichthys [531]. plecostomus [404]. Pleuronectes [608, 534]. PmGTH [159]. pollutants [611]. polymorphism [94]. polypeptide [513, 71]. Polyunsaturated [622, 35, 675, 282, 127, 485, 348, 163, 238, 30, 613, 28, 29, 265]. Pomacentridae [445]. Poor [394, 312]. population [22]. porcine [217]. porgy [373, 382, 685, 135, 510]. possesses [304]. possible [251, 449, 668, 199, 349]. post [294]. post-smolt [294]. postlarval [610]. postovulatory [349]. Potassium [399, 439]. potency [493]. potential [70]. pout [446]. practical [673]. pre [635]. pre-treatment [635]. precociously [518]. precursor [504, 656]. Preface [177]. pregn [77]. pregnen [375, 502, 77, 78, 573, 396, 356, 501, 600, 179, 510]. pregnene [600]. Preliminary [4]. Preparation [216, 410]. prepared [508]. Presence [375, 109, 681, 544, 285]. pressure [235, 10, 540]. presumptive [270]. pretreated [475]. Primary [367, 558, 602, 648, 222, 227]. probable [311]. processes [288]. produces [617]. Production [426, 287, 340, 608, 549, 158, 152, 573, 492, 574, 205, 600]. products [266]. profile [433, 320, 297]. profiles [577, 158, 606, 607, 225]. progestin [77]. Progestogen [115]. proglucagon [514]. prolactin [139, 218, 644, 522, 191, 521, 661]. proliferation [556]. proliferative [91]. promoting [375, 408]. pronephric [434]. pronephros [27]. Properties [285, 599, 567, 216, 48, 98, 487, 334, 622, 232, 317, 60, 612, 347, 624, 52, 59, 366]. Propylthiouracil [604, 19]. Propylthiouracil-induced [604]. prostaglandin [340]. prostanoids [301]. protandrous [373, 382, 685, 510]. proteases [599, 482, 347]. Protein [255, 345, 334, 239, 120, 621, 368, 465, 81, 142, 386, 108, 544, 569, 619, 181, 659, 14, 622, 371, 611, 219, 522, 620, 406, 22, 672, 333, 521, 637, 210, 282, 545]. protein-magnesium [81]. Protein-nitrogen [255]. protein-turnover [545]. Proteins [53, 683, 415, 218, 679, 364, 504, 413, 109]. Proteoglycans [350]. Proteolysis [364]. proteolytic [232]. Prothrombin [684]. protogynous [478]. Proton [532]. Psetta [629]. pseudobranch [358]. Pseudolabrus [573]. Pseudopleuronectes [158]. Pubertal [451, 407]. Publisher [678]. pump [4]. pumps [532]. punctatus [136, 33, 655, 268, 667]. Pungitius [51]. pur [337]. Purification [315, 116, 100, 484, 543, 550, 52, 59, 159, 228, 262, 138, 645, 366, 13, 660]. putative [315, 77, 444]. pyloric [484, 59, 110]. pyrene [83, 286, 479].

Qualitative [129, 389]. quality [7, 646, 553]. Quantification [270, 84, 243, 227]. Quantitative [108, 389, 126].

R. [4]. R5020 [77]. rabbitfish [347]. radioimmunoassay [159, 15]. radiolabelled [623]. Rafinesque [624]. Rainbow [602, 648, 641, 310, 39, 139, 2, 658, 625, 236, 38, 442, 529, 58, 259, 230, 376, 430, 341, 84, 381, 469, 223, 352, 244, 365, 27, 358, 549, 567, 36, 666, 124, 142, 95, 4, 386, 32, 485, 558, 390, 111, 165, 271, 129, 487, 560, 100, 137, 402, 476, 626, 252, 237, 119, 392, 494, 457, 597, 369, 76, 663, 453, 207, 195, 385, 263, 5, 31, 570, 450, 319, 611, 551, 546, 96, 396, 202, 280, 676, 219, 91, 418, 480, 503, 605, 576, 67, 623, 632, 314, 359, 16, 221, 123, 274, 444, 191]. rainbow [631, 635, 278, 289, 330, 662, 671, 377, 455, 288, 55, 112, 616, 206, 167, 23, 192, 668, 356, 333, 104, 564, 344, 452, 28, 409, 25, 307, 460, 556, 600, 162, 132, 269, 660, 65, 64, 395, 179]. raised [85]. Raja [350]. rat [63]. rate [236, 235, 10, 388, 318, 638, 561, 135, 422]. rates [322, 620, 234, 333, 545]. ratio [119, 117]. ration [330, 672, 90]. ratios [340]. raven [283, 257]. re [142]. re-feeding [142]. Reabsorption [570]. reactivity [63]. reared [659, 561, 297, 631]. rearing [424, 598]. receptor [315, 410, 513, 27, 99, 169, 503, 501, 542, 575, 660]. receptors [454, 505, 493, 203, 494, 643, 77, 441, 444, 642]. recombinant [216, 217]. Recovery [439]. recruitment [307]. red [454, 332, 674, 673, 487, 484, 543, 50, 248, 213, 261, 496, 550, 377, 167, 134, 311, 159, 156, 269, 520, 109]. redox [283]. reduced [208, 141, 338]. reduction [424]. reef [235]. refeeding [359, 435, 585]. Referees [21, 171, 276, 325, 378, 438]. reflex [472]. regard [42]. regimes [625, 128, 345, 320]. Regional [16, 23]. regions [55]. regulate [629]. regulated [585]. regulating [167]. Regulation [331, 503, 497, 642, 495, 190, 236, 515, 511, 211, 516, 34, 491, 494, 453, 563, 3, 89, 423, 676, 591, 586, 632, 209, 22, 534, 189, 354, 501, 355, 670, 212]. regulatory [185, 52]. related [566, 425, 568, 590, 635, 55, 112, 210]. relation [527, 166, 414, 499, 363, 554]. Relationship [260, 365, 595, 135, 576, 445, 398]. relationships [415, 597, 219]. release [190, 180, 494, 568, 183, 618, 191, 640, 575]. releasing [608, 188, 666, 493, 182, 158, 488, 185, 258, 661, 186, 575]. relevant [252]. renal [555, 534]. renin [133, 528]. renin-angiotensin [133, 528]. repeat [445]. repeated [509]. replacement [242]. reproduction [477, 200, 42, 324, 41, 490]. reproductive [625, 448, 203, 476, 187, 147, 323, 458, 634]. Requirement [121, 106, 248, 149, 104]. rerio [542]. reserves [597]. reservoir [67]. resistance [231]. resistant [359]. resorption [314, 359]. resources [146]. respect [322]. respectively [578]. respiration [404, 354]. respiratory [229, 591, 335]. respirometer [7]. respond [548]. Response [346, 673, 409, 592, 188, 388, 434, 417, 264, 447, 31, 268, 546, 91, 616, 167, 46, 637, 565, 90]. responses [391, 351, 424, 271, 51, 164, 472, 371, 519, 480, 591, 406, 335, 69]. responsible [283]. responsiveness [407, 257]. restricted [214, 90]. restructuring [402]. resulting [79]. reticulum [412]. retinoid [642]. reversal [373, 478]. review [200, 194]. RH [54]. Rhodeus [316]. rhythm [143, 37]. rhythms [289, 330]. rich [287, 436, 581]. Richardson [36, 65, 64]. ring [559]. river [540]. RNA [322, 371]. roach [601, 115]. Role [402, 316, 480, 419, 113, 114, 537, 251, 515, 34, 42, 450, 133, 449, 167]. Root [262]. rose [316]. rostrata [34]. routine [591]. RU [91]. rudd [115]. Russian [470]. Rutilus [601, 115]. rutin [136].

S [48, 220, 501, 555, 101]. S-transferase [555, 101]. sablefish [54]. sac [465]. saffron [522]. salar [683, 329, 322, 588, 255, 294, 384, 539, 584, 424, 429, 386, 279, 388, 98, 266, 544, 273, 337, 552, 414, 147, 131, 117, 250, 300, 47, 524, 239, 644, 548, 295, 583, 60, 620, 598, 672, 46, 149, 637, 435, 275, 35, 321, 339, 245, 580, 231, 512, 677, 343]. salinities [644]. Salinity [56, 166, 332, 69, 590, 521, 321, 339, 338]. Salmo [683, 459, 329, 322, 2, 588, 255, 294, 384, 27, 539, 584, 36, 424, 429, 4, 386, 279, 32, 388, 98, 266, 544, 273, 337, 552, 414, 646, 147, 131, 195, 117, 250, 300, 232, 47, 5, 31, 524, 643, 239, 644, 548, 295, 583, 324, 60, 620, 598, 672, 46, 149, 637, 435, 275, 28, 35, 321, 339, 25, 245, 580, 231, 335, 512, 677, 65, 64, 343]. salmon [683, 633, 201, 495, 329, 322, 127, 647, 489, 588, 255, 294, 384, 539, 515, 584, 218, 604, 424, 429, 386, 279, 388, 98, 266, 544, 273, 336, 337, 488, 552, 414, 147, 241, 131, 117, 250, 300, 47, 524, 239, 644, 548, 169, 305, 295, 148, 583, 217, 518, 281, 509, 492, 60, 304, 522, 290, 620, 664, 79, 298, 221, 196, 478, 598, 672, 46, 149, 479, 637, 435, 299, 275, 35, 321, 339, 536, 245, 580, 231, 512, 677, 184, 225, 343]. salmonid [313]. salmonids [211, 194, 90, 87]. salt [36, 543, 526]. salt-activated [543]. salts [237]. saltwater [253]. Salvelinus [253, 291, 686, 94, 145, 557, 141, 77, 78, 3, 371, 80, 105, 613, 652, 226, 422, 265]. sampling [291]. sapidissima [627]. saturation [237]. saxatilis [68, 399, 387, 277, 56, 228]. scale [359]. Scardinius [115]. scavenger [27]. scavenger- [27]. schlegeli [373, 382, 685, 510]. schlosseri [578]. Sciaenops [332, 674, 673, 248, 213, 261, 134]. Scophthalamus [166]. Scophthalmus [287, 362, 340, 636, 122, 462, 595, 587, 284, 621, 624, 29, 321, 339]. Scopthalmus [303]. sculpin [540, 338]. Scyliorhinus [82]. sea [331, 346, 368, 596, 584, 403, 425, 283, 517, 448, 462, 75, 484, 543, 264, 464, 337, 131, 364, 232, 458, 426, 11, 12, 163, 551, 550, 406, 483, 586, 296, 133, 254, 257, 156, 109]. Seabream [547, 599, 658, 246, 572, 617, 393, 496, 553, 311, 159]. season [381]. Seasonal [625, 45, 169, 296, 22, 278, 452, 446, 469, 487, 557]. Seasonality [471]. seatrout [501]. seawater [168, 563, 195, 277, 148, 583, 561, 226, 354]. seawater-adapted [563]. secreted [619]. secreting [519, 43]. secretion [633, 410, 562, 38, 430, 429, 560, 476, 181, 185, 141, 353, 396, 443, 189, 18, 227, 520]. sedimentation [93]. segments [284, 303]. semen [260, 381]. seminal [681, 608, 398, 53, 161, 446]. Senegal [682, 405]. senegalensis [682, 405]. sensitive [650]. sensitivity [169]. September [547]. Sequence [656, 367, 176, 644]. sequestration [25]. series [596, 35]. serotonergic [641]. serotonin [582, 447, 119, 112, 23, 306, 575]. serotonin-induced [447]. Sertoli [407]. Serum [225, 641, 108, 334, 19, 37, 572]. set [638]. sex [633, 373, 158, 622, 117, 353, 478]. sexual [329, 451, 188, 144, 239, 396, 492, 196]. shad [627]. shape [595, 601, 598]. shark [302, 235, 591, 24]. shift [369]. shifts [153]. shock [637]. Short [54, 477, 260]. short-term [260]. sialic [161]. Siberian [357, 428, 88]. Siganus [347]. signal [175]. significance [526]. Sillago [37]. Siluriforms [679]. Silurus [157]. Similarities [562]. Singi [160]. single [530]. Site [252]. sites [425, 182, 492, 313, 274, 581, 338, 397]. situ [529]. Size [67, 625, 229, 597, 614, 556]. skate [350]. skeletal [99, 266, 72]. skin [566, 485, 558, 207]. slices [442]. slightly [612]. slow [72, 556]. slow-growing [556]. smelt [549, 278]. smolt [294, 384, 539, 295, 245, 225]. smoltification [515, 273, 524, 196, 479, 536]. smolts [304, 620]. smooth [259, 230, 391, 71, 677]. snailfish [334]. snapper [401]. social [511]. Sodium [535, 388, 399]. soft [335]. sole [682, 405]. Solea [682, 405]. somatic [576]. somatolactin [517, 336, 290, 196, 661, 520, 519]. somatolactin- [519]. somatostatin [430, 671, 43]. somatostatin- [43]. somatostatin-14 [671]. somatotrope [141]. somatotropin [217]. Some [232, 347, 381, 272, 193, 264, 107, 617, 571, 366]. Sound [571]. source [47, 46, 149]. sources [368]. southern [45, 19]. soybean [652, 635]. sp [251, 419]. Sparidae [445, 246, 653]. sparids [599]. Sparus [599, 658, 596, 425, 517, 246, 154, 617, 11, 12, 163, 551, 393, 296, 553]. Spawning [369, 445, 143, 336, 654, 606, 607, 316, 37, 573, 290]. special [42, 385, 404]. species [63, 402, 645, 1, 71, 313, 590, 461]. Specific [432, 179, 322, 152, 20]. specificity [367, 575]. spectrometry [227]. spent [290]. sperm [647, 669, 608, 398, 501, 446]. spermatogenesis [456, 209]. spermatozoa [97, 146, 398, 53]. spermiation [478, 510]. spermine [613, 483]. spermmotility [540]. Sphyrna [591]. spined [477, 432, 619]. spleen [621]. splenic [434, 67]. spontaneous [360]. spotted [501]. Squalus [209]. stable [620]. stage [294, 448, 654, 12, 406, 210]. stage-related [210]. stages [459, 561, 662, 600]. Stanniocalcin [168, 199]. starch [148]. starvation [653, 519, 576, 90]. state [153]. static [410]. status [81, 584, 511, 211, 476, 568, 385, 428, 635, 354, 269, 65]. steelhead [669]. stenohaline [293]. Steroid [227, 633, 116, 158, 626, 622, 606, 607, 77, 37, 574, 16, 444, 356, 205]. Steroidogenesis [157, 649, 470, 451, 605, 671]. steroidogenic [662]. steroids [495, 451, 349, 445, 208, 374, 42, 353]. stickleback [51, 318, 432, 619]. sticklebacks [477]. stimulate [145, 396]. stimulated [410, 562, 430, 137, 69, 45, 120, 510]. stimulates [74, 373, 294, 223, 685]. Stimulation [160, 680, 292]. Stimulatory [209, 476]. Stizostedion [323]. stocking [617]. stocks [664]. stomach [410, 249]. Storage [343, 260, 557, 414]. Strain [94]. strains [148, 556]. strategies [389]. Stress [424, 274, 592, 291, 566, 351, 124, 659, 268, 617, 267, 613, 123, 565]. Stress-induced [274]. stress-related [566]. stresses [280]. striped [68, 399, 387, 277, 56, 228]. strong [304]. Structural [541, 601]. studied [410]. Studies [38, 520, 658, 513, 374, 449, 209]. study [559, 426, 313, 23, 372, 73, 109, 536, 204]. sturgeon [649, 470, 357, 256, 361, 360, 472, 428, 394, 443, 88]. sub [611]. sub-lethal [611]. subdermal [334]. subforms [9]. Sublethal [512, 335, 132]. submitted [21, 171, 276, 325, 378, 438]. subsequent [653]. substance [681, 249, 71]. substances [152]. Substrate [308]. substrates [358]. subunit [367, 533, 384, 569, 500, 536, 670]. subunits [509, 492, 498]. sulfatase [251]. sulfate [251, 47, 220, 240, 419, 46]. sulfates [208]. Sulfation [356]. sulfur [395]. sulphate [396]. summer [253]. Summertime [107]. superoxide [49]. supplementation [462, 483, 635, 545]. supplemented [673]. supply [346]. supplying [249]. suppresses [373]. suppression [141]. surface [440]. survey [1]. survival [387, 246]. swelling [421]. swim [311]. swim-bladder [311]. swimbladder [71]. Swimming [132, 7, 583, 591, 422, 64]. Synbranchidae [478]. Synbranchus [478]. synergistic [136]. synthase [589]. Synthesis [502, 355, 86, 386, 345, 208, 14, 120, 620, 209, 455, 672, 192, 333, 210, 73]. synthetase [401]. synthetic [77, 571]. system [495, 98, 147, 301, 133, 528, 401]. systems [641, 352, 211, 418].

T [128, 32, 409, 585]. tanks [424]. target [432]. tarichi [103]. Tartrate [359]. Tc1 [655]. Tc1-like [655]. technique [620]. Teleost [658, 541, 562, 415, 86, 8, 272, 193, 491, 309, 33, 502, 312, 614, 436, 71, 323, 651, 37, 573, 618, 89, 219, 135, 418, 480, 278, 571, 474]. teleostean [566, 559]. Teleostei [229, 146, 478]. teleosts [6, 486, 10, 72, 1, 258, 497, 684, 133, 151, 73, 57]. Temperature [153, 259, 230, 577, 469, 279, 402, 320, 143, 83, 286, 33, 597, 19, 131, 548, 611, 613, 91, 56, 22, 288, 624, 667, 333, 545]. temperature-induced [402]. temperatures [631, 598, 677, 85]. Temporal [142, 628]. tension [50]. tensions [647, 669]. term [260, 519]. testes [208, 157, 600]. Testicular [208, 407, 649, 506, 373, 456, 499, 52, 501, 227]. testis [451, 632, 210, 282]. testosterone [365, 685, 117, 499, 298, 156, 225]. testudineus [315]. tetrachlorobiphenyl [480]. their [118, 454, 251, 211, 208, 502, 458, 11, 405, 96, 440, 57]. themale [267]. therapy [242]. thermal [51, 129, 318, 389]. thermally [402]. thermally-acclimated [402]. thewhite [545]. thirty [1]. thirty-three [1]. three [477, 118, 10, 94, 645, 432, 619, 1, 313, 624, 461]. three-spined [477, 432, 619]. threespine [318]. thrombocytes [676]. thromboplastin [684]. Thyroid [211, 324, 629, 57, 299, 315, 142, 19, 45, 241, 213, 261, 56, 631, 113, 114, 205, 642, 670]. thyroid/retinoid [642]. thyroidal [584, 211]. thyronine [686, 32, 651, 662, 162]. thyrotropin [185, 629, 670]. thyrotropin-releasing [185]. thyroxine [641, 604, 19, 162, 225]. tiburo [591]. tidepool [338]. tilapia [190, 331, 74, 481, 454, 176, 420, 475, 569, 243, 183, 370, 342, 431, 353, 499, 561, 297, 630, 267, 113, 114, 498, 530, 521, 150, 366, 581, 412, 523, 585, 30]. time [684, 289, 330]. timing [627]. tipped [235]. Tissue [218, 371, 20, 322, 636, 341, 451, 513, 367, 384, 318, 154, 241, 155, 628, 428, 444, 631, 52, 227]. Tissue-specific [20]. tissues [32, 285, 273, 370, 96]. titration [410]. toadfish [83, 286]. tocopherol [544, 300, 572, 613]. tocopherolbinding [544]. tolerance [332, 148, 590]. tolerate [26, 578]. tonic [235, 24]. total [604, 464, 12, 405, 612, 339]. toxicokinetics [479]. tPRL [521]. trace [47]. tract [361, 552, 284, 303]. training [95]. transaminase [9]. transcription [509]. Transcriptional [516]. transduction [175]. Transepithelial [70]. transfer [215, 252, 337, 644]. transferase [352, 555, 101]. transferases [48]. transferred [253, 354]. Transferrin [414, 459]. transformation [683, 295, 245, 225]. translocator [542]. transmontanus [256, 443]. Transport [300, 78, 683, 244, 76, 614, 534, 245, 461, 338]. transporters [535]. transposable [655]. treated [84, 3]. treatment [608, 9, 285, 14, 5, 31, 499, 428, 359, 635]. treatments [312, 564]. Trematomus [351]. triacylglycerol [100, 303]. trigger [540]. triglycerides [673]. trihydroxy [573, 501, 510]. triiodo [686, 32, 651, 662, 162]. triiodo-L-thyronine [662]. Triiodothyronine [195, 160, 604, 387, 19]. trimethylamine [278, 571]. triphosphatases [69]. triphosphate [50]. triphosphates [237]. triploid [291, 217, 422]. triploids [297]. tritiated [88]. tropical [461]. trout [641, 459, 310, 39, 139, 2, 669, 291, 658, 625, 236, 38, 442, 529, 58, 259, 230, 391, 376, 430, 341, 84, 381, 469, 223, 352, 244, 365, 27, 262, 358, 567, 36, 666, 124, 142, 95, 4, 386, 138, 535, 32, 485, 558, 390, 111, 165, 271, 129, 487, 99, 560, 100, 137, 402, 421, 476, 626, 145, 252, 237, 119, 392, 494, 457, 597, 369, 76, 663, 97, 646, 453, 207, 563, 195, 385, 152, 53, 232, 263, 317, 141, 5, 31, 643, 77, 78, 3, 570, 170, 371, 120, 450, 319, 423, 611, 551, 546]. trout [96, 301, 324, 396, 202, 280, 466, 676, 219, 91, 418, 480, 503, 605, 576, 67, 623, 632, 314, 359, 16, 221, 123, 274, 444, 191, 631, 635, 289, 330, 662, 671, 270, 377, 455, 288, 55, 112, 616, 206, 167, 23, 192, 602, 648, 668, 356, 354, 333, 104, 564, 422, 344, 452, 28, 409, 25, 307, 460, 556, 600, 335, 162, 132, 269, 660, 65, 64, 395, 179]. trutta [646, 232, 459, 195, 232, 643, 324, 335]. trypsin [429, 406, 586, 598, 672]. tryptophan [121, 119]. tshawytscha [647, 305, 148, 290]. turbot [287, 362, 340, 636, 166, 122, 462, 610, 595, 587, 284, 303, 629, 621, 624, 29, 321, 339]. Turkey [103]. Turnover [544, 465, 119, 385, 611, 306, 545]. two [527, 599, 188, 243, 71, 573, 295, 618, 202, 418, 664, 590, 631, 662, 85]. two-year [295]. type [600, 585]. types [610]. Tyrosine [385, 203, 192].

Ultrastructural [202, 563, 652]. ultrastructure [562, 407]. unchanged [266]. undersulfated [240]. underyearling [156]. unfed [405]. unilateral [307]. unique [433]. unsaturated [553]. Uptake [411, 298, 481, 538, 294, 8, 420, 475, 286, 220, 623, 151, 314, 667, 356, 199, 412]. Urea [467, 103, 570, 278, 571, 338]. ureogenesis [680]. ureogenic [89]. uric [681]. urine [316]. use [645]. using [410, 128, 4, 77, 620, 478, 23, 149]. utilisation [639]. Utilization [148, 653, 383, 577, 164, 394].

values [684]. Variable [320]. variants [478]. variation [452]. variations [322, 625, 557, 654, 53, 278, 112, 306, 147]. variety [344]. various [647, 436, 280, 55, 57, 212]. varying [598]. Vascular [525]. vasoactive [529, 513, 582, 152, 71]. Vasoactivity [301]. vasoconstriction [447]. vasopressin [474]. vasotocin [198, 663]. velocity [93]. venous [244]. ventilation [236, 170]. ventilatory [472]. versus [320, 193]. vertebral [350]. very [433]. vesicle [161]. vessels [582]. via [27]. vitamin [136, 365, 272, 266, 47, 46, 149, 275, 231]. vitellogenesis [633, 300, 524, 449]. vitellogenic [25]. Vitellogenin [243, 382, 645, 458, 328, 455, 228, 15, 25, 13]. vitreum [323]. vitro [649, 647, 669, 560, 476, 185, 573, 671, 677, 470, 391, 451, 357, 153, 138, 390, 145, 208, 157, 308, 374, 303, 1, 220, 426, 396, 135, 605, 449, 191, 407, 460, 269]. viviparus [14]. vivo [272, 551, 629, 348, 447, 76, 374, 426, 163, 238, 96, 30, 620, 449, 377]. Vol [326, 62, 173, 379]. Volume [3, 614, 96, 267, 82].

Walbaum [259, 230, 341, 158, 277, 450, 546, 404, 274, 288, 616]. warm [254]. warm- [254]. waste [103]. water [331, 315, 411, 468, 584, 424, 7, 337, 312, 131, 125, 561, 267, 293, 133, 104, 634, 580, 65, 64]. water-borne [104]. waterborne [459]. waters [335]. weaned [595]. weaning [368, 462]. weights [87]. wels [157]. wheat [581]. White [304, 138, 130, 256, 361, 679, 597, 369, 374, 317, 394, 443]. whiting [37]. whole [56, 132]. whole-animal [56]. wild [596, 606, 607, 628, 238, 664, 126, 299, 536]. wild-caught [238]. window [60]. winter [253, 158, 611, 22, 534]. without [150]. wrasse [573].

x [150, 8]. X-cell [8]. xenobiotic [16].

year [295]. yellowtail [608]. Yolk [665, 465, 113, 85, 383]. yolk-sac [465]. yolksac [12]. young [290, 440].

Zealand [606, 607]. zebrafish [542]. zinc [635, 640]. zinc-related [635]. Zn [412, 285]. Zoarces [14]. zonation [247].

References

Leatherland:1990:HMA

[1] J. F. Leatherland, P. K. Reddy, and T. J. Lam. Hepatic 5′-monodeiodinase activity in teleosts in vitro: A survey of thirty-three species. Fish Physiology and Biochemistry, 8(1):1–10, January 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004426.

Bell:1990:OGD

[2] J. Gordon Bell and Colin B. Cowey. Oxidation of glutathione during hydroperoxide metabolism in isolated hepatocytes of rainbow trout (Salmo gairdneri). Fish Physiology and Biochemistry, 8(1):11–17, January 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004427.

Marshall:1990:VRG

[3] William S. Marshall, Sharon E. Bryson, and M. Michelle Sapp. Volume regulation in glutathione-treated brook trout (Salvelinus fontinalis) erythrocytes. Fish Physiology and Biochemistry, 8(1):19–28, January 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004428.

Fauconneau:1990:MPL

[4] Benoit Fauconneau and Sophie Tesseraud. Measurement of plasma leucine flux in rainbow trout (Salmo gairneri R.) using osmotic pump. Preliminary investigations on influence of diet. Fish Physiology and Biochemistry, 8(1):29–44, January 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004429.

Madsen:1990:CTI

[5] Steffen S. Madsen. Cortisol treatment improves the development of hypoosmoregulatory mechanisms in the euryhaline rainbow trout, Salmo gairdneri. Fish Physiology and Biochemistry, 8(1):45–52, January 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004430.

Boyd:1990:ABI

[6] R. B. Boyd, J. Atkin, and A. L. DeVries. Absence of binding and impermeability to ferritins of gill endothelium in marine teleosts. Fish Physiology and Biochemistry, 8(1):53–60, January 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004431.

Gehrke:1990:RCW

[7] Peter C. Gehrke, Larry E. Fidler, and David J. Randall. A respirometer with controlled water quality and computerized data acquisition for experiments with swimming fish. Fish Physiology and Biochemistry, 8 (1):61–67, January 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004432.

Davison:1990:OUA

[8] William Davison, Craig E. Franklin, and Peter W. Carey. Oxygen uptake in the Antarctic teleost Pagothenia borchgrevinki. Limitations imposed by X-cell gill disease. Fish Physiology and Biochemistry, 8(1):69–77, January 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004433.

Deer:1990:EPT

[9] K. Aranka Deér, János Nemcsók, and Béla Szajáni. Effects of paraquat treatment on fish transaminase molecular subforms. Fish Physiology and Biochemistry, 8(1):79–83, January 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004434.

Fritsche:1990:EHB

[10] Regina Fritsche. Effects of hypoxia on blood pressure and heart rate in three marine teleosts. Fish Physiology and Biochemistry, 8(1): 85–92, January 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004435.

Mourente:1990:EBDa

[11] G. Mourente and J. M. Odriozola. Effect of broodstock diets on lipid classes and their fatty acid composition in eggs of gilthead sea bream (Sparus aurata L.). Fish Physiology and Biochemistry, 8(2):93–101, March 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004436.

Mourente:1990:EBDb

[12] G. Mourente and J. M. Odriozola. Effect of broodstock diets on total lipids and fatty acid composition of larvae of gilthead sea bream (Sparus aurata L.) during yolksac stage. Fish Physiology and Biochemistry, 8 (2):103–110, March 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004437.

Tyler:1990:PPC

[13] Charles R. Tyler and John P. Sumpter. The purification and partial characterization of carp, Cyprinus carpio, vitellogenin. Fish Physiology and Biochemistry, 8(2):111–120, March 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004438.

Korsgaard:1990:ETI

[14] Bodil Korsgaard. Estrogen treatment and its influence on protein synthesis and amino acid metabolism in Zoarces viviparus (L) males. Fish Physiology and Biochemistry, 8(2):121–127, March 1990. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004439.

Tyler:1990:DRC

[15] Charles R. Tyler and John P. Sumpter. The development of a radioimmunoassay for carp, Cyprinus carpio, vitellogenin. Fish Physiology and Biochemistry, 8(2):129–140, March 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004440.

Pesonen:1990:RDM

[16] Maija Pesonen, Tiiu Hansson, and Tommy Andersson. Regional distribution of microsomal xenobiotic and steroid metabolism in kidney microsomes from rainbow trout. Fish Physiology and Biochemistry, 8 (2):141–145, March 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004441.

McKenzie:1990:DEA

[17] David J. McKenzie and David J. Randall. Does Amia calva aestivate? Fish Physiology and Biochemistry, 8(2):147–158, March 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004442.

Suzuki:1990:GHS

[18] Reiko Suzuki, Mitsuyo Kishida, and Tetsuya Hirano. Growth hormone secretion during longterm incubation of the pituitary of the Japanese eel, Anguilla japonica. Fish Physiology and Biochemistry, 8 (2):159–165, March 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004443.

Leatherland:1990:CST

[19] J. F. Leatherland, R. W. Hilliard, and I. C. Potter. Changes in serum thyroxine and triiodothyronine concentrations during metamorphosis of the Southern Hemisphere lamprey Geotria australis, and the effect of propylthiouracil, triiodothyronine and environmental temperature on serum thyroid hormone concentrations of ammocoetes. Fish Physiology and Biochemistry, 8(2):167–177, March 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004444.

Yardley:1990:TSD

[20] Darrell G. Yardley. Tissue-specific distribution of amylase in the mosquitofish (Gambusia affinis holbrooki). Fish Physiology and Biochemistry, 8(2):179–183, March 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004445.

Anonymous:1990:RAS

[21] Anonymous. Referees of articles submitted to the journal in 1989. Fish Physiology and Biochemistry, 8(2):185–186, March 1990. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004446.

Price:1990:SCR

[22] Jeffrey L. Price, Catherine E. Lyons, and Ru Chih C. Huang. Seasonal cycle and regulation by temperature of antifreeze protein mRNA in a long Island population of winter flounder. Fish Physiology and Biochemistry, 8(3):187–198, May 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004457.

Saligaut:1990:DCS

[23] C. Saligaut, T. Bailhache, and P. Jego. Dynamic characteristics of serotonin and dopamine metabolism in the rainbow trout brain: a regional study using liquid chromatography with electrochemical detection. Fish Physiology and Biochemistry, 8(3):199–205, May 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/BF00004458.

Watsky:1990:IDT

[24] Mitchell A. Watsky and Samuel H. Gruber. Induction and duration of tonic immobility in the lemon shark, Negaprion brevirostris. Fish Physiology and Biochemistry, 8(3):207–210, May 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004459.

Tyler:1990:DVS

[25] C. R. Tyler, J. P. Sumpter, and R. M. Handford. The dynamics of vitellogenin sequestration into vitellogenic ovarian follicles of the rainbow trout, Salmo gairdneri. Fish Physiology and Biochemistry, 8 (3):211–219, May 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004460.

Chew:1990:CME

[26] S. F. Chew, A. L. L. Lim, and Y. K. Ip. Can the mudskipper, Periophthalmus chrysospilos, tolerate acute environmental hypoxic exposure? Fish Physiology and Biochemistry, 8(3):221–227, May 1990. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004461.

Dannevig:1990:EES

[27] B. H. Dannevig, G. Struksnæs, and T. Berg. Endocytosis via the scavenger- and the mannose-receptor in rainbow trout (Salmo gairdneri) pronephros is carried out by nonphagocytic cells. Fish Physiology and Biochemistry, 8(3):229–238, May 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004462.

Tocher:1990:IMPa

[28] Douglas R. Tocher. Incorporation and metabolism of (n-3) and (n-6) polyunsaturated fatty acids in phospholipid classes in cultured rainbow trout (Salmo gairdneri) cells. Fish Physiology and Biochemistry, 8(3):239–249, May 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004463.

Tocher:1990:IMPb

[29] Douglas R. Tocher and Elizabeth E. Mackinlay. Incorporation and metabolism of (n-3) and (n-6) polyunsaturated fatty acids in phospholipid classes in cultured turbot (Scophthalmus maximus) cells. Fish Physiology and Biochemistry, 8(3):251–260, May 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004464.

Olsen:1990:CLA

[30] R. E. Olsen, R. J. Henderson, and B. J. McAndrew. The conversion of linoleic acid and linolenic acid to longer chain polyunsaturated fatty acids by Tilapia (Oreochromis) nilotica in vivo. Fish Physiology and Biochemistry, 8(3):261–270, May 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004465.

Madsen:1990:EHR

[31] Steffen S. Madsen. Enhanced hypoosmoregulatory response to growth hormone after cortisol treatment in immature rainbow trout, Salmo gairdneri. Fish Physiology and Biochemistry, 8(4):271–279, July 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00003422.

Fok:1990:DTT

[32] Pearl Fok, J. G. Eales, and S. B. Brown. Determination of 3,5,3′-triiodo-L-thyronine (T_{3}) levels in tissues of rainbow trout (Salmo gairdneri) and the effect of low ambient pH and aluminum. Fish Physiology and Biochemistry, 8(4):281–290, July 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003423.

Khan:1990:DEP

[33] I. A. Khan and K. P. Joy. Differential effects of photoperiod and temperature on hypothalamic monoaminergic activity in the teleost Channa punctatus (Bloch). Fish Physiology and Biochemistry, 8(4):291–297, July 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00003424.

Foster:1990:RGP

[34] Glen D. Foster and T. W. Moon. The role of glycogen phosphorylase in the regulation of glycogenolysis by insulin and glucagon in isolated eel (Anguilla rostrata) hepatocytes. Fish Physiology and Biochemistry, 8(4):299–309, July 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003425.

Tocher:1990:PFA

[35] Douglas R. Tocher and James R. Dick. Polyunsaturated fatty acid metabolism in cultured fish cells: Incorporation and metabolism of (n-3) and (n-6) series acids by Atlantic salmon (Salmo salar) cells. Fish Physiology and Biochemistry, 8(4):311–319, July 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00003426.

Eddy:1990:CIE

[36] F. B. Eddy, N. F. Smith, and C. Grierson. Circulatory and ionoregulatory effects of atrial natriuretic peptide on rainbow trout (Salmo gairdneri Richardson) fed normal or high levels of dietary salt. Fish Physiology and Biochemistry, 8(4):321–327, July 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00003427.

Matsuyama:1990:DRS

[37] Michiya Matsuyama, Shinji Adachi, and Shuhei Matsura. Diurnal rhythm of serum steroid hormone levels in the Japanese whiting, Sillago japonica, a daily-spawning teleost. Fish Physiology and Biochemistry, 8 (4):329–338, July 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003428.

Breton:1990:SMA

[38] B. Breton, T. Mikolajczyk, and H. Vaudry. Studies on the mode of action of neuropeptide Y (NPY) on maturational gonadotropin (GtH) secretion from perifused rainbow trout pituitary glands. Fish Physiology and Biochemistry, 8(4):339–346, July 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00003429.

Baldwin:1990:IPE

[39] Linda A. Baldwin, Edward J. Calabrese, and Jae-Ho Yang. Isolation of peroxisomal enoyl-CoA hydratase in rainbow trout and immunochemical identification with the bifunctional enzyme. Fish Physiology and Biochemistry, 8(4):347–351, July 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003430.

Honma:1990:DHA

[40] Y. Honma, A. Chiba, and U. Welsch. Development of the hypophysis of the Arctic lamprey, Lampetra japonica. Fish Physiology and Biochemistry, 8(5):355–364, September 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003367.

Sower:1990:NCR

[41] Stacia A. Sower. Neuroendocrine control of reproduction in lampreys. Fish Physiology and Biochemistry, 8(5):365–374, September 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00003368.

Larsen:1990:REI

[42] Lis Olesen Larsen. The role of external and internal factors during the phase of reproduction in lampreys, with special regard to the interplay between gonadal and extragonadal (interrenal) steroids. Fish Physiology and Biochemistry, 8(5):375–388, September 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00003369.

Youson:1990:MSI

[43] John H. Youson and Richard Cheung. Morphogenesis of somatostatin- and insulin-secreting cells in the lamprey endocrine pancreas. Fish Physiology and Biochemistry, 8(5):389–397, September 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00003370.

Noorden:1990:GHC

[44] Susan Van Noorden. Gut hormones in cyclostomes. Fish Physiology and Biochemistry, 8(5):399–408, September 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00003371.

Leatherland:1990:SES

[45] J. F. Leatherland, D. J. Macey, and I. C. Potter. Seasonal and estradiol-17β-stimulated changes in thyroid function of adult Geotria australis, a southern hemisphere lamprey. Fish Physiology and Biochemistry, 8(5): 409–417, September 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003372.

Sandnes:1990:ASD

[46] K. Sandnes, T. Hansen, and R. Waagbø. Ascorbate-2-sulfate as a dietary vitamin C source for Atlantic salmon (Salmo salar): 1. Growth, bioactivity, haematology and humoral immune response. Fish Physiology and Biochemistry, 8(6):419–427, November 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00003398.

Maage:1990:ASD

[47] Amund Maage, Rune Waagbø, and Kjartan Sandnes. Ascorbate-2-sulfate as a dietary vitamin C source for Atlantic salmon (Salmo salar): 2. Effects of dietary levels and immunization on the metabolism of trace elements. Fish Physiology and Biochemistry, 8(6):429–436, November 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00003399.

George:1990:IPI

[48] Stephen G. George and Gordon Buchanan. Isolation, properties and induction of plaice liver cytosolic glutathione-s-transferases. Fish Physiology and Biochemistry, 8(6):437–449, November 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00003400.

Harris:1990:ICH

[49] L. R. Harris, M. H. Cake, and I. C. Potter. An increase in the concentration of hepatic iron during the metamorphosis of the lamprey Geotria australis is accompanied by increased superoxide dismutase activity. Fish Physiology and Biochemistry, 8(6):451–457, November 1990. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00003401.

Jensen:1990:IBR

[50] Frank B. Jensen, Niels A. Andersen, and Norbert Heisler. Interrelationships between red cell nucleoside triphosphate content, and blood pH, O_{2}-tension and haemoglobin concentration in the carp, Cyprinus carpio. Fish Physiology and Biochemistry, 8(6):459–464, November 1990. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00003402.

Guderley:1990:AMR

[51] Helga Guderley and Luc Foley. Anatomic and metabolic responses to thermal acclimation in the ninespine stickleback, Pungitius pungitius. Fish Physiology and Biochemistry, 8(6):465–473, November 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00003403.

Skorkowski:1990:MNP

[52] Edward F. Skorkowski and Kenneth B. Storey. Mitochondrial NAD(P)- dependent malic enzyme from herring testicular tissue: Purification, kinetic behaviour and regulatory properties. Fish Physiology and Biochemistry, 8(6):475–484, November 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00003404.

Loir:1990:PSF

[53] M. Loir, C. Labbé, and F. Chambeyron. Proteins of seminal fluid and spermatozoa in the trout (Oncorhynchus mykiss): Partial characterization and variations. Fish Physiology and Biochemistry, 8(6):485–495, November 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00003405.

Solar:1990:SCI

[54] I. I. Solar, E. McLean, and E. M. Donaldson. Short communication: Induced ovulation of sablefish (Anoplopoma fimbria) following oral administration of des Gly^{10}-(D-Ala^{6})LH-RH ethylamide. Fish Physiology and Biochemistry, 8(6):497–499, November 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00003406.

Rozas:1990:DHR

[55] G. Rozas, P. Rey, and M. Aldegunde. Distribution of 5-hydroxytryptamine and related compounds in various brain regions of rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 8(6):501– 506, November 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003407.

Parker:1990:STE

[56] Steven J. Parker and Jennifer L. Specker. Salinity and temperature effects on whole-animal thyroid hormone levels in larval and juvenile striped bass, Morone saxatilis. Fish Physiology and Biochemistry, 8(6): 507–514, November 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003408.

Tagawa:1990:THE

[57] Masatomo Tagawa, Masaru Tanaka, and Tetsuya Hirano. Thyroid hormones in eggs of various freshwater, marine and diadromous teleosts and their changes during egg development. Fish Physiology and Biochemistry, 8(6):515–520, November 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003409.

Burka:1990:ECC

[58] John F. Burka, Rosalind M. J. Blair, and Jane E. Hogan. Effects of calcium channel blockers on pharmacologically induced contractions of rainbow trout (Oncorhynchus mykiss) intestine. Fish Physiology and Biochemistry, 8(6):521–527, November 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003410.

Straetkvern:1990:PPP

[59] Knut O. Strætkvern, Arnt J. Raae, and Bernt T. Walther. Purification and physicochemical properties of deoxyribonuclease from pyloric caeca of Atlantic cod (Gadus morhua L.). Fish Physiology and Biochemistry, 8 (6):529–539, November 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003411.

Nordtug:1990:OPP

[60] Trond Nordtug and Ole Kristia Berg. Optical properties of the pineal window of Atlantic salmon (Salmo salar L.). Fish Physiology and Biochemistry, 8(6):541–546, November 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00003412.

Anonymous:1990:IA

[61] Anonymous. Index of authors. Fish Physiology and Biochemistry, 8(6): 547–548, November 1990. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003413.

Anonymous:1990:CVN

[62] Anonymous. Contents, vol. 8, nos. 1–6, 1990. Fish Physiology and Biochemistry, 8(6):549–551, November 1990. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00003414.

Goksoyr:1991:ICR

[63] Anders Goksøyr, Tommy Andersson, and Lars Förlin. Immunochemical cross-reactivity of β-naphthoflavone-inducible cytochrome P450 (P450IA) in liver microsomes from different fish species and rat. Fish Physiology and Biochemistry, 9(1):1–13, March 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF01987606.

Ye:1991:EWP

[64] Xuemin Ye and D. J. Randall. The effect of water pH on swimming performance in rainbow trout (Salmo gairdneri, Richardson). Fish Physiology and Biochemistry, 9(1):15–21, March 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF01987607.

Ye:1991:EAW

[65] Xuemin Ye, D. J. Randall, and Xiqin He. The effect of acid water on oxygen consumption, circulating catecholamines and blood ionic and acidbase status in rainbow trout (Salmo gairdneri, Richardson). Fish Physiology and Biochemistry, 9(1):23–30, March 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF01987608.

Yardley:1991:EDA

[66] Darrell G. Yardley and Stacey E. Wild. Effects of diet on amylase expression in the mosquitofish. Fish Physiology and Biochemistry, 9 (1):31–37, March 1991. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01987609.

Pearson:1991:SHI

[67] M. P. Pearson and E. D. Stevens. Size and hematological impact of the splenic erythrocyte reservoir in rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 9(1):39–50, March 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/BF01987610.

Berlinsky:1991:CGH

[68] David L. Berlinsky and Jennifer L. Specker. Changes in gonadal hormones during oocyte development in the striped bass, Morone saxatilis. Fish Physiology and Biochemistry, 9(1):51–62, March 1991. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01987611.

Ip:1991:OME

[69] Y. K. Ip, C. G. L. Lee, and T. J. Lam. Osmoregulation in the mudskipper, Boleophthalmus boddaerti I. Responses of branchial cation activated and anion stimulated adenosine triphosphatases to changes in salinity. Fish Physiology and Biochemistry, 9(1):63–68, March 1991. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01987612.

Lee:1991:OME

[70] C. G. L. Lee, W. P. Low, and Y. K. Ip. Osmoregulation in the mudskipper, Boleophthalmus boddaerti II. Transepithelial potential and hormonal control. Fish Physiology and Biochemistry, 9(1):69–75, March 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01987613.

Lundin:1991:EVI

[71] Kersti Lundin. Effects of vasoactive intestinal polypeptide, substance p, 5hydroxytryptamine, met-enkephalin and neurotensin on the swimbladder smooth muscle of two teleost species, Gadus morhua and Anguilla anguilla. Fish Physiology and Biochemistry, 9(1):77–82, March 1991. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01987614.

Johnson:1991:AEC

[72] T. P. Johnson, T. W. Moon, and I. A. Johnston. Actions of epinephrine on the contractility of fast and slow skeletal muscle fibres in teleosts. Fish Physiology and Biochemistry, 9(1):83–89, March 1991. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01987615.

Szabo:1991:CSA

[73] András Szabó, János Nemcsók, and Dénes Budai. Comparative study of acetylcholine synthesis in organs of freshwater teleosts. Fish Physiology and Biochemistry, 9(2):93–99, April 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF02265124.

Chang:1991:EAS

[74] Ching-Fong Chang and Shih-Jung Lin. An enkephalin analog stimulates growth of tilapia. Fish Physiology and Biochemistry, 9(2):101–106, April 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF02265125.

Gutierrez:1991:CPG

[75] J. Gutiérrez, J. Pérez, and M. Carrillo. Changes in plasma glucagon and insulin associated with fasting in sea bass (Dicentrarchus labrax). Fish Physiology and Biochemistry, 9(2):107–112, April 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02265126.

Kindberg:1991:ITO

[76] Grete M. Kindberg, Birgit H. Dannevig, and Trond Berg. Intracellular transport of ovalbumin after in vivo endocytosis in rainbow trout liver. Fish Physiology and Biochemistry, 9(2):113–121, April 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02265127.

Maneckjee:1991:DPM

[77] A. Maneckjee, D. R. Idler, and M. Weisbart. Demonstration of putative membrane and cytosol steroid receptors for 17 α, 20 β-dihydroxy-4-pregnen-3-one in brook trout Salvelinus fontinalis oocytes by photoaffinity labelling using synthetic progestin 17,21-dimethyl-19-nor-pregn-4,9-diene-3,20-dione (R5020). Fish Physiology and Biochemistry, 9(2): 123–135, April 1991. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02265128.

Maneckjee:1991:TEA

[78] A. Maneckjee, D. R. Idler, and M. Weisbart. Transport of 17 α, 20 βdihydroxy-4-pregnen-3-one in brook trout Salvelinus fontinalis ovarian follicles. Fish Physiology and Biochemistry, 9(2):137–144, April 1991. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF02265129.

Piferrer:1991:DDD

[79] Francese Piferrer and Edward M. Donaldson. Dosage-dependent differences in the effect of aromatizable and nonaromatizable androgens on the resulting phenotype of Coho salmon (Oncorhynchus kisutch). Fish Physiology and Biochemistry, 9(2):145–150, April 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF02265130.

Olsen:1991:LAC

[80] R. E. Olsen, R. J. Henderson, and E. Ringø. Lipids of Arctic charr, Salvelinus alpinus (L.) I. Dietary induced changes in lipid class and fatty acid composition. Fish Physiology and Biochemistry, 9(2):151–164, April 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF02265131.

Dabrowska:1991:MSF

[81] Henryka Dabrowska, Klaus H. Meyer-Burgdorff, and Klaus-Dietrich Gunther. Magnesium status in freshwater fish, common carp (Cyprinus carpio, L.) and the dietary protein-magnesium interaction. Fish Physiology and Biochemistry, 9(2):165–172, April 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF02265132.

Tort:1991:BVM

[82] Lluis Tort, Ferran González-Arch, and Juan Hidalgo. On the blood volume of the Mediterranean dogfish, Scyliorhinus canicula. Fish Physiology and Biochemistry, 9(2):173–177, April 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF02265133.

Kennedy:1991:ETB

[83] Christopher J. Kennedy and Patrick J. Walsh. The effects of temperature on benzo[a]pyrene metabolism and adduct formation in the gulf toadfish, Opsanus beta. Fish Physiology and Biochemistry, 9(2):179–187, April 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF02265134.

Celander:1991:CAI

[84] Malin Celander and Lars Förlin. Catalytic activity and immunochemical quantification of hepatic cytochrome P-450 in β-naphthoflavone and isosafrol treated rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 9(3):189–197, June 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF02265140.

Wiegand:1991:IYF

[85] Murray D. Wiegand, Cheryl L. Kitchen, and Joy M. Hataley. Incorporation of yolk fatty acids into body lipids of goldfish (Carassius auratus L.) larvae raised at two different temperatures. Fish Physiology and Biochemistry, 9(3):199–213, June 1991. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02265141.

Dabrowski:1991:AGD

[86] Konrad Dabrowski. Administration of gulonolactone does not evoke ascorbic acid synthesis in teleost fish. Fish Physiology and Biochemistry, 9 (3):215–221, June 1991. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02265142.

Sundby:1991:PLI

[87] Anne Sundby, K. Eliassen, and Erika M. Plisetskaya. Plasma levels of insulin, glucagon and glucagon-like peptide in salmonids of different weights. Fish Physiology and Biochemistry, 9(3):223–230, June 1991. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02265143.

Pelissero:1991:PKI

[88] C. Pelissero, F. Le Menn, and J. F. Narbonne. Plasma kinetics of ingested tritiated estradiol and the influence on estradiol plasma levels in the cultured Siberian sturgeon Acipenser baeri. Fish Physiology and Biochemistry, 9(3):231–245, June 1991. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02265144.

Mommsen:1991:HRM

[89] Thomas P. Mommsen, Eva Danulat, and Patrick J. Walsh. Hormonal regulation of metabolism in hepatocytes of the ureogenic teleost opsanus beta. Fish Physiology and Biochemistry, 9(3):247–252, June 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02265145.

Sundby:1991:PIG

[90] A. Sundby, K. A. Eliassen, and T. Asgard. Plasma insulin, glucagon, glucagon-like peptide and glucose levels in response to feeding, starvation and life long restricted feed ration in salmonids. Fish Physiology and Biochemistry, 9(3):253–259, June 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF02265146.

Oostrom:1991:ITP

[91] J. A. Van Oostrom and N. C. Bols. Influence of temperature on the proliferative response of rainbow trout gonadal fibroblasts to cortisol and RU 486. Fish Physiology and Biochemistry, 9(3):261–269, June 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF02265147.

Arnold-Reed:1991:BAA

[92] D. Arnold-Reed, N. Hazon, and R. J. Balment. Biological actions of atrial natriuretic factor in flatfish. Fish Physiology and Biochemistry, 9(3):271–277, June 1991. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02265148.

Houston:1991:EFV

[93] Arthur Houston, Gwyneth Freeman, and Doreen Korcock. Erythrocyte fractionation by velocity sedimentation and discontinuous density gradient centrifugation. Fish Physiology and Biochemistry, 9(3):279–289, June 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF02265149.

Giles:1991:SDH

[94] M. A. Giles. Strain differences in hemoglobin polymorphism, oxygen consumption, and blood oxygen equilibria in three hatchery broodstocks of Arctic charr, Salvelinus alpinus. Fish Physiology and Biochemistry, 9 (4):291–301, December 1991. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02265150.

Farrell:1991:EET

[95] A. P. Farrell, J. A. Johansen, and R. K. Suarez. Effects of exercisetraining on cardiac performance and muscle enzymes in rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 9(4):303– 312, December 1991. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02265151.

Munger:1991:EFV

[96] R. S. Munger, S. D. Reid, and Chris M. Wood. Extracellular fluid volume measurements in tissues of the rainbow trout (Oncorhynchus mykiss) in vivo and their effects on intracellular pH and ion calculations. Fish Physiology and Biochemistry, 9(4):313–323, December 1991. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02265152.

Labbe:1991:PMT

[97] C. Labbé and M. Loir. Plasma membrane of trout spermatozoa: I. Isolation and partial characterization. Fish Physiology and Biochemistry, 9 (4):325–338, December 1991. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02265153.

Goksoyr:1991:CPS

[98] Anders Goksøyr and Håvard E. Larsen. The cytochrome P450 system of Atlantic salmon (Salmo salar): I. Basal properties and induction of P450 1A1 in liver of immature and mature fish. Fish Physiology and Biochemistry, 9(4):339–349, December 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF02265154.

Gutierrez:1991:IRB

[99] Joaquim Gutiérrez, Torbjørn Åsgård, and Erika M. Plisetskaya. Insulinreceptor binding in skeletal muscle of trout. Fish Physiology and Biochemistry, 9(4):351–360, December 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF02265155.

Harmon:1991:PCH

[100] Jamie S. Harmon, Kim G. Michelsen, and Mark A. Sheridan. Purification and characterization of hepatic triacylglycerol lipase isolated from rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 9(4): 361–368, December 1991. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02265156.

Veld:1991:GTI

[101] Peter A. Van Veld, Unchu Ko, and Donna J. Westbrook. Glutathione Stransferase in intestine, liver and hepatic lesions of mummichog (Fundulus heteroclitus) from a creosote-contaminated environment. Fish Physiology and Biochemistry, 9(4):369–376, December 1991. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF02265157.

Anonymous:1991:A

[102] Anonymous. Announcement. Fish Physiology and Biochemistry, 9 (4):i–ii, December 1991. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02265158.

Danulat:1992:NWE

[103] Eva Danulat and Stephan Kempe. Nitrogenous waste excretion and accumulation of urea and ammonia in Chalcalburnus tarichi (Cyprinidae), endemic to the extremely alkaline Lake Van (Eastern Turkey). Fish Physiology and Biochemistry, 9(5-6):377–386, February 1992. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02274218.

Shearer:1992:EWB

[104] Karl D. Shearer and Torbjørn Åsgård. The effect of water-borne magnesium on the dietary magnesium requirement of the rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 9(5-6): 387–392, February 1992. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02274219.

Olsen:1992:LAC

[105] R. E. Olsen and E. Ringø. Lipids of Arctic charr, Salvelinus alpinus (L.) II. Influence of dietary fatty acids on the elongation and desaturation of linoleic and linolenic acid. Fish Physiology and Biochemistry, 9 (5-6):393–399, February 1992. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02274220.

Borlongan:1992:EFA

[106] I. G. Borlongan. The essential fatty acid requirement of milkfish (Chanos chanos Forsskål). Fish Physiology and Biochemistry, 9(5-6): 401–407, February 1992. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02274221.

Lind:1992:SEA

[107] Ylva Lind. Summertime and early autumn activity of some enzymes in the carbohydrate and fatty acid metabolism of the crucian carp. Fish Physiology and Biochemistry, 9(5-6):409–415, February 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02274222.

Filosa:1992:QCS

[108] Michael F. Filosa, Misa A. Ito, and John H. Youson. Quantitative changes in a serum protein, AS, during the life cycle of the lamprey, Petromyzon marinus L. Fish Physiology and Biochemistry, 9(5-6):417–426, February 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF02274223.

Uematsu:1992:POI

[109] K. Uematsu, M. Kitano, and N. Iijima. Presence and ontogeny of intestinal and pancreatic phospholipase A_{2}-like proteins in the Red Sea bream, Pagrus major. An immunocytochemical study. Fish Physiology and Biochemistry, 9(5-6):427–438, February 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF02274224.

Straetkvern:1992:CPD

[110] Knut O. Strætkvern, Arnt J. Raae, and Bernt T. Walther. Characterization of a pancreatic DNase from pyloric caeca of Atlantic cod (Gadus morhua L.). Fish Physiology and Biochemistry, 9(5-6):439–452, February 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF02274225.

Gjoen:1992:MLD

[111] Tor Gjøen and Trond Berg. Metabolism of low density lipoproteins in rainbow trout. Fish Physiology and Biochemistry, 9(5-6):453–461, February 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF02274226.

Rozas:1992:AVB

[112] G. Rozas, P. Rey, and M. Aldegunde. Annual variations in brain serotonin and related compounds of domesticated rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 9(5-6):463–471, February 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF02274227.

Reddy:1992:RTHa

[113] P. K. Reddy and T. J. Lam. Role of thyroid hormones in tilapia larvae (Oreochromis mossambicus): I. Effects of the hormones and an antithyroid drug on yolk absorption, growth and development. Fish Physiology and Biochemistry, 9(5-6):473–485, February 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF02274228.

Reddy:1992:RTHb

[114] P. K. Reddy, C. L. Brown, and T. J. Lam. Role of thyroid hormones in tilapia larvae (Oreochromis mossambicus): II. Changes in the hormones and 5′-monodeiodinase activity during development. Fish Physiology and Biochemistry, 9(5-6):487–496, February 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF02274229.

Kime:1992:PMO

[115] David E. Kime. Progestogen metabolism by ovaries of the roach (Rutilus rutilus L.) and the rudd (Scardinius erythrophthalmus L.). Fish Physiology and Biochemistry, 9(5-6):497–504, February 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02274230.

Haider:1992:OME

[116] S. Haider and N. V. Rao. Oocyte maturation in Clarias batrachus. III. purification and characterization of maturation-inducing steroid. Fish Physiology and Biochemistry, 9(5-6):505–512, February 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02274231.

Lewis:1992:EDT

[117] Kim M. Lewis and Stacia A. Sower. Effects of dietary testosterone on growth and sex ratio in juvenile Atlantic salmon (Salmo salar). Fish Physiology and Biochemistry, 9(5-6):513–517, February 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02274232.

Cake:1992:DEA

[118] Max H. Cake, Ian C. Potter, and Mayamin Tajbakhsh. Digestive enzyme activities and their distribution in the alimentary canal of larvae of the three extant lamprey families. Fish Physiology and Biochemistry, 10(1):1–10, May 1992. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004649.

Johnston:1992:EFF

[119] W. L. Johnston and N. T. Glanville. Effect of feeding and fasting on plasma tryptophan and tryptophan to large neutral amino acid ratio, and on brain serotonin turnover in rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 10(1):11–22, May 1992. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004650.

McMillan:1992:PST

[120] D. Nelson McMillan and Dominic F. Houlihan. Protein synthesis in trout liver is stimulated by both feeding and fasting. Fish Physiology and Biochemistry, 10(1):23–34, May 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004651.

Coloso:1992:RTM

[121] Relicardo M. Coloso, Lillian B. Tiro, and Lita V. Benitez. Requirement for tryptophan by milkfish (Chanos chanos Forsskål) juveniles. Fish Physiology and Biochemistry, 10(1):35–41, May 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004652.

George:1992:MIC

[122] Stephen George, David Burgess, and Nick Frerichs. Metallothionein induction in cultured fibroblasts and liver of a marine flatfish, the turbot, Scophthalmus maximus. Fish Physiology and Biochemistry, 10(1):43–54, May 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004653.

Pottinger:1992:BAC

[123] T. G. Pottinger, T. A. Moran, and P. A. Cranwell. The biliary accumulation of corticosteroids in rainbow trout, Oncorhynchus mykiss, during acute and chronic stress. Fish Physiology and Biochemistry, 10(1):55–66, May 1992. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004654.

Farbridge:1992:PGH

[124] K. J. Farbridge and J. F. Leatherland. Plasma growth hormone levels in fed and fasted rainbow trout (Oncorhynchus mykiss) are decreased following handling stress. Fish Physiology and Biochemistry, 10(1):67–73, May 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004655.

Mayer-Gostan:1992:EAI

[125] N. Mayer-Gostan and R. Naon. Effects of ambient ion concentrations on gill ATPases in fresh water eel, Anguilla anguilla. Fish Physiology and Biochemistry, 10(1):75–89, May 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004656.

Schwalme:1992:QCB

[126] Karl Schwalme. A quantitative comparison between diet and body fatty acid composition in wild northern pike (Esox lucius L.). Fish Physiology and Biochemistry, 10(2):91–98, August 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004520.

Bell:1992:IMP

[127] J. Gordon Bell and John R. Sargent. The incorporation and metabolism of polyunsaturated fatty acids in phospholipids of cultured cells from chum salmon (Oncorhynchus keta). Fish Physiology and Biochemistry, 10(2):99–109, August 1992. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004521.

Boujard:1992:CPH

[128] T. Boujard and J. F. Leatherland. Circadian pattern of hepatosomatic index, liver glycogen and lipid content, plasma non-esterified fatty acid, glucose, T_{3}, T_{4}, growth hormone and cortisol concentrations in Oncorhynchus mykiss held under different photoperiod regimes and fed using demandfeeders. Fish Physiology and Biochemistry, 10(2):111–122, August 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004522.

Guderley:1992:QMM

[129] Helga Guderley and Anna Gawlicka. Qualitative modification of muscle metabolic organization with thermal acclimation of rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 10(2):123– 132, August 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004523.

Focant:1992:MPM

[130] Bruno Focant, Françoise Huriaux, and Guy Goessens. Myosin, parvalbumin and myofibril expression in barbel (Barbus barbus L.) lateral white muscle during development. Fish Physiology and Biochemistry, 10 (2):133–143, August 1992. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004524.

Lega:1992:ELS

[131] Yu. V. Lega, A. G. Chernitsky, and N. M. Belkovsky. Effect of low sea water temperature on water balance in the Atlantic salmon, (Salmo salar L.). Fish Physiology and Biochemistry, 10(2):145–148, August 1992. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004525.

Wilson:1992:SPW

[132] Rod W. Wilson and Chris M. Wood. Swimming performance, whole body ions, and gill Al accumulation during acclimation to sublethal aluminium in juvenile rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 10(2):149–159, August 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004526.

Perrott:1992:DBS

[133] M. N. Perrott, C. E. Grierson, and R. J. Balment. Drinking behaviour in sea water and fresh water teleosts, the role of the renin-angiotensin system. Fish Physiology and Biochemistry, 10(2):161–168, August 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004527.

Stevens:1992:GMR

[134] E. Don Stevens. Gill morphometry of the red drum, Sciaenops ocellatus. Fish Physiology and Biochemistry, 10(2):169–176, August 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004528.

Oikawa:1992:RBM

[135] Shin Oikawa and Yasuo Itazawa. Relationship between metabolic rate in vitro and body mass in a marine teleost, porgy Pagrus major. Fish Physiology and Biochemistry, 10(3):177–182, October 1992. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004511.

Bai:1992:DRL

[136] Sungchul C. Bai and Delbert M. Gatlin III. Dietary rutin has limited synergistic effects on vitamin C nutrition of fingerling channel catfish (Ictalurus punctatus). Fish Physiology and Biochemistry, 10(3):183–188, October 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004512.

Harmon:1992:GSL

[137] Jamie S. Harmon and Mark A. Sheridan. Glucose-stimulated lipolysis in rainbow trout, Oncorhynchus mykiss, liver. Fish Physiology and Biochemistry, 10(3):189–199, October 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004513.

Ferguson:1992:GTE

[138] R. A. Ferguson and K. B. Storey. Gluconeogenesis in trout (Oncorhynchus mykiss) white muscle: purification and characterization of fructose-1,6-bisphosphatase activity in vitro. Fish Physiology and Biochemistry, 10 (3):201–212, October 1992. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004514.

Beckman:1992:AAM

[139] Bo Beckman and Tariq Mustafa. Arachidonic acid metabolism in gill homogenate and isolated gill cells from rainbow trout, Oncorhynchus mykiss: the effect of osmolality, electrolytes and prolactin. Fish Physiology and Biochemistry, 10(3):213–222, October 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004515.

Degani:1992:EID

[140] Gad Degani and M. Abraham. Effect of insulin in the diet on the growth of European eels (Anguilla anguilla L.). Fish Physiology and Biochemistry, 10(3):223–227, October 1992. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004516.

Mackett:1992:HCI

[141] D. B. Mackett, W. H. Tam, and J. N. Fryer. Histological changes in insulinimmunoreactive pancreatic β-cells, and suppression of insulin secretion and somatotrope activity in brook trout (Salvelinus fontinalis) maintained on reduced food intake or exposed to acidic environment. Fish Physiology and Biochemistry, 10(3):229–243, October 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004517.

Farbridge:1992:TCP

[142] K. J. Farbridge and J. F. Leatherland. Temporal changes in plasma thyroid hormone, growth hormone and free fatty acid concentrations, and hepatic 5′-monodeiodinase activity, lipid and protein content during chronic fasting and re-feeding in rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 10(3):245–257, October 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004518.

Hsiao:1992:FCS

[143] Shyh-Min Hsiao and Albert H. Meier. Freerunning circasemilunar spawning rhythm of Fundulus grandis and its temperature compensation. Fish Physiology and Biochemistry, 10(3):259–265, October 1992. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004519.

Marte:1992:HCA

[144] Clarissa L. Marte and T. J. Lam. Hormonal changes accompanying sexual maturation in captive milkfish (Chanos chanos Forsskål). Fish Physiology and Biochemistry, 10(4):267–275, December 1992. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004475.

Hsu:1992:ASE

[145] Sheau-Yu Hsu and Frederick Willia Goetz. Angiotensins stimulate in vitro ovulation and contraction of brook trout (Salvelinus fontinalis) follicles. Fish Physiology and Biochemistry, 10(4):277–282, December 1992. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004476.

Lahnsteiner:1992:MER

[146] F. Lahnsteiner, R. A. Patzner, and T. Weismann. Monosaccharides as energy resources during motility of spermatozoa in Leuciscus cephalus (Cyprinidae, Teleostei). Fish Physiology and Biochemistry, 10(4):283– 289, December 1992. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004477.

Larsen:1992:CPS

[147] Håvard E. Larsen, Malin Celander, and Anders Goksøyr. The cytochrome P450 system of Atlantic salmon (Salmo salar): II. Variations in hepatic catalytic activities and isozyme patterns during an annual reproductive cycle. Fish Physiology and Biochemistry, 10(4):291–301, December 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004478.

Mazur:1992:UDS

[148] Carol Nelson Mazur, D. A. Higgs, and B. E. March. Utilization of dietary starch and glucose tolerance in juvenile Chinook salmon (Oncorhynchus tshawytscha) of different strains in seawater. Fish Physiology and Biochemistry, 10(4):303–313, December 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004479.

Sandnes:1992:MDR

[149] Kjartan Sandnes, Ole Torrissen, and Rune Waagbø. The minimum dietary requirement of vitamin C in Atlantic salmon (Salmo salar) fry using Ca ascorbate-2-monophosphate as dietary source. Fish Physiology and Biochemistry, 10(4):315–319, December 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004480.

Shiau:1992:IDE

[150] Shi-Yen Shiau and Sez-Yuan Chen. The influence of dietary energy levels with and without PCB induction on the growth of tilapia, Oreochromis niloticos x O. aureus. Fish Physiology and Biochemistry, 10 (4):321–326, December 1992. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004481.

Perry:1992:IBG

[151] S. F. Perry, G. G. Goss, and J. C. Fenwick. Interrelationships between gill chloride cell morphology and calcium uptake in freshwater teleosts. Fish Physiology and Biochemistry, 10(4):327–337, December 1992. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004482.

Lipke:1992:SIM

[152] D. W. Lipke and K. R. Olson. A specific inhibitor of mammalian kallikrein, Phe-Phe-arg-chloromethyl ketone, inhibits the production of vasoactive substances from trout plasma by kallikrein and blocks endogenous kallikrein-like activity in trout gills. Fish Physiology and Biochemistry, 10 (4):339–346, December 1992. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004483.

Dey:1992:TSI

[153] Indranil Dey and Tibor Farkas. Temperature shifts induce adaptive changes in the physical state of carp (Cyprinus carpio L.) erythrocyte plasma membranes in vitro. Fish Physiology and Biochemistry, 10(4): 347–355, December 1992. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004484.

Koven:1993:EDL

[154] W. M. Koven, S. Kolkovski, and D. Sklan. The effect of dietary lecithin and lipase, as a function of age, on n-9 fatty acid incorporation in the tissue lipids of Sparus aurata larvae. Fish Physiology and Biochemistry, 10(5):357–364, March 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004502.

Lyndon:1993:CTP

[155] A. R. Lyndon, I. Davidson, and D. F. Houlihan. Changes in tissue and plasma free amino acid concentrations after feeding in Atlantic cod. Fish Physiology and Biochemistry, 10(5):365–375, March 1993. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004503.

Woo:1993:IOA

[156] N. Y. S. Woo, A. S. B. Chung, and T. B. Ng. Influence of oral administration of estradiol-17β and testosterone on growth, digestion, food conversion and metabolism in the underyearling red sea bream, Chrysophrys major. Fish Physiology and Biochemistry, 10(5):377–387, March 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004504.

Kime:1993:SOT

[157] David E. Kime, Shelley Bhattacharya, and Krzysztof Bieniarz. Steroidogenesis by ovaries and testes of the European catfish, the wels (Silurus glanis), in vitro. Fish Physiology and Biochemistry, 10(5):389–398, March 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004505.

Harmin:1993:IGH

[158] S. A. Harmin and L. W. Crim. Influence of gonadotropic hormonereleasing hormone analog (GnRH-A) on plasma sex steroid profiles and milt production in male winter flounder, Pseudopleuronectes americanus (Walbaum). Fish Physiology and Biochemistry, 10(5):399–407, March 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004506.

Tanaka:1993:PGP

[159] Hideki Tanaka, Hirohiko Kagawa, and Keiji Hirose. Purification of gonadotropins (PmGTH I and II) from red seabream (Pagrus major) and development of a homologous radioimmunoassay for PmGTH II. Fish Physiology and Biochemistry, 10(5):409–418, March 1993. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004507.

De:1993:SAA

[160] Srabani De, A. K. Dasmahapatra, and A. K. Medda. Stimulation of acetylcholinesterase activity by triiodothyronine in the brain of singi fish, Heteropneustes fossilis (Bloch). Fish Physiology and Biochemistry, 10 (5):419–424, March 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004508.

Senthilkumaran:1993:ACC

[161] B. Senthilkumaran and K. P. Joy. Annual cyclic, and castration and cyproterone acetate-induced, changes in sialic acid content of the seminal vesicle of the catfish, Heteropneustes fossilis (Bloch). Fish Physiology and Biochemistry, 10(5):425–430, March 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004509.

Whitaker:1993:CTT

[162] Alison Whitaker and J. G. Eales. Comparison of 3,5,3′-triiodo-L-thyronine and L-thyroxine absorption from the intestinal lumen of the fasted rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 10 (5):431–441, March 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004510.

Mourente:1993:IMCa

[163] Gabriel Mourente and Douglas R. Tocher. Incorporation and metabolism of ^{14}C-labelled polyunsaturated fatty acids in juvenile gilthead sea bream Sparus aurata L. in vivo. Fish Physiology and Biochemistry, 10(6): 443–453, April 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004599.

Hemre:1993:DCU

[164] Gro-Ingunn Hemre, Øyvind Lie, and Anne Sundby. Dietary carbohydrate utilization in cod (Gadus morhua): metabolic responses to feeding and fasting. Fish Physiology and Biochemistry, 10(6):455–463, April 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004600.

Gjoen:1993:ILD

[165] Tor Gjøen and Trond Berg. Interaction of low density lipoproteins with liver cells in rainbow trout. Fish Physiology and Biochemistry, 10 (6):465–473, April 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004601.

Brown:1993:HLT

[166] J. Anne Brown and Peter Tytler. Hypoosmoregulation of larvae of the turbot, Scophthalamus maximus: drinking and gut function in relation to environmental salinity. Fish Physiology and Biochemistry, 10 (6):475–483, April 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004602.

Salama:1993:RCR

[167] Annika Salama. The role of cAMP in regulating the β-adrenergic response of rainbow trout (Oncorhynchus mykiss) red blood cells. Fish Physiology and Biochemistry, 10(6):485–490, April 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004603.

Hanssen:1993:SKF

[168] R. G. J. M. Hanssen, N. Mayer-Gostan, and S. E. Wendelaar Bonga. Stanniocalcin kinetics in freshwater and seawater European eel (Anguilla anguilla). Fish Physiology and Biochemistry, 10(6):491–496, April 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004604.

Maule:1993:SCC

[169] Alec G. Maule, Carl B. Schreck, and Cameron Sharpe. Seasonal changes in cortisol sensitivity and glucocorticoid receptor affinity and number in leukocytes of Coho salmon. Fish Physiology and Biochemistry, 10 (6):497–506, April 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004605.

McKenzie:1993:ECP

[170] D. J. McKenzie, D. J. Randall, and S. Aota. Effects of changes in plasma pH, CO_{2} and ammonia on ventilation in trout. Fish Physiology and Biochemistry, 10(6):507–515, April 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004606.

Anonymous:1993:RAS

[171] Anonymous. Referees of articles submitted to the journal in 1992. Fish Physiology and Biochemistry, 10(6):517–519, April 1993. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004607.

Anonymous:1993:Ea

[172] Anonymous. Erratum. Fish Physiology and Biochemistry, 10(6):520, April 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004608.

Anonymous:1993:CVN

[173] Anonymous. Contents, vol. 10, nos. 1–6, April 1993. Fish Physiology and Biochemistry, 10(6):i–ii, April 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004609.

Anonymous:1993:IA

[174] Anonymous. Index of authors. Fish Physiology and Biochemistry, 10(6):iii–iv, April 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004610.

Perry:1993:BAS

[175] Steve F. Perry and Scott D. Reid. β-adrenergic signal transduction in fish: interactive effects of catecholamines and cortisol. Fish Physiology and Biochemistry, 11(1-6):??, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004567.

Groneveld:1993:CSA

[176] Diet Gröneveld, Mark J. Hut, and Sjoerd E. Wendelaar Bonga. Cloning and sequence analysis of hypothalamus cDNA encoding tilapia melaninconcentrating hormone. Fish Physiology and Biochemistry, 11(1-6): ??, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-

5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004557.

Anonymous:1993:P

[177] Anonymous. Preface. Fish Physiology and Biochemistry, 11(1-6):1–2, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004544.

Nagahama:1993:MEO

[178] Yoshitaka Nagahama, Michiyasu Yoshikuni, and Minoru Tanaka. Molecular endocrinology of oocyte growth and maturation in fish. Fish Physiology and Biochemistry, 11(1-6):3–14, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004545.

Yoshikuni:1993:SBH

[179] Michiyasu Yoshikuni, Naoki Shibata, and Yoshitaka Nagahama. Specific binding of [^{3}H] 17 α, 20 β-dihydroxy-4-pregnen-3-one to oocyte cortices of rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 11(1-6):15–24, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004546.

Chang:1993:IMM

[180] John P. Chang, Richard M. Jobin, and Anderson O. L. Wong. Intracellular mechanisms mediating gonadotropin and growth hormone release in the goldfish, Carassius auratus. Fish Physiology and Biochemistry, 11(1-6):25–33, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004547.

Jobin:1993:IPK

[181] Richard M. Jobin and John P. Chang. Involvement of protein kinase C in the modulation of gonadotropin and growth hormone secretion from dispersed goldfish pituitary cells. Fish Physiology and Biochemistry, 11(1-6):35–42, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004548.

Habibi:1993:EGR

[182] Hamid R. Habibi and Debananda Pati. Extrapituitary gonadotropinreleasing hormone (GnRH) binding sites in goldfish. Fish Physiology and Biochemistry, 11(1-6):43–49, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004549.

Levavi-Sivan:1993:IMG

[183] Berta Levavi-Sivan and Zvi Yaron. Intracellular mediation of GnRH action on GTH release in tilapia. Fish Physiology and Biochemistry, 11(1-6):51–59, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004550.

Xiong:1993:CSP

[184] Fei Xiong, Ronald Chin, and Choy L. Hew. Control of salmon pituitary hormone gene expression. Fish Physiology and Biochemistry, 11 (1-6):63–70, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004551.

Lin:1993:RET

[185] Xin-Wei Lin, Hao-Ren Lin, and Richard E. Peter. The regulatory effects of thyrotropin-releasing hormone on growth hormone secretion from the pituitary of common carp in vitro. Fish Physiology and Biochemistry, 11(1-6):71–76, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004552.

Wong:1993:DFG

[186] Anderson O. L. Wong, John P. Chang, and Richard E. Peter. Dopamine functions as a growth hormone-releasing factor in the goldfish, Carassius auratus. Fish Physiology and Biochemistry, 11(1-6):77–84, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004553.

Kah:1993:RBF

[187] Olivier Kah, Isabelle Anglade, and David de Monbrison. The reproductive brain in fish. Fish Physiology and Biochemistry, 11(1-6):85–98, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004554.

Dufour:1993:DDR

[188] Sylvie Dufour, Maïté Montero, and Yves-Alain Fontaine. Differential distribution and response to experimental sexual maturation of two forms of brain gonadotropin-releasing hormone (GnRH) in the European eel, Anguilla anguilla. Fish Physiology and Biochemistry, 11(1-6):99–106, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004555.

Schulz:1993:FRP

[189] R. W. Schulz, H. Paczoska-Eliasiewicz, and H. J. Th. Goos. The feedback regulation of pituitary GTH-II secretion in male African catfish (Clarias gariepinus): Participation of 11-ketotestosterone. Fish Physiology and Biochemistry, 11(1-6):107–115, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004556.

Balm:1993:NRA

[190] P. H. M. Balm, P. Pepels, and S. E. Wendelaar Bonga. Neuroimmunological regulation of α -MSH release in tilapia (Oreochromis mossambicus).

Fish Physiology and Biochemistry, 11(1-6):125–130, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004558.

Prunet:1993:GEC

[191] Patrick Prunet, Jean-François Gonnard, and Gilles Paboeuf. GABA-ergic control of prolactin release in rainbow trout (Oncorhynchus mykiss) pituitaries in vitro. Fish Physiology and Biochemistry, 11(1-6):131–137, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004559.

Saligaut:1993:CSR

[192] C. Saligaut, S. Bennani, and T. Bailhache. Catecholamine synthesis in the rainbow trout (Oncorhynchus mykiss) brain: modulation of tyrosine hydroxylase activity. Fish Physiology and Biochemistry, 11(1-6):139–144, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004560.

Fontaine:1993:AEV

[193] Y.-A. Fontaine. Adaptations versus accommodations: some neuroendocrine aspects in teleost fish. Fish Physiology and Biochemistry, 11 (1-6):147–154, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004561.

Sakamoto:1993:OAG

[194] Tatsuya Sakamoto, Stephen D. McCormick, and Tetsuya Hirano. Osmoregulatory actions of growth hormone and its mode of action in salmonids: A review. Fish Physiology and Biochemistry, 11(1-6): 155–164, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004562.

Leloup:1993:TNA

[195] Jacques Leloup and Jean-Marc Lebel. Triiodothyronine is necessary for the action of growth hormone in acclimation to seawater of brown (Salmo trutta) and rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 11(1-6):165–173, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004563.

Rand-Weaver:1993:PSL

[196] Mariann Rand-Weaver and Penny Swanson. Plasma somatolactin levels in Coho salmon (Oncorhynchus kisutch) during smoltification and sexual maturation. Fish Physiology and Biochemistry, 11(1-6):175–182, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004564.

Takei:1993:BPF

[197] Yoshio Takei and Richard J. Balment. Biochemistry and physiology of a family of eel natriuretic peptides. Fish Physiology and Biochemistry, 11(1-6):183–188, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004565.

Balment:1993:AVF

[198] R. J. Balment, J. M. Warne, and N. Hazon. Arginine vasotocin and fish osmoregulation. Fish Physiology and Biochemistry, 11(1-6):189–194, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004566.

Verbost:1993:BCU

[199] Pieter M. Verbost, Gert Flik, and Sjoerd E. Wendelaar Bonga. Branchial calcium uptake: possible mechanisms of control by stanniocalcin. Fish Physiology and Biochemistry, 11(1-6):205–215, July 1993. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004568.

Gac:1993:GHG

[200] Florence Le Gac, Odile Blaise, and Claudine Weil. Growth hormone (GH) and reproduction: a review. Fish Physiology and Biochemistry, 11(1-6):219–232, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004569.

Amano:1993:CLG

[201] Masafumi Amano, Katsumi Aida, and Yoshihisa Hasegawa. Changes in levels of GnRH in the brain and pituitary and GTH in the pituitary in male masu salmon, Oncorhynchus masou, from hatching to maturation. Fish Physiology and Biochemistry, 11(1-6):233–240, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004570.

Naito:1993:UCT

[202] Nobuko Naito, Kunimasa Suzuki, and Yasumitsu Nakai. Ultrastructural characteristics of two distinct gonadotropes (GTH I- and GTH II-cells) in the pituitary of rainbow trout Oncorhynchus mykiss. Fish Physiology and Biochemistry, 11(1-6):241–246, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004571.

Gutierrez:1993:IIR

[203] J. Gutiérrez, M. Párrizas, and J. Planas. Insulin and IGF-i receptors and tyrosine kinase activity in carp ovaries: changes with reproductive cycle. Fish Physiology and Biochemistry, 11(1-6):247–254, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004572.

Zandbergen:1993:GCP

[204] M. A. Zandbergen, C. A. van Branden, and J. Peute. GTH-cells in the pituitary of the African catfish, Clarias gariepinus, during gonadal maturation: an immuno-electron microscopical study. Fish Physiology and Biochemistry, 11(1-6):255–263, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004573.

Soyano:1993:ETH

[205] Kiyoshi Soyano, Toshihiko Saito, and Kohei Yamauchi. Effects of thyroid hormone on gonadotropin-induced steroid production in medaka, Oryzias latipes, ovarian follicles. Fish Physiology and Biochemistry, 11 (1-6):265–272, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004574.

Sakai:1993:IER

[206] Noriyoshi Sakai, Minoru Tanaka, and Yoshitaka Nagahama. Isolation and expression of rainbow trout (Oncorhynchus mykiss) ovarian cDNA encoding 3β-hydroxysteroid dehydrogenase/Δ^{5−4}-isomerase. Fish Physiology and Biochemistry, 11(1-6):273–279, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004575.

Latz:1993:AMS

[207] M. Latz and R. Reinboth. Androgen metabolism in the skin of the rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 11(1-6):281–286, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004576.

Kime:1993:EVS

[208] D. E. Kime and A. P. Scott. In vitro synthesis of 20α-reduced and of 11- and 21-oxygenated steroids and their sulfates by testes of the goldfish (Carassius auratus): Testicular synthesis of corticosteroids. Fish Physiology and Biochemistry, 11(1-6):287–292, July 1993. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004577.

Piferrer:1993:SIR

[209] Francesc Piferrer, Michael Redding, and Gloria Callard. Stimulatory and inhibitory regulation of DNA synthesis during spermatogenesis: studies in Squalus acanthias. Fish Physiology and Biochemistry, 11 (1-6):293–298, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004578.

Sourdaine:1993:CCI

[210] Pascal Sourdaine and Bernard Jégou. Cell-cell interactions in the testis of the dogfish: stage-related changes in protein synthesis. Fish Physiology and Biochemistry, 11(1-6):299–309, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004579.

Eales:1993:THD

[211] J. G. Eales, D. L. MacLatchy, and R. M. Sweeting. Thyroid hormone deiodinase systems in salmonids, and their involvement in the regulation of thyroidal status. Fish Physiology and Biochemistry, 11(1-6):313–321, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004580.

deJesus:1993:FMR

[212] Evelyn Grace de Jesus, Tetsuya Hirano, and Yasao Inui. Flounder metamorphosis: its regulation by various hormones. Fish Physiology and Biochemistry, 11(1-6):323–328, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004581.

MacKenzie:1993:DET

[213] Duncan S. MacKenzie, Hae Young Moon, and Lisa R. Perez. Dietary effects on thyroid hormones in the red drum, Sciaenops ocellatus. Fish Physiology and Biochemistry, 11(1-6):329–335, July 1993. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004582.

Boujard:1993:ERA

[214] T. Boujard, S. Brett, and J. F. Leatherland. Effect of restricted access to demand-feeders on diurnal pattern of liver composition, plasma metabolites and hormone levels in Oncorhynchus mykiss. Fish Physiology and Biochemistry, 11(1-6):337–344, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004583.

Cavari:1993:AFG

[215] Benzion Cavari, Yunhan Hong, and Manfred Schartl. All-fish gene constructs for growth hormone gene transfer in fish. Fish Physiology and Biochemistry, 11(1-6):345–352, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004584.

Fine:1993:PCB

[216] Mira Fine, Edna Sakal, and Arieh Gertler. Preparation and comparison of biological properties of recombinant carp (Cyprinus carpio) growth hormone and its Cys-123 to Ala mutant. Fish Physiology and Biochemistry, 11(1-6):353–361, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004585.

McLean:1993:GEF

[217] Ewen McLean, Edward M. Donaldson, and Lawrence M. Souza. Growth enhancement following dietary delivery of recombinant porcine somatotropin to diploid and triploid Coho salmon (Oncorhynchus kisutch). Fish Physiology and Biochemistry, 11(1-6):363–369, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004586.

Duan:1993:ILG

[218] Cumming Duan, Stephen J. Duguay, and Erika M. Plisetskaya. Insulin-like growth factor I (IGF-I) mRNA expression in Coho salmon, Oncorhynchus kisutch: Tissue distribution and effects of growth hormone/prolactin family proteins. Fish Physiology and Biochemistry, 11(1-6):371–379, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004587.

Niu:1993:DPB

[219] Ping-De Niu, Jaime Perez-Sanchez, and Pierre-Yves Le Bail. Development of a protein binding assay for teleost insulin-like growth factor (IGF)-like: relationships between growth hormone (GH) and IGF-like in the blood of rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 11(1-6):381–391, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004588.

Marchant:1993:HIE

[220] Tracy A. Marchant and Brenda M. Moroz. Hormonal influences on in vitro [^{35} s]-sulfate uptake by gill arches from the goldfish (Carassius auratus L.). Fish Physiology and Biochemistry, 11(1-6):393–399, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004589.

Plisetskaya:1993:IBI

[221] Erika M. Plisetskaya, Elena Fabbri, and Celestina Ottolenghi. Insulin binding to isolated hepatocytes of Atlantic salmon and rainbow trout. Fish Physiology and Biochemistry, 11(1-6):401–409, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004590.

Segner:1993:BBI

[222] Helmut Segner, Ralf Böhm, and Werner Kloas. Binding and bioactivity of insulin in primary cultures of carp (Cyprinus carpio) hepatocytes. Fish Physiology and Biochemistry, 11(1-6):411–420, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004591.

Cowley:1993:ISH

[223] Darrin J. Cowley and Mark A. Sheridan. Insulin stimulates hepatic lipogenesis in rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 11(1-6):421–428, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004592.

Mommsen:1993:MEF

[224] Thomas P. Mommsen and Erika M. Plisetskaya. Metabolic and endocrine functions of glucagon-like peptides — evolutionary and biochemical perspectives. Fish Physiology and Biochemistry, 11(1-6):429–438, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004593.

Yamada:1993:STE

[225] Hideaki Yamada, Hiromi Ohta, and Kohei Yamauchi. Serum thyroxine, estradiol-17β, and testosterone profiles during the parr-smolt transformation of masu salmon, Oncorhynchus masou. Fish Physiology and Biochemistry, 12(1):1–9, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004317.

Schmitz:1993:EAS

[226] Monika Schmitz and Ian Mayer. Effect of androgens on seawater adaptation in Arctic char, Salvelinus alpinus. Fish Physiology and Biochemistry, 12(1):11–20, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004318.

Vermeulen:1993:SHS

[227] G. J. Vermeulen, J. G. D. Lambert, and H. J. Th. Goos. Steroid hormone secretion by testicular tissue from African catfish, Clarias gariepinus, in primary culture: identification and quantification by gas chromatography — mass spectrometry. Fish Physiology and Biochemistry, 12(1):21–30, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004319.

Tao:1993:PCI

[228] Yunxia Tao, Akihiko Hara, and Craig V. Sullivan. Purification, characterization and immunoassay of striped bass (Morone saxatilis) vitellogenin. Fish Physiology and Biochemistry, 12(1):31–46, July 1993. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004320.

Kalinin:1993:DBS

[229] Ana Lúcia Kalinin, Francisco Tadeu Rantin, and Mogens Lesner Glass. Dependence on body size of respiratory function in Hoplias malabaricus (Teleostei, Erythrinidae) during graded hypoxia. Fish Physiology and Biochemistry, 12(1):47–51, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004321.

Burka:1993:EAT

[230] John F. Burka, Heather A. Briand, and William P. Ireland. The effects of acute temperature change on smooth muscle contractility of rainbow trout (Oncorhynchus mykiss Walbaum) intestine. Fish Physiology and Biochemistry, 12(1):53–60, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004322.

Waagbo:1993:DVC

[231] Rune Waagbø, Johan Glette, and Kjartan Sandnes. Dietary vitamin C, immunity and disease resistance in Atlantic salmon (Salmo salar). Fish Physiology and Biochemistry, 12(1):61–73, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/BF00004323.

Luberda:1993:SPP

[232] Zofia Luberda, Jerzy Strzezek, and Miroslaw Luczynski. Some proteolytic properties of hatching liquid in the sea trout, Salmo trutta m. trutta L. Fish Physiology and Biochemistry, 12(1):75–80, July 1993. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004324.

Anonymous:1993:Eb

[233] Anonymous. Errata. Fish Physiology and Biochemistry, 12(1):i, July 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004325.

Pelletier:1993:DAC

[234] Dany Pelletier, Helga Guderley, and Jean-Denis Dutil. Does the aerobic capacity of fish muscle change with growth rates? Fish Physiology and Biochemistry, 12(2):83–93, August 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004373.

Davie:1993:BPH

[235] Peter S. Davie, Craig E. Franklin, and Gordon C. Grigg. Blood pressure and heart rate during tonic immobility in the black tipped reef shark, Carcharhinus melanoptera. Fish Physiology and Biochemistry, 12 (2):95–100, August 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004374.

Borch:1993:CAV

[236] Kristian Borch, Frank B. Jensen, and Bent B. Andersen. Cardiac activity, ventilation rate and acid-base regulation in rainbow trout exposed to hypoxia and combined hypoxia and hypercapnia. Fish Physiology and Biochemistry, 12(2):101–110, August 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004375.

Jensen:1993:INT

[237] Frank B. Jensen. Influence of nucleoside triphosphates, inorganic salts, NADH, catecholamines, and oxygen saturation on nitrite-induced oxidation of rainbow trout haemoglobin. Fish Physiology and Biochemistry, 12(2):111–117, August 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004376.

Mourente:1993:IMCb

[238] Gabriel Mourente and Douglas R. Tocher. Incorporation and metabolism of ^{14}C-labelled polyunsaturated fatty acids in wild-caught juveniles of golden grey mullet, Liza aurata, in vivo. Fish Physiology and Biochemistry, 12(2):119–130, August 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004377.

Martin:1993:PMD

[239] N. B. Martin, D. F. Houlihan, and R. M. Palmer. Protein metabolism during sexual maturation in female Atlantic salmon (Salmo salar L.). Fish Physiology and Biochemistry, 12(2):131–141, August 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004378.

Pfeiler:1993:CDU

[240] Edward Pfeiler. Characterization and distribution of undersulfated chondroitin sulfate and chondroitin in leptocephalous larvae of elopomorph fishes. Fish Physiology and Biochemistry, 12(2):143–148, August 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004379.

Leatherland:1993:IDP

[241] J. F. Leatherland and S. B. Barrett. Investigations into the development of the pituitary gland-thyroid tissue axis and distribution of tissue thyroid hormone content in embryonic Coho salmon (Oncorhynchus kisutch) from Lake Ontario. Fish Physiology and Biochemistry, 12(2):149–159, August 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004380.

Cuil:1993:HRT

[242] Zongbin Cuil and Zuoyan Zhu. Hormonal replacement therapy in fish:human growth hormone gene function in hypophysectomized carp. Fish Physiology and Biochemistry, 12(2):161–169, August 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004381.

Kishida:1993:VTE

[243] Mitsuyo Kishida and Jennifer L. Specker. Vitellogenin in tilapia (Oreochromis mossambicus): Induction of two forms by estradiol, quantification in plasma and characterization in oocyte extract. Fish Physiology and Biochemistry, 12(3):171–182, October 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004365.

Currie:1993:ACD

[244] S. Currie and B. L. Tufts. An analysis of carbon dioxide transport in arterial and venous blood of the rainbow trout, Oncorhynchus mykiss, following exhaustive exercise. Fish Physiology and Biochemistry, 12(3): 183–192, October 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004366.

Veillette:1993:CIF

[245] Philip A. Veillette, Ronald J. White, and Jennifer L. Specker. Changes in intestinal fluid transport in Atlantic salmon (Salmo salar L.) during parr-smolt transformation. Fish Physiology and Biochemistry, 12 (3):193–202, October 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004367.

Kolkovski:1993:EDE

[246] S. Kolkovski, A. Tandler, and A. Gertler. The effect of dietary exogenous digestive enzymes on ingestion, assimilation, growth and survival of gilthead seabream (Sparus aurata, Sparidae, Linnaeus) larvae. Fish Physiology and Biochemistry, 12(3):203–209, October 1993. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004368.

Foster:1993:CMH

[247] Glen D. Foster, J. Zhang, and T. W. Moon. Carbohydrate metabolism and hepatic zonation in the Atlantic hagfish, Myxine glutinosa liver: effects of hormones. Fish Physiology and Biochemistry, 12(3):211–219, October 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004369.

Lochmann:1993:EFA

[248] Rebecca T. Lochmann and Delbert M. Gatlin III. Essential fatty acid requirement of juvenile red drum (Sciaenops ocellatus). Fish Physiology and Biochemistry, 12(3):221–235, October 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004370.

Jensen:1993:ESP

[249] Jörgen Jensen, Paul Karila, and Susanne Holmgren. Effects of substance P and distribution of substance P-like immunoreactivity in nerves supplying the stomach of the cod, Gadus morhua. Fish Physiology and Biochemistry, 12(3):237–247, October 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004371.

Lie:1993:IDF

[250] Øyvind Lie, Aslaug Sandvin, and Rune Waagbø. Influence of dietary fatty acids on the lipid composition of lipoproteins in farmed Atlantic salmon (Salmo salar). Fish Physiology and Biochemistry, 12(3):249–260, October 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004372.

Diaz:1993:GSA

[251] Raúl E. Diaz and Edward Pfeiler. Glycosidase and sulfatase activities and their possible role in keratan sulfate degradation in metamorphosing bonefish (Albula sp.) leptocephali. Fish Physiology and Biochemistry, 12 (4):261–268, December 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004411.

Iwama:1993:SAB

[252] George K. Iwama, Atsushi Ishimatsu, and Norbert Heisler. Site of acidbase relevant ion transfer in the gills of rainbow trout (Oncorhynchus mykiss) exposed to environmental hypercapnia. Fish Physiology and Biochemistry, 12(4):269–280, December 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004412.

Arnesen:1993:FIG

[253] Arne M. Arnesen, Even H. Jørgensen, and Malcolm Jobling. Feed intake, growth and osmoregulation in Arctic charr, Salvelinus alpinus (L.), transferred from freshwater to saltwater at 8^{o}C during summer and winter. Fish Physiology and Biochemistry, 12(4):281–292, December 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004413.

Ventrella:1993:LCM

[254] V. Ventrella, A. Pagliarani, and A. R. Borgatti. Lipid composition and microsomal ATPase activities in gills and kidneys of warm- and coldacclimated sea bass (Dicentrarchus labrax L.). Fish Physiology and Biochemistry, 12(4):293–304, December 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004414.

Carter:1993:PNF

[255] C. G. Carter, D. F. Houlihan, and A. I. Mitchell. Protein-nitrogen flux and protein growth efficiency of individual Atlantic salmon (Salmo salar L.). Fish Physiology and Biochemistry, 12(4):305–315, December 1993. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004415.

Fynn-Aikins:1993:EFH

[256] Kofi Fynn-Aikins, Silas S. O. Hung, and Steven G. Hughes. Effects of feeding a high level of D-glucose on liver function in juvenile white sturgeon (Acipenser transmontanus). Fish Physiology and Biochemistry, 12 (4):317–325, December 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004416.

Vijayan:1993:ECH

[257] M. M. Vijayan, G. D. Foster, and T. W. Moon. Effects of cortisol on hepatic carbohydrate metabolism and responsiveness to hormones in the sea raven, Hemitripterus americanus. Fish Physiology and Biochemistry, 12 (4):327–335, December 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004417.

Okuzawa:1993:CII

[258] Koichi Okuzawa, Masafumi Amano, and Hirohiko Kagawa. Chromatographic and immunological identification of gonadotropin-releasing hormone in five marine teleosts. Fish Physiology and Biochemistry, 12(4):

337–345, December 1993. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004418.

Burka:1993:CSM

[259] John F. Burka, Heather A. Briand, and William P. Ireland. Changes in smooth muscle contractility of rainbow trout (Oncorhynchus mykiss Walbaum) intestine during acclimation to altered temperature. Fish Physiology and Biochemistry, 12(4):347–355, December 1993. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004419.

Ciereszko:1994:RBB

[260] Andrzej Ciereszko and Konrad Dabrowski. Relationship between biochemical constituents of fish semen and fertility: the effect of short-term storage. Fish Physiology and Biochemistry, 12(5):357–367, January 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004300.

Moon:1994:EDT

[261] HaeYoung Lee Moon, Duncan S. MacKenzie, and Delbert M. Gatlin III. Effects of dietary thyroid hormones on the red drum (Sciaenops ocellatus). Fish Physiology and Biochemistry, 12(5):369–380, January 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004301.

Dimberg:1994:CAI

[262] Kenth Dimberg. The carbonic anhydrase inhibitor in trout plasma: purification and its effect on carbonic anhydrase activity and the root effect. Fish Physiology and Biochemistry, 12(5):381–386, January 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004302.

Lumsden:1994:IPC

[263] J. S. Lumsden and H. W. Ferguson. Isolation and partial characterization of rainbow trout (Oncorhynchus mykiss) gill mucin. Fish Physiology and Biochemistry, 12(5):387–398, January 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004303.

Infante:1994:DRD

[264] J. L. Zambonino Infante and C. Cahu. Development and response to a diet change of some digestive enzymes in sea bass (Dicentrarchus labrax) larvae. Fish Physiology and Biochemistry, 12(5):399–408, January 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004304.

Yang:1994:EDN

[265] X. Yang, J. L. Tabachek, and T. A. Dick. Effects of dietary n − 3 polyunsaturated fatty acids on lipid and fatty acid composition and haematology of juvenile Arctic charr Salvelinus alpinus (L.). Fish Physiology and Biochemistry, 12(5):409–420, January 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004305.

Hamre:1994:DCH

[266] Kristin Hamre, Brit Hjeltnes, and Øyvind Lie. Decreased concentration of hemoglobin, accumulation of lipid oxidation products and unchanged skeletal muscle in Atlantic salmon (Salmo salar) fed low dietary vitamin E. Fish Physiology and Biochemistry, 12(5):421–429, January 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004306.

Okimoto:1994:PVM

[267] Darren K. Okimoto, Joseph J. DiStefano III, and E. Gordon Grau. On plasma volume measurement and the effect of experimental stress in the themale tilapia, Oreochromis mossambicus, maintained in fresh water. Fish Physiology and Biochemistry, 12(5):431–438, January 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004307.

Mazik:1994:EDA

[268] Patricia M. Mazik, Steven M. Plakas, and Guy R. Stehly. Effects of dorsal aorta cannulation on the stress response of channel catfish (Ictalurus punctatus). Fish Physiology and Biochemistry, 12(5):439–444, January 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004308.

Wood:1994:CPH

[269] Chris M. Wood and Heather Simmons. The conversion of plasma HCO_{3}^{−} to CO_{2} by rainbow trout red blood cells in vitro: adrenergic inhibition and the influence of oxygenation status. Fish Physiology and Biochemistry, 12(6):445–454, March 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004447.

Reid:1994:QPH

[270] Scott D. Reid and Steve F. Perry. Quantification of presumptive Na^{+} /H^{+} antiporters of the erythrocytes of trout and eel. Fish Physiology and Biochemistry, 12(6):455–463, March 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004448.

Goss:1994:MRR

[271] Greg G. Goss, Chris M. Wood, and Steve F. Perry. Morphological responses of the rainbow trout (Oncorhynchus mykiss) gill to hyperoxia, base (NaHCO_{3}) and acid (HCl) infusions. Fish Physiology and Biochemistry, 12(6):465–477, March 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004449.

Fenwick:1994:VCE

[272] J. C. Fenwick, W. Davison, and M. E. Forster. In vivo calcitropic effect of some vitamin D compounds in the marine Antarctic teleost, Pagothenia bernacchii. Fish Physiology and Biochemistry, 12(6):479–484, March 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004450.

Johnston:1994:CAP

[273] C. E. Johnston, B. S. Horney, and R. Angus. Changes in alkaline phosphatase isoenzyme activity in tissues and plasma of Atlantic salmon (Salmo salar) before and during smoltification and gonadal maturation. Fish Physiology and Biochemistry, 12(6):485–497, March 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004451.

Pottinger:1994:SIC

[274] T. G. Pottinger, F. R. Knudsen, and J. Wilson. Stress-induced changes in the affinity and abundance of cytosolic cortisol-binding sites in the liver of rainbow trout, Oncorhynchus mykiss (Walbaum), are not accompanied by changes in measurable nuclear binding. Fish Physiology and Biochemistry, 12(6):499–511, March 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004452.

Thompson:1994:EDV

[275] I. Thompson, T. C. Fletcher, and C. J. Secombes. The effect of dietary vitamin A on the immunocompetence of Atlantic salmon (Salmo salar L.).

Fish Physiology and Biochemistry, 12(6):513–523, March 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004453.

Anonymous:1994:RASa

[276] Anonymous. Referees of articles submitted to the journal in 1993. Fish Physiology and Biochemistry, 12(6):525–527, March 1994. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004454.

Madsen:1994:PSA

[277] Steffen S. Madsen, Stephen D. McCormick, and Howard A. Bern. Physiology of seawater acclimation in the striped bass, Morone saxatilis (Walbaum). Fish Physiology and Biochemistry, 13(1):1–11, May 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004114.

Raymond:1994:SVT

[278] James A. Raymond. Seasonal variations of trimethylamine oxide and urea in the blood of a cold-adapted marine teleost, the rainbow smelt. Fish Physiology and Biochemistry, 13(1):13–22, May 1994. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004115.

Floysand:1994:ETA

[279] R. Fløysand and K. B. Helle. Effects of temperature and adrenaline on the atrial myocardium of the cultured Atlantic salmon (Salmo salar). Fish Physiology and Biochemistry, 13(1):23–30, May 1994. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004116.

Narnaware:1994:EVS

[280] Yuwaraj K. Narnaware, Bridget I. Baker, and Mike G. Tomlinson. The effect of various stresses, corticosteroids and adrenergic agents on phagocytosis in the rainbow trout Oncorhynchus mykiss. Fish Physiology and Biochemistry, 13(1):31–40, May 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004117.

Nagae:1994:ECA

[281] Masaki Nagae, Hirotoshi Fuda, and Kohei Yamauchi. The effect of cortisol administration on blood plasma immunoglobulin M (IgM) concentrations in masu salmon (Oncorhynchus masou). Fish Physiology and Biochemistry, 13(1):41–48, May 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004118.

Wade:1994:PFA

[282] Michael G. Wade, Patrice M. Jacobson, and Glen Van Der Kraak. Polyunsaturated fatty acids do not activate protein kinase C in the testis of the goldfish (Carassius auratus). Fish Physiology and Biochemistry, 13(1):49–57, May 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004119.

Foster:1994:CRL

[283] Glen D. Foster, J. Zhang, and T. W. Moon. Are cell redox or lactate dehydrogenase kinetics responsible for the absence of gluconeogenesis from lactate in sea raven, hepatocytes? Fish Physiology and Biochemistry, 13(1):59–67, May 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004120.

Koven:1994:LDTa

[284] W. M. Koven, R. J. Henderson, and J. R. Sargent. Lipid digestion in turbot (Scophthalmus maximus). I: Lipid class and fatty acid composition of digesta from different segments of the digestive tract. Fish Physiology and Biochemistry, 13(1):69–79, May 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004121.

Hylland:1994:PCM

[285] K. Hylland, C. Haux, and R. A. Andersen. Properties of cod metallothionein, its presence in different tissues and effects of Cd and Zn treatment. Fish Physiology and Biochemistry, 13(1):81–91, May 1994. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004122.

Kennedy:1994:ETU

[286] C. J. Kennedy and P. J. Walsh. The effects of temperature on the uptake and metabolism of benzo[a]pyrene in isolated gill cells of the gulf toadfish, Opsanus beta. Fish Physiology and Biochemistry, 13(2):93–103, June 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004335.

Bell:1994:EDR

[287] J. Gordon Bell, Douglas R. Tocher, and John R. Sargent. Effects of diets rich in linoleic (18:2n − 6) and α-linolenic (18:3n − 3) acids on the growth, lipid class and fatty acid compositions and eicosanoid production in juvenile turbot (Scophthalmus maximus L.). Fish Physiology and Biochemistry, 13(2):105–118, June 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004336.

Rosjo:1994:ETD

[288] Camilla Røsjø, Trond Berg, and Magny S. Thomassen. Effects of temperature and dietary n-3 and n-6 fatty acids on endocytic processes in isolated rainbow trout (Oncorhynchus mykiss, Walbaum) hepatocytes. Fish Physiology and Biochemistry, 13(2):119–132, June 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004337.

Reddy:1994:DTF

[289] P. K. Reddy and J. F. Leatherland. Does the time of feeding affect the diurnal rhythms of plasma hormone and glucose concentration and hepatic glycogen content of rainbow trout? Fish Physiology and Biochemistry, 13(2):133–140, June 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004338.

Olivereau:1994:ISC

[290] Madeleine Olivereau and Mariann Rand-Weaver. Immunoreactive somatolactin cells in the pituitary of young, migrating, spawning and spent Chinook salmon, Oncorhynchus tshawytscha. Fish Physiology and Biochemistry, 13(2):141–151, June 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004339.

Biron:1994:CGH

[291] Michel Biron and Tillmann J. Benfey. Cortisol, glucose and hematocrit changes during acute stress, cohort sampling, and the diel cycle in diploid and triploid brook trout (Salvelinus fontinalis Mitchill). Fish Physiology and Biochemistry, 13(2):153–160, June 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004340.

Winkoop:1994:SGG

[292] A. Van Winkoop, L. P. M. Timmermans, and H. J. Th. Goos. Stimulation of gonadal and germ cell development in larval and juvenile carp (Cyprinus carpio L.) by homologous pituitary extract. Fish Physiology and Biochemistry, 13(2):161–171, June 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004341.

Parwez:1994:OSF

[293] I. Parwez, Fauzia A. Sherwani, and S. V. Goswami. Osmoregulation in the stenohaline freshwater catfish, Heteropneustes fossilis (Bloch) in deionized water. Fish Physiology and Biochemistry, 13(2):173–181, June 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004342.

Cornell:1994:CSI

[294] Sean C. Cornell, David M. Portesi, and Jennifer L. Specker. Cortisol stimulates intestinal fluid uptake in Atlantic salmon (Salmo salar) in the post-smolt stage. Fish Physiology and Biochemistry, 13(3):183–190, July 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004356.

Mayer:1994:PLI

[295] Ian Mayer, Bertil Borg, and Erika M. Plisetskaya. Plasma levels of insulin and liver glycogen contents in one- and two-year old Atlantic salmon (Salmo salar L.) during the period of parr-smolt transformation. Fish Physiology and Biochemistry, 13(3):191–197, July 1994. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004357.

Perez-Sanchez:1994:SCC

[296] J. Pérez-Sánchez, H. Marti-Palanca, and P-Y. Le Bail. Seasonal changes in circulating growth hormone (GH), hepatic GH-binding and plasma insulin-like growth factor-i immunoreactivity in a marine fish, gilthead sea bream, Sparus aurata. Fish Physiology and Biochemistry, 13 (3):199–208, July 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004358.

Mol:1994:HPG

[297] K. Mol, N. Byamungu, and E. R. Kühn. Hormonal profile of growing male and female diploids and triploids of the blue tilapia, Oreochromis aureus, reared in intensive culture. Fish Physiology and Biochemistry, 13(3):209–218, July 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004359.

Piferrer:1994:UCE

[298] Francesc Piferrer and Edward M. Donaldson. Uptake and clearance of exogenous estradiol-17β and testosterone during the early development of Coho salmon (Oncorhynchus kisutch), including eggs, alevins and fry. Fish Physiology and Biochemistry, 13(3):219–232, July 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004360.

Tagawa:1994:THC

[299] M. Tagawa, T. Ogasawara, and T. Hirano. Thyroid hormone concentrations in the gonads of wild chum salmon during maturation. Fish Physiology and Biochemistry, 13(3):233–240, July 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004361.

Lie:1994:TAT

[300] Øyvind Lie, Aslaug Sandvin, and Rune Waagbø. Transport of alphatocopherol in Atlantic salmon (Salmo salar) during vitellogenesis. Fish Physiology and Biochemistry, 13(3):241–247, July 1994. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004362.

Mustafa:1994:VPT

[301] T. Mustafa and C. Agnisola. Vasoactivity of prostanoids in the trout (Oncorhynchus mykiss) coronary system: modification by noradrenaline. Fish Physiology and Biochemistry, 13(3):249–261, July 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004363.

Abel:1994:EPF

[302] Daniel C. Abel, William R. Lowelll, and Melody A. Lipke. Elasmobranch pericardial function. 3. The pericardioperitoneal canal in the horn shark Heterodontus francisci. Fish Physiology and Biochemistry, 13(3):263–274, July 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004364.

Koven:1994:LDTb

[303] W. M. Koven, R. J. Henderson, and J. R. Sargent. Lipid digestion in turbot (Scopthalmus maximus) 11: Lipolysis in vitro of ^{14}C-labelled triacylglycerol, cholesterol ester and phosphatidylcholine by digesta from different segments of the digestive tract. Fish Physiology and Biochemistry, 13(4):275–283, October 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003431.

Ogata:1994:WMM

[304] Hiroshi Ogata and Takeshi Murai. White muscle of masu salmon, Oncorhynchus masou masou, smolts possesses a strong buffering capacity due to a high level of anserine. Fish Physiology and Biochemistry, 13 (4):285–293, October 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003432.

Mayer:1994:AIG

[305] Ian Mayer, Ewen McLean, and Edward M. Donaldson. Antisomatostatininduced growth acceleration in Chinook salmon (Oncorhynchus tshawytscha). Fish Physiology and Biochemistry, 13(4):295–300, October 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https:/

/link.springer.com/article/10.1007/BF00003433.

Senthilkumaran:1994:EPA

[306] B. Senthilkumaran and K. P. Joy. Effects of photoperiod alterations on day-night variations in hypothalamic serotonin content and turnover, and monoamine oxidase activity in the female catfish, Heteropneustes fossilis (Bloch). Fish Physiology and Biochemistry, 13(4):301–307, October 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00003434.

Tyler:1994:EUO

[307] C. R. Tyler, J. J. Nagler, and M. A. Turner. Effects of unilateral ovariectomy on recruitment and growth of follicles in the rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 13(4): 309–316, October 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003435.

Kime:1994:SCA

[308] David E. Kime, Mohammad A. S. Abdullah, and Piotr Epler. Substrate concentration affects the in vitro metabolism of 17-hydroxyprogesterone by ovaries of the carp, Cyprinus carpio. Fish Physiology and Biochemistry, 13(4):317–324, October 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003436.

Hwang:1994:CBE

[309] Pung-Pung Hwang, Ya-Ni Tsai, and Yu-Chi Tung. Calcium balance in embryos and larvae of the freshwater-adapted teleost, Oreochromis mossambicus. Fish Physiology and Biochemistry, 13(4):325–333, October 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00003437.

Andersson:1994:DIC

[310] Tommy Andersson and Anders Goksøyr. Distribution and induction of cytochrome P450 1A1 in the rainbow trout brain. Fish Physiology and Biochemistry, 13(4):335–342, October 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00003438.

Suzuki:1994:IHB

[311] Tohru Suzuki, Tadahide Kurokawa, and Makoto Asashima. Identification of a heparin-binding, mesoderm-inducing peptide in the swim-bladder of the red seabream, Pagrus major: a probable fish fibroblast growth factor. Fish Physiology and Biochemistry, 13(4):343–352, October 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00003439.

Laurent:1994:GEC

[312] Pierre Laurent, Suzanne Dunel-Erb, and Jacques Lignon. Gill epithelial cells kinetics in a freshwater teleost, Oncorhynchus mykiss during adaptation to ion-poor water and hormonal treatments. Fish Physiology and Biochemistry, 13(5):353–370, November 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00003415.

Pang:1994:CSP

[313] C. S. Pang, M. A. Ali, and S. F. Pang. A comparative study of picomolar affinity 2-[^{125}I]iodomelatonin binding sites in the hearts of three salmonid species. Fish Physiology and Biochemistry, 13(5):371–378, November 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00003416.

Persson:1994:EIC

[314] Petra Persson, Kristina Sundell, and Björn Th Björnsson. Estradiol-17β-induced calcium uptake and resorption in juvenile rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 13(5):379– 386, November 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003417.

Bandyopadhyay:1994:PPT

[315] Arun Bandyopadhyay and Samir Bhattacharya. Purification of putative thyroid hormone receptor from the ovarian nuclei of fresh water perch, Anabas testudineus. Fish Physiology and Biochemistry, 13(5): 387–398, November 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003418.

Matsubara:1994:RUS

[316] Takahiro Matsubara. Role of urine in the spawning of female rose bitterling, Rhodeus ocellatus ocellatus. Fish Physiology and Biochemistry, 13 (5):399–405, November 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003419.

Lushchak:1994:IED

[317] V. I. Lushchak and K. B. Storey. Influence of exercise on the distribution of enzymes in trout white muscle and kinetic properties of AMP-deaminase from free and bound fractions. Fish Physiology and Biochemistry, 13 (5):407–418, November 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003420.

Guderley:1994:IAA

[318] Helga Guderley, Brigitte A. Lavoie, and Nicole Dubois. The interaction among age, thermal acclimation and growth rate in determining muscle metabolic capacities and tissue masses in the threespine stickleback, Gasterosteus aculeatus. Fish Physiology and Biochemistry, 13(5): 419–431, November 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003421.

Monod:1994:BEA

[319] G. Monod, D. Saucier, and L. Astic. Biotransformation enzyme activities in the olfactory organ of rainbow trout (Oncorhynchus mykiss).

Immunocytochemical localization of cytochrome P4501A1 and its induction by β-naphthoflavone. Fish Physiology and Biochemistry, 13(6): 433–444, December 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004326.

Houston:1994:VVC

[320] A. H. Houston and Julie H. Gingras-Bedard. Variable versus constant temperature acclimation regimes: Effects on hemoglobin isomorph profile in goldfish, Carassius auratus. Fish Physiology and Biochemistry, 13(6): 445–450, December 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004327.

Tocher:1994:ESG

[321] Douglas R. Tocher, John D. Castell, and John R. Sargent. Effects of salinity on the growth and lipid composition of Atlantic salmon (Salmo salar) and turbot (Scophthalmus maximus) cells in culture. Fish Physiology and Biochemistry, 13(6):451–461, December 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004328.

Arndt:1994:CRC

[322] Steven K. A. Arndt, Tillmann J. Benfey, and Richard A. Cunjak. A comparison of RNA concentrations and ornithine decarboxylase activity in Atlantic salmon (Salmo salar) muscle tissue, with respect to specific growth rates and diel variations. Fish Physiology and Biochemistry, 13 (6):463–471, December 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004329.

Malison:1994:EGC

[323] Jeffrey A. Malison, Lynne S. Procarione, and Terrence B. Kayes. Endocrine and gonadal changes during the annual reproductive cycle of the freshwater teleost, Stizostedion vitreum. Fish Physiology and Biochemistry, 13(6):473–484, December 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004330.

Mylonas:1994:THB

[324] Costadinos C. Mylonas, Craig V. Sullivan, and Jeffrey M. Hinshaw. Thyroid hormones in brown trout (Salmo trutta) reproduction and early development. Fish Physiology and Biochemistry, 13(6):485–493, December 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004331.

Anonymous:1994:RASb

[325] Anonymous. Referees of articles submitted to the journal in 1994. Fish Physiology and Biochemistry, 13(6):495–497, December 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004332.

Anonymous:1994:CVN

[326] Anonymous. Contents, vol. 13, nos. 1–6, 1994. Fish Physiology and Biochemistry, 13(6):499–501, December 1994. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004333.

Anonymous:1994:IA

[327] Anonymous. Index of authors. Fish Physiology and Biochemistry, 13 (6):503–504, December 1994. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004334.

Norberg:1995:AHE

[328] Birgitta Norberg. Atlantic halibut (Hippoglossus hippoglossus) vitellogenin: induction, isolation and partial characterization. Fish Physiology and Biochemistry, 14(1):1–13, February 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004286.

Antonopoulou:1995:EAI

[329] Efthimia Antonopoulou, Ian Mayer, and Bertil Borg. Effects of aromatase inhibitors on sexual maturation in Atlantic salmon, Salmo salar, male parr. Fish Physiology and Biochemistry, 14(1):15–24, February 1995. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004287.

Reddy:1995:ICT

[330] P. K. Reddy and J. F. Leatherland. Influence of the combination of time of feeding and ration level on the diurnal hormone rhythms in rainbow trout. Fish Physiology and Biochemistry, 14(1):25–36, February 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004288.

Balm:1995:RIF

[331] P. H. M. Balm, H. E. M. G. Haenen, and S. E. Wendelaar Bonga. Regulation of interrenal function in freshwater and sea water adapted tilapia (Oreochromis mossambicus). Fish Physiology and Biochemistry, 14(1):37–47, February 1995. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004289.

Craig:1995:EDL

[332] Steven R. Craig, William H. Neill, and Delbert M. Gatlin III. Effects of dietary lipid and environmental salinity on growth, body composition, and cold tolerance of juvenile red drum (Sciaenops ocellatus). Fish Physiology and Biochemistry, 14(1):49–61, February 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004290.

Sephton:1995:LTA

[333] D. H. Sephton and W. R. Driedzic. Low temperature acclimation decreases rates of protein synthesis in rainbow trout (Oncorhynchus mykiss) heart. Fish Physiology and Biochemistry, 14(1):63–69, February 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004291.

Jung:1995:PCF

[334] A. Jung, P. Johnson, and A. L. DeVries. Protein content and freezing avoidance properties of the subdermal extracellular matrix and serum of the Antarctic snailfish, Paraliparis devriesi. Fish Physiology and Biochemistry, 14(1):71–80, February 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004292.

Waring:1995:IRR

[335] Colin P. Waring and J. Anne Brown. Ionoregulatory and respiratory responses of brown trout, Salmo trutta, exposed to lethal and sublethal aluminium in acidic soft waters. Fish Physiology and Biochemistry, 14 (1):81–91, February 1995. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004293.

Kakizawa:1995:CPS

[336] Sho Kakizawa, Toyoji Kaneko, and Tetsuya Hirano. Changes in plasma somatolactin levels during spawning migration of chum salmon (Oncorhynchus keta). Fish Physiology and Biochemistry, 14(2):93–101, April 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00002453.

Knoph:1995:EMA

[337] M. B. Knoph. Effects of metomidate anaesthesia or transfer to pur sea water on plasma parameters in ammonia-exposed Atlantic salmon (Salmo salar L.) in sea water. Fish Physiology and Biochemistry, 14 (2):103–109, April 1995. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00002454.

Wright:1995:AUE

[338] P. A. Wright, P. Part, and C. M. Wood. Ammonia and urea excretion in the tidepool sculpin (Oligocottus maculosus): sites of excretion, effects of reduced salinity and mechanisms of urea transport. Fish Physiology and Biochemistry, 14(2):111–123, April 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00002455.

Tocher:1995:ESF

[339] Douglas R. Tocher, John D. Castell, and John R. Sargent. Effects of salinity on the fatty acid compositions of total lipid and individual glycerophospholipid classes of Atlantic salmon (Salmo salar) and turbot (Scophthalmus maximus) cells in culture. Fish Physiology and Biochemistry, 14 (2):125–137, April 1995. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00002456.

Bell:1995:EDD

[340] J. Gordon Bell, John D. Castell, and John R. Sargent. Effects of different dietary arachidonic acid: docosahexaenoic acid ratios on phospholipid fatty acid compositions and prostaglandin production in juvenile turbot (Scophthalmus maximus). Fish Physiology and Biochemistry, 14(2):139–151, April 1995. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00002457.

Carter:1995:EFT

[341] C. G. Carter, Z-Y. He, and I. Davidson. Effect of feeding on the tissue free amino acid concentrations in rainbow trout (Oncorhynchus mykiss Walbaum). Fish Physiology and Biochemistry, 14(2):153–164, April 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00002458.

Lin:1995:HEA

[342] Jui-Hsing Lin and Shi-Yen Shiau. Hepatic enzyme adaptation to different dietary carbohydrates in juvenile tilapia Oreochromis niloticus × O. aureus. Fish Physiology and Biochemistry, 14(2):165–170, April 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00002459.

Zhol:1995:SLM

[343] Shengying Zhol, Robert G. Ackman, and Carol Morrison. Storage of lipids in the myosepta of Atlantic salmon (Salmo salar). Fish Physiology and Biochemistry, 14(2):171–178, April 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00002460.

Tantikitti:1995:DPF

[344] Chutima Tantikitti and B. E. March. Dynamics of plasma free amino acids in rainbow trout (Oncorhynchus mykiss) under variety of dietary conditions. Fish Physiology and Biochemistry, 14(3):179–194, June 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004309.

Houlihan:1995:PSG

[345] D. F. Houlihan, B. H. Pedersen, and J. Brechin. Protein synthesis, growth and energetics in larval herring (Clupea harengus) at different feeding regimes. Fish Physiology and Biochemistry, 14(3):195–208, June 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004310.

Cahu:1995:EMF

[346] C. L. Cahu and J. L. Zambonino Infante. Effect of the molecular form of dietary nitrogen supply in sea bass larvae: Response of pancreatic enzymes and intestinal peptidases. Fish Physiology and Biochemistry, 14(3):209–214, June 1995. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004311.

Sabapathy:1995:SPI

[347] Uma Sabapathy and Leng-Hong Teo. Some properties of the intestinal proteases of the rabbitfish, Siganus canaliculatus (Park). Fish Physiology and Biochemistry, 14(3):215–221, June 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004312.

Henderson:1995:DEC

[348] R. James Henderson, Moira T. Park, and John R. Sargent. The desaturation and elongation of ^{14}C-labelled polyunsaturated fatty acids by pike (Esox lucius L.) in vivo. Fish Physiology and Biochemistry, 14 (3):223–235, June 1995. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004313.

Formacion:1995:OOE

[349] M. J. Formacion, B. Venkatesh, and T. J. Lam. Overripening of ovulated eggs in goldfish, Carassius auratus: II. Possible involvement of postovulatory follicles and steroids. Fish Physiology and Biochemistry, 14(3):237–246, June 1995. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004314.

Gelsleichter:1995:PVC

[350] James J. Gelsleichter, John A. Musick, and Peter Van Veld. Proteoglycans from the vertebral cartilage of the clearnose skate, Raja eglanteria: Inhibition of hydroxyapatite formation. Fish Physiology and Biochemistry, 14(3):247–251, June 1995. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004315.

Davison:1995:CRA

[351] William Davison, Michael Axelsson, and Stefan Nilsson. Cardiovascular responses to acute handling stress in the Antarctic fish Trematomus bernacchii are not mediated by circulatory catecholamines. Fish Physiology and Biochemistry, 14(3):253–257, June 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004316.

Cravedi:1995:IGH

[352] J. P. Cravedi, A. Paris, and P. Prunet. Influence of growth hormone on the hepatic mixed function oxidase and transferase systems of rainbow trout. Fish Physiology and Biochemistry, 14(4):259–266, August 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004064.

Melamed:1995:EGD

[353] Philippa Melamed, Noa Eliahu, and Zvi Yaron. The effects of gonadal development and sex steroids on growth hormone secretion in the male tilapia hybrid (Oreochromis niloticus × O. aureus). Fish Physiology and Biochemistry, 14(4):267–277, August 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004065.

Seddiki:1995:EGH

[354] H. Seddiki, V. Maxime, and C. Peyraud. Effects of growth hormone on plasma ionic regulation, respiration and extracellular acid-base status in trout (Oncorhynchus mykiss) transferred to seawater. Fish Physiology and Biochemistry, 14(4):279–288, August 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004066.

Vizziano:1995:SRH

[355] Denise Vizziano, Florence Le Gac, and Alexis Fostier. Synthesis and regulation of 17α-hydroxy-20β-dihydroprogesterone in immature males of Oncorhynchus mykiss. Fish Physiology and Biochemistry, 14(4): 289–299, August 1995. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004067.

Scott:1995:SUM

[356] A. P. Scott, Y. Nagahama, and J. J. Nagler. Sulfation and uptake of the maturation-inducing steroid, 17α, 20β-dihydroxy-4-pregnen-3-one by rainbow trout ovarian follicles. Fish Physiology and Biochemistry, 14 (4):301–311, August 1995. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004068.

Cuisset:1995:OEV

[357] B. Cuisset, A. Fostier, and F. Le Menn. Occurrence and in vitro biosynthesis of 11-ketotestosterone in Siberian sturgeon, Acipenser baeri Brandt maturing females. Fish Physiology and Biochemistry, 14(4): 313–322, August 1995. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004069.

Dimberg:1995:IPO

[358] Kenth Dimberg. Investigation of the pseudobranch organ in rainbow trout (Oncorhyncus mykiss): endogenous substrates and activities of carbonic anhydrase, lactate dehydrogenase, and 3-hydroxy-acyl CoA dehydrogenase. Fish Physiology and Biochemistry, 14(4):323–327, August 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004070.

Persson:1995:TRA

[359] Petra Persson, Yasuaki Takagi, and Björn Thrandur Björnsson. Tartrate resistant acid phosphatase as a marker for scale resorption in rainbow trout, Oncorhynchus mykiss: effects of estradiol-17β treatment and refeeding. Fish Physiology and Biochemistry, 14(4):329–339, August 1995. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004071.

McKenzie:1995:EDS

[360] D. J. McKenzie, G. Piraccini, and E. W. Taylor. Effects of diet on spontaneous locomotor activity and oxygen consumption in Adriatic sturgeon (Acipenser naccarii). Fish Physiology and Biochemistry, 14(5): 341–355, October 1995. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003373.

Gawlicka:1995:HHC

[361] A. Gawlicka, S. J. Teh, and J. de la Noüe. Histological and histochemical changes in the digestive tract of white sturgeon larvae during ontogeny. Fish Physiology and Biochemistry, 14(5):357–371, October 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00003374.

Bell:1995:EDB

[362] J. Gordon Bell, Douglas R. Tocher, and John R. Sargent. Effects of dietary borage oil [enriched in γ-linolenic acid,18:3(n-6)] or marine fish oil [enriched in eicosapentaenoic acid,20:5(n-3)] on growth, mortalities, liver histopathology and lipid composition of juvenile turbot (Scophthalmus maximus). Fish Physiology and Biochemistry, 14(5):373–383, October 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00003375.

Segner:1995:MEA

[363] Helmut Segner and Johan Verreth. Metabolic enzyme activities in larvae of the African catfish, Clarias gariepinus: changes in relation to age and nutrition. Fish Physiology and Biochemistry, 14(5):385–398, October 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00003376.

Lindner:1995:PJS

[364] P. Lindner, A. Eshell, and S. Harpaz. Proteolysis by juvenile sea bass (Dicentrarchus labrax) gastrointestinal enzymes as a method for the evaluation of feed proteins. Fish Physiology and Biochemistry, 14(5): 399–407, October 1995. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003377.

Dabrowski:1995:RBV

[365] K. Dabrowski, R. E. Ciereszko, and J. S. Ottobre. Relationship between vitamin C and plasma concentrations of testosterone in female rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 14 (5):409–414, October 1995. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00003378.

Takemura:1995:LOT

[366] Akihiro Takemura and Kazunori Takano. Lysozyme in the ovary of tilapia (Oreochromis mossambicus): its purification and some biological properties. Fish Physiology and Biochemistry, 14(5):415–421, October 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00003379.

Cutler:1995:PST

[367] Christopher P. Cutler, Ian L. Sanders, and Gordon Cramb. Primary sequence, tissue specificity and mRNA expression of the Na^{+}, K^{+}-ATPase β 1 subunit in the European eel (Anguilla anguilla). Fish Physiology and Biochemistry, 14(5):423–429, October 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00003380.

Cahu:1995:MPI

[368] C. L. Cahu and J. L. Zambonino Infante. Maturation of the pancreatic and intestinal digestive functions in sea bass (Dicentrarchus labrax): effect of weaning with different protein sources. Fish Physiology and Biochemistry, 14(6):431–437, December 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004343.

Kiessling:1995:SIS

[369] A. Kiessling, L. Larsson, and S. S. S. Hung. Spawning induces a shift in energy metabolism from glucose to lipid in rainbow trout white muscle. Fish Physiology and Biochemistry, 14(6):439–448, December 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004344.

Lin:1995:EEA

[370] Tsai-Chu Lin, Jia-Cheng Hsieh, and Cheng-I Lin. Electromechanical effects of acetylcholine on the atrial tissues of the cultured tilapia (Oreochromis nilotica × O. aureus). Fish Physiology and Biochemistry, 14 (6):449–457, December 1995. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004345.

McLaughlin:1995:TCR

[371] Robert L. McLaughlin, Moira M. Ferguson, and David L. G. Noakes. Tissue concentrations of RNA and protein for juvenile brook trout (Salvelinus fontinalis): lagged responses to fluctuations in food availability. Fish Physiology and Biochemistry, 14(6):459–469, December 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004346.

Senthilkumaran:1995:EMP

[372] B. Senthilkumaran and K. P. Joy. Effects of melatonin, p-chlorophenylalanine, and α-methylparatyrosine on plasma gonadotropin level and ovarian activity in the catfish, Heteropneustes fossilis: A study correlating changes in hypothalamic monoamines. Fish Physiology and Biochemistry, 14(6): 471–480, December 1995. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004347.

Chang:1995:EST

[373] Ching-Fong Chang, En-Lieng Lau, and Bih-Yun Lin. Estradiol-17β suppresses testicular development and stimulates sex reversal in protandrous black porgy, Acanthopagrus schlegeli. Fish Physiology and Biochemistry, 14(6):481–488, December 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004348.

King:1995:IGS

[374] William King V, David L. Berlinsky, and Craig V. Sullivan. Involvement of gonadal steroids in final oocyte maturation of white perch (Morone americana) and white bass (M. chrysops): in vivo and in vitro studies. Fish Physiology and Biochemistry, 14(6):489–500, December 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004349.

Haider:1995:PMP

[375] S. Haider and K. Balamurugan. Presence of maturation-promoting factor in 17α, 20β-dihydroxy-4-pregnen-3-one-induced oocytes of catfish, Clarias batrachus. Fish Physiology and Biochemistry, 14(6):501–508, December 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF00004350.

Byrne:1995:BCA

[376] P. J. Byrne, V. E. Ostland, and H. W. Ferguson. Blood chemistry and acidbase balance in rainbow trout Oncorhynchus mykiss with experimentallyinduced acute bacterial gill disease. Fish Physiology and Biochemistry, 14 (6):509–518, December 1995. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004351.

Reid:1995:EHE

[377] Steve G. Reid and Steve F. Perry. The effects of hypoxia, in vivo, on red blood cell β-adrenoceptors in the rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 14(6):519–524, December 1995. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004352.

Anonymous:1995:RAS

[378] Anonymous. Referees of articles submitted to the journal between October 1994 and October 1995. Fish Physiology and Biochemistry, 14(6): 525–528, December 1995. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004353.

Anonymous:1995:CVN

[379] Anonymous. Contents, vol. 14, nos. 1–6, 1995. Fish Physiology and Biochemistry, 14(6):529–531, December 1995. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF00004354.

Anonymous:1995:IA

[380] Anonymous. Index of authors. Fish Physiology and Biochemistry, 14 (6):533–534, December 1995. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004355.

Ciereszko:1996:ESD

[381] Andrzej Ciereszko, Li Liu, and Konrad Dabrowski. Effects of season and dietary ascorbic acid on some biochemical characteristics of rainbow trout (Oncorhynchus mykiss) semen. Fish Physiology and Biochemistry, 15(1):1–10, February 1996. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01874832.

Chang:1996:CVI

[382] Ching-Fong Chang, En-Lieng Lau, and Shan-Ru Jeng. Characterization of vitellogenin induced by estradiol-17β in protandrous black porgy, Acanthopagrus schlegeli. Fish Physiology and Biochemistry, 15 (1):11–19, February 1996. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01874833.

Wiegand:1996:UYF

[383] Murray D. Wiegand. Utilization of yolk fatty acids by goldfish embryos and larvae. Fish Physiology and Biochemistry, 15(1):21–27, February 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF01874834.

DCotta:1996:CBP

[384] Helena C. D’Cotta, Claudiane Gallais, and Patrick Prunet. Comparison between parr and smolt Atlantic salmon (Salmo salar) α subunit gene expression of Na^{+} /K^{+}-ATPase in gill tissue. Fish Physiology and Biochemistry, 15(1):29–39, February 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF01874835.

Linard:1996:THA

[385] Boris Linard, Sanae Bennani, and Christian Saligaut. Tyrosine hydroxylase activity and dopamine turnover of rainbow trout (Oncorhynchus mykiss) brain: the special status of the hypothalamus. Fish Physiology and Biochemistry, 15(1):41–48, February 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF01874836.

Fauconneau:1996:EGH

[386] Benoit Fauconneau, Marie Paule Mady, and Pierre Yves LeBail. Effect of growth hormone on muscle protein synthesis in rainbow trout (Oncorhynchus mykiss) and Atlantic salmon (Salmo salar). Fish Physiology and Biochemistry, 15(1):49–56, February 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF01874837.

Huang:1996:ETG

[387] Liyue Huang, Jennifer L. Specker, and David A. Bengtson. Effect of triiodothyronine on the growth and survival of larval striped bass (Morone saxatilis). Fish Physiology and Biochemistry, 15(1):57–64, February 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01874838.

Fuentes:1996:DRJ

[388] Juan Fuentes, J. C. McGeer, and F. B. Eddy. Drinking rate in juvenile Atlantic salmon, Salmo salar L fry in response to a nitric oxide donor, sodium nitroprusside and an inhibitor of angiotensin converting enzyme, enalapril. Fish Physiology and Biochemistry, 15(1):65–69, February 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01874839.

Lin:1996:QQS

[389] Jen jen Lin, Shirley MacLeod, and Ching ming Kuo. Qualitative and quantitative strategies of thermal adaptation of grass carp (Ctenopharyngodon idella) cytoplasmic malate dehydrogenases. Fish Physiology and Biochemistry, 15(1):71–81, February 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF01874840.

Gilmour:1996:EEA

[390] Kathleen M. Gilmour and Steve F. Perry. The effects of experimental anaemia on CO_{2} excretion in vitro in rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 15(1):83–94, February 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01874841.

Burka:1996:EMA

[391] John F. Burka, Heather A. Briand, and William P. Ireland. Effects of modulatory agents on neurally-mediated responses of trout intestinal smooth muscle in vitro. Fish Physiology and Biochemistry, 15(2):95–104, April 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF01875589.

Kaczanowski:1996:DEA

[392] T. C. Kaczanowski and F. W. H. Beamish. Dietary essential amino acids and heat increment in rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 15(2):105–120, April 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/BF01875590.

Moyano:1996:CDE

[393] F. J. Moyano, M. Díaz, and M. C. Sarasquete. Characterization of digestive enzyme activity during larval development of gilthead seabream (Sparus aurata). Fish Physiology and Biochemistry, 15(2):121–130, April 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF01875591.

Ng:1996:PUD

[394] Wing K. Ng, Silas S. O. Hung, and Mark A. Herold. Poor utilization of dietary free amino acids by white sturgeon. Fish Physiology and Biochemistry, 15(2):131–142, April 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF01875592.

Yokoyama:1996:IIS

[395] Masahito Yokoyama and Jun ichi Nakazoe. Intraperitoneal injection of sulfur amino acids enhance the hepatic cysteine dioxygenase activity in rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 15(2):143–148, April 1996. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875593.

Nagler:1996:GID

[396] J. J. Nagler, A. P. Scott, and J. P. Sumpter. Gonadotropins I and II do not stimulate the in vitro secretion of 17α, 20β-dihydroxy-4-pregnen-3-one 20-sulphate by rainbow trout gonads during final sexual maturation. Fish Physiology and Biochemistry, 15(2):149–156, April 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01875594.

Zhang:1996:CGH

[397] Yushi Zhang and Tracy A. Marchant. Characterization of growth hormone binding sites in the goldfish, Carassius auratus: effects of hypophysectomy and hormone injection. Fish Physiology and Biochemistry, 15 (2):157–165, April 1996. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875595.

Lahnsteiner:1996:MSE

[398] F. Lahnsteiner, B. Berger, and R. A. Patzner. Motility of spermatozoa of Alburnus alburnus (Cyprinidae) and its relationship to seminal plasma composition and sperm metabolism. Fish Physiology and Biochemistry, 15(2):167–179, April 1996. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875596.

Grizzle:1996:PFJ

[399] John M. Grizzle and Keith A. Cummins. Potassium flux in juvenile striped bass (Morone saxatilis): influence of external concentrations of sodium chloride and calcium. Fish Physiology and Biochemistry, 15 (2):181–186, April 1996. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875597.

Anonymous:1996:E

[400] Anonymous. Erratum. Fish Physiology and Biochemistry, 15(2):187, April 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF01875598.

Walker:1996:DGS

[401] S. P. Walker, D. Keast, and S. McBride. Distribution of glutamine synthetase in the snapper (Pagrus auratus) and implications for the immune system. Fish Physiology and Biochemistry, 15(3):187–194, June 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01875569.

Hazel:1996:RMS

[402] Jeffrey R. Hazel. Role of molecular species catabolism in the temperatureinduced restructuring of phosphatidylcholines in liver microsomes of thermally-acclimated rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 15(3):195–204, June 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF01875570.

Estevez:1996:CMM

[403] A. Estévez, A. Delgado, and M. J. Alejandre. Characterization of mevalonate metabolism in the sea bass (Dicentrarchus labrax L.) liver. Fish Physiology and Biochemistry, 15(3):205–211, June 1996. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01875571.

Perna:1996:GMF

[404] Sonia Aparecida Perna and Marisa Narciso Fernandes. Gill morphometry of the facultative air-breathing loricariid fish, Hypostomus plecostomus (Walbaum) with, special emphasis on aquatic respiration. Fish Physiology and Biochemistry, 15(3):213–220, June 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF01875572.

Mourente:1996:CCT

[405] Gabriel Mourente and Rosa Vázquez. Changes in the content of total lipid, lipid classes and their fatty acids of developing eggs and unfed larvae of the Senegal sole, Solea senegalensis Kaup. Fish Physiology and Biochemistry, 15(3):221–235, June 1996. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875573.

Peres:1996:ATR

[406] A. Péres, C. L. Cahu, and P. Quazuguel. Amylase and trypsin responses to intake of dietary carbohydrate and protein depend on the developmental stage in sea bass (Dicentrarchus labrax) larvae. Fish Physiology and Biochemistry, 15(3):237–242, June 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF01875574.

Schulz:1996:TRG

[407] W. Schulz, Karlien Lubberink, and Henk J. Th. Goos. Testicular responsiveness to gonadotropic hormone in vitro and Leydig and Sertoli cell ultrastructure during pubertal development of male African catfish (Clarias gariepinus). Fish Physiology and Biochemistry, 15(3):243–254, June 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF01875575.

Haider:1996:ICM

[408] S. Haider and K. Balamurugan. Identification and characterization of maturation-promoting factor from catfish, Clarias batrachus. Fish Physiology and Biochemistry, 15(3):255–263, June 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1007/BF01875576.

Trombetti:1996:RRT

[409] Fabiana Trombetti, Vittoria Ventrella, and Anna Rosa Borgatti. Response of rainbow trout gill (Na^{+} +K^{+})-ATPase and chloride cells to T_{3} and NaCl administration. Fish Physiology and Biochemistry, 15(3):265–274, June 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF01875577.

Bomgren:1996:BHR

[410] Peter Bomgren and Ann-Cathrine Jönsson. Basal H_{2}-receptor stimulated and pH-dependent acid secretion from an isolated stomach mucosa preparation of the cod, Gadus morhua, studied using a modified pH-static titration method. Fish Physiology and Biochemistry, 15(4):275–285, August 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF02112354.

Bijvelds:1996:UDE

[411] Marcel J. C. Bijvelds, Gert Flik, and Sjoerd E. Wendelaar Bonga. Uptake, distribution and excretion of magnesium in Oreochromis mossambicus: dependence on magnesium in diet and water. Fish Physiology and Biochemistry, 15(4):287–298, August 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF02112355.

Verbost:1996:EZC

[412] P. M. Verbost, M. A. Salah El-Deen, and S. E. Wendelaar Bonga. Effects of zn^{2+} on ca^{2+} uptake by mitochondria and endoplasmic reticulum in permeabilized tilapia gill cells. Fish Physiology and Biochemistry, 15 (4):299–305, August 1996. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02112356.

Tsai:1996:CPC

[413] Jong-Chang Tsai. Characterization of the pattern of cytokeratin proteins in the epidermal cells of loach, Misgurnus anguillicaudatus (Cyprinformes). Fish Physiology and Biochemistry, 15(4):307–316, August 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF02112357.

Kvingedal:1996:TMR

[414] Ane Marit Kvingedal, Anette Dehli, and Kjell-Arne Rørvik. Transferrin mRNA in relation to liver iron storage in farmed Atlantic salmon Salmo salar. Fish Physiology and Biochemistry, 15(4):317–322, August 1996.

CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF02112358.

Celander:1996:IRC

[415] Malin Celander, Donald R. Buhler, and John J. Stegeman. Immunochemical relationships of cytochrome P4503A-like proteins in teleost fish. Fish Physiology and Biochemistry, 15(4):323–332, August 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02112359.

Inaba:1996:EPA

[416] Kozo Inaba, Masaharu Seno, and Hisaya Morino. Expression patterns of acidic and basic fibroblast growth factor in loach fish embryos. Fish Physiology and Biochemistry, 15(4):333–338, August 1996. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF02112360.

Houston:1996:NHR

[417] A. H. Houston, N. Dobric, and R. Kahurananga. The nature of hematological response in fish. Fish Physiology and Biochemistry, 15(4): 339–347, August 1996. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF02112361.

Otto:1996:EAS

[418] Diana M. E. Otto and Thomas W. Moon. Endogenous antioxidant systems of two teleost fish, the rainbow trout and the black bullhead, and the effect of age. Fish Physiology and Biochemistry, 15(4):349–358, August 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF02112362.

Pfeiler:1996:EMB

[419] Edward Pfeiler. Energetics of metamorphosis in bonefish (Albula sp.) leptocephali: Role of keratan sulfate glycosaminoglycan. Fish Physiology and Biochemistry, 15(4):359–362, August 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF02112363.

Hwang:1996:EEC

[420] Pung-Pung Hwang, Yu-Chi Tung, and Min-Hwang Chang. Effect of environmental calcium levels on calcium uptake in tilapia larvae Oreochromis mossambicus. Fish Physiology and Biochemistry, 15(5):363– 370, November 1996. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875578.

Holk:1996:EIS

[421] Karin Holk. Effects of isotonic swelling on the intracellular Bohr factor and the oxygen affinity of trout and carp blood. Fish Physiology and Biochemistry, 15(5):371–375, November 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF01875579.

Stillwell:1996:HLM

[422] Evelyn J. Stillwell and Tillmann J. Benfey. Hemoglobin level, metabolic rate, opercular abduction rate and swimming efficiency in female triploid brook trout (Salvelinus fontinalis). Fish Physiology and Biochemistry, 15 (5):377–383, November 1996. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875580.

Morgan:1996:CIC

[423] John D. Morgan and George K. Iwama. Cortisol-induced changes in oxygen consumption and ionic regulation in coastal cutthroat trout (Oncorhynchus clarki clarki) parr. Fish Physiology and Biochemistry, 15(5): 385–394, November 1996. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875581.

Einarsdottir:1996:SRA

[424] Ingibörg E. Einarsdóttir and Kjell J. Nilssen. Stress responses of Atlantic salmon (Salmo salar L.) elicited by water level reduction in rearing tanks. Fish Physiology and Biochemistry, 15(5):395–400, November 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01875582.

Falcon:1996:ARC

[425] Jack Falcón, Miguel Molina-Borja, and Sol Oaknin. Age-related changes in 2-[^{125}I]-iodomelatonin binding sites in the brain of sea breams (Sparus aurata, L.). Fish Physiology and Biochemistry, 15(5):401–411, November 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF01875583.

Molina-Borja:1996:PMG

[426] Miguel Molina-Borja, Jack Falcón, and J.-P. Ravault. Production of melatonin by the gilthead sea bream pineal: an in vivo and in vitro study. Fish Physiology and Biochemistry, 15(5):413–419, November 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01875584.

McCord:1996:CME

[427] Anne McCord, Nicola Dunlop, and John A. Craft. Characterization of the microsomal epoxide hydrolase of hepatic microsomes of the common dab, Limanda limanda. Fish Physiology and Biochemistry, 15(5): 421–430, November 1996. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875585.

Moreau:1996:AAS

[428] Régis Moreau, Sadasivam J. Kaushik, and Konrad Dabrowski. Ascorbic acid status as affected by dietary treatment in the Siberian sturgeon (Acipenser baeri Brandt): tissue concentration, mobilisation and L-gulonolactone oxidase activity. Fish Physiology and Biochemistry, 15 (5):431–438, November 1996. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875586.

Einarsson:1996:EFS

[429] Sigfús Einarsson, P. Spencer Davies, and Clive Talbot. The effect of feeding on the secretion of pepsin, trypsin and chymotrypsin in the Atlantic salmon, Salmo salar L. Fish Physiology and Biochemistry, 15(5): 439–446, November 1996. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01875587.

Carneiro:1996:LSS

[430] Nancy M. Carneiro, Carmen D. Eilertson, and Mark A. Sheridan. Lipidstimulated somatostatin secretion in rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 15(5):447–452, November 1996. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01875588.

Mathiyalagan:1996:ECG

[431] A. Mathiyalagan, P. K. Reddy, and T. J. Lam. Effects of cortisol on growth and development in tilapia larvae, Oreochromis mossambicus. Fish Physiology and Biochemistry, 15(6):453–458, December 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF01874919.

Jakobsson:1996:SBK

[432] Staffan Jakobsson, Ian Mayer, and Bertil Borg. Specific binding of 11ketotestosterone in an androgen target organ, the kidney of the male three-spined stickleback, Gasterosteus aculeatus. Fish Physiology and Biochemistry, 15(6):459–467, December 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF01874920.

Ando:1996:ULP

[433] Seiichi Ando and Masao Matsuzaki. A unique lipoprotein profile found in the plasma of cultured Japanese eel Anguilla japonica: very low density lipoprotein, but not high density lipoprotein, is the main component of plasma. Fish Physiology and Biochemistry, 15(6):469–479, December 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic).

URL https://link.springer.com/article/10.1007/BF01874921.

Houston:1996:HRF

[434] A. H. Houston, W. C. Roberts, and J. A. Kennington. Hematological response in fish: pronephric and splenic involvements in the goldfish, Carassius auratus L. Fish Physiology and Biochemistry, 15(6): 481–489, December 1996. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01874922.

Soengas:1996:FDR

[435] J. L. Soengas, E. F. Strong, and M. D. Andrés. Food deprivation and refeeding in Atlantic salmon, Salmo salar: effects on brain and liver carbohydrate and ketone bodies metabolism. Fish Physiology and Biochemistry, 15(6):491–511, December 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1007/BF01874923.

Lee:1996:MRC

[436] Tsung-Han Lee, Pung-Pung Hwang, and Fore-Lien Huang. Mitochondriarich cells in the branchial epithelium of the teleost, Oreochromis mossambicus, acclimated to various hypotonic environments. Fish Physiology and Biochemistry, 15(6):513–523, December 1996. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF01874924.

Gutierrez:1996:JPM

[437] Joaquim Gutiérrez and Jesús Palomeque. José planas mestres (1926– 1995). Fish Physiology and Biochemistry, 15(6):525–526, December 1996.

CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF01874925.

Anonymous:1996:RAS

[438] Anonymous. Referees of articles submitted to the journal between October 1995 and October 1996. Fish Physiology and Biochemistry, 15(6): 527–530, December 1996. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF01874926.

Knudsen:1997:RNI

[439] P. K. Knudsen and F. B. Jensen. Recovery from nitrite-induced methaemoglobinaemia and potassium balance disturbances in carp. Fish Physiology and Biochemistry, 16(1):1–10, January 1997. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004535.

Schutz:1997:YMC

[440] M. Schütz and G. W. Barlow. Young of the Midas cichlid get biologically active nonnutrients by eating mucus from the surface of their parents. Fish Physiology and Biochemistry, 16(1):11–18, January 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004536.

Poli:1997:BPC

[441] A. Poli, B. Pavan, and R. Rossi. Biochemical and pharmacological characterization of adenosine A1 receptors in eel (Anguilla anguilla) brain. Fish Physiology and Biochemistry, 16(1):19–27, January 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004537.

Brett:1997:ENE

[442] S. E. Brett and J. F. Leatherland. Epinephrine and norepinephrine elevate 5′-monodeiodinase activity in rainbow trout liver slices. Fish Physiology and Biochemistry, 16(1):29–34, January 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1007/BF00004538.

Pavlick:1997:DIG

[443] R. J. Pavlick, Jr. and G. P. Moberg. Dopaminergic influence on gonadotropin secretion in white sturgeon (Acipenser transmontanus). Fish Physiology and Biochemistry, 16(1):35–43, January 1997. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004539.

Pottinger:1997:CPS

[444] T. G. Pottinger and A. Moore. Characterization of putative steroid receptors in the membrane, cytosol and nuclear fractions from the olfactory tissue of brown and rainbow trout. Fish Physiology and Biochemistry, 16(1):45–63, January 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004540.

Hobby:1997:RBP

[445] A. C. Hobby and N. W. Pankhurst. The relationship between plasma and ovarian levels of gonadal steroids in the repeat spawning marine fishes Pagrus auratus (Sparidae) and Chromis dispilus (Pomacentridae). Fish Physiology and Biochemistry, 16(1):65–75, January 1997. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004541.

Wang:1997:SCB

[446] Z. Wang and L. W. Crim. Seasonal changes in the biochemistry of seminal plasma and sperm motility in the ocean pout, Macrozoarces americanus. Fish Physiology and Biochemistry, 16(1):77–83, January 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004542.

Janvier:1997:MSI

[447] J.-J. Janvier. Mediation of serotonin-induced branchial vasoconstriction by a cholinergic and muscarinic response in vivo in European eel Anguilla anguilla. Fish Physiology and Biochemistry, 16(1):85–92, January 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1007/BF00004543.

Garcia-Rejon:1997:FBA

[448] L. Garcia-Rejón, M. J. Sanchez-Muros, and M. de la Higuera. Fructose 1,6 bisphosphatase activity in liver and gonads of sea bass (Dicentrarchus labrax). Influence of diet composition and stage of the reproductive cycle. Fish Physiology and Biochemistry, 16(2):93–105, March 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004667.

Peyon:1997:IPR

[449] P. Peyon, S. Baloche, and E. Burzawa-Gérard. Investigation into the possible role of androgens in the induction of hepatic vitellogenesis in the European eel: in vivo and in vitro studies. Fish Physiology and Biochemistry, 16(2):107–118, March 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004668.

Milligan:1997:RCA

[450] C. L. Milligan. The role of cortisol in amino acid mobilization and metabolism following exhaustive exercise in rainbow trout (Oncorhynchus mykiss Walbaum). Fish Physiology and Biochemistry, 16 (2):119–128, March 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004669.

Cavaco:1997:PDM

[451] J. E. B. Cavaco, J. G. D. Lambert, and H. J. Th. Goos. Pubertal development of male African catfish, Clarias gariepinus. In vitro steroidogenesis by testis and interrenal tissue and plasma levels of sexual steroids. Fish Physiology and Biochemistry, 16(2):129–138, March 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004670.

Thibault:1997:SVM

[452] M. Thibault, P. U. Blier, and H. Guderley. Seasonal variation of muscle metabolic organization in rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 16(2):139–155, March 1997. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1007/BF00004671.

Larsen:1997:IIC

[453] B. K. Larsen and F. B. Jensen. Influence of ionic composition on acid-base regulation in rainbow trout (Oncorhynchus mykiss) exposed to environmental hypercapnia. Fish Physiology and Biochemistry, 16 (2):157–170, March 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1007/BF00004672.

Chen:1997:DRT

[454] K. M. Chen, W. K. Chan, and A. D. Munro. Dexamethasone receptors and their distribution in the brain of the red tilapia. Fish Physiology and Biochemistry, 16(3):171–179, May 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007733706311.

Reis-Henriques:1997:MEV

[455] M. A. Reis-Henriques, M. M. Cruz, and J. O. Pereira. The modulating effect of vitellogenin on the synthesis of 17β-estradiol by rainbow trout (Oncorhynchus mykiss) ovary. Fish Physiology and Biochemistry, 16(3):181–186, May 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007784626784.

Chiba:1997:DDL

[456] H. Chiba, T. Miura, and K. Yamauchi. Differentiation and development of Leydig cell, and induction of spermatogenesis during testicular differentiation in the Japanese eel, Anguilla japonica. Fish Physiology and Biochemistry, 16(3):187–195, May 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007753115397.

Khan:1997:ECM

[457] M. N. Khan, P. K. Reddy, and J. F. Leatherland. Effect of cortisol on the metabolism of 17-hydroxyprogesterone by Arctic charr and rainbow trout embryos. Fish Physiology and Biochemistry, 16(3):197–209, May 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007725832235.

Mananos:1997:PMR

[458] E. L. Mañanós, S. Zanuy, and M. Carrillo. Photoperiodic manipulations of the reproductive cycle of sea bass (Dicentrarchus labrax) and their effects on gonadal development, and plasma 17ß-estradiol and vitellogenin levels. Fish Physiology and Biochemistry, 16(3):211–222, May 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007777816306.

Andersen:1997:AWI

[459] Ø. Andersen. Accumulation of waterborne iron and expression of ferritin and transferrin in early developmental stages of brown trout (Salmo trutta). Fish Physiology and Biochemistry, 16(3):223–231, May 1997. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007729900376.

Underhay:1997:EPC

[460] J. R. Underhay and J. F. Burka. Effects of pH on contractility of rainbow trout (Oncorhynchus mykiss) intestinal muscle in vitro. Fish Physiology and Biochemistry, 16(3):233–246, May 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007781917214.

Wells:1997:BOT

[461] R. M. G. Wells, J. Baldwin, and R. E. Weber. Blood oxygen transport and hemoglobin function in three tropical fish species from northern Australian freshwater billabongs. Fish Physiology and Biochemistry, 16(3):247–258, May 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007734001284.

Geurden:1997:EDP

[462] I. Geurden, P. Coutteau, and P. Sorgeloos. Effect of a dietary phospholipid supplementation on growth and fatty acid composition of European sea bass (Dicentrarchus labrax L.) and turbot (Scophthalmus maximus L.) juveniles from weaning onwards. Fish Physiology and Biochemistry, 16(4):259–272, July 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007785128042.

Luizi:1997:PLE

[463] F. Luizi, B. Korsgaard, and I. M. Petersen. Plasma lipoproteins in European eels (Anguilla anguilla): effects of estradiol. Fish Physiology and Biochemistry, 16(4):273–280, July 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007746912113.

Kao:1997:DTL

[464] Yung hsi Kao, John H. Youson, and Mark A. Sheridan. Differences in the total lipid and lipid class composition of larvae and metamorphosing sea lampreys, Petromyzon marinus. Fish Physiology and Biochemistry, 16(4):281–290, July 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007799028951.

Conceicao:1997:FGP

[465] L. E. C. Conceição, D. F. Houlihan, and J. A. J. Verreth. Fast growth, protein turnover and costs of protein metabolism in yolk-sac larvae of the African catfish (Clarias gariepinus). Fish Physiology and Biochemistry, 16(4):291–302, July 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007751130768.

Navarro:1997:EFF

[466] I. Navarro, J. Blasco, and J. Gutiérrez. Effects of fasting and feeding on plasma amino acid levels in brown trout. Fish Physiology and Biochemistry, 16(4):303–309, July 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007759316656.

Terjesen:1997:UAE

[467] B. F. Terjesen, J. Verreth, and H. J. Fyhn. Urea and ammonia excretion by embryos and larvae of the African catfish Clarias gariepinus (Burchell 1822). Fish Physiology and Biochemistry, 16(4):311–321, July 1997. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007719618473.

Bijvelds:1997:MBE

[468] M. J. C. Bijvelds, G. Flik, and S. E. Wendelaar Bonga. Mineral balance in Oreochromis mossambicus: dependence on magnesium in diet and water. Fish Physiology and Biochemistry, 16(4):323–331, July 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007727819382.

Cordiner:1997:EST

[469] S. Cordiner and S. Egginton. Effects of seasonal temperature acclimatization on muscle metabolism in rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 16(4):333–343, July 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007732003452.

Bukovskaya:1997:LVS

[470] O. Bukovskaya, J. G. D. Lambert, and D. E. Kime. In vitro steroidogenesis by gonads of the Russian sturgeon, Acipenser gueldenstaedti Brandt. Fish Physiology and Biochemistry, 16(4):345–353, July 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007784020290.

Aho:1997:SAA

[471] E. Aho and M. Vornamen. Seasonality of ATPase activities in crucian carp (Carassius carassius L.) heart. Fish Physiology and Biochemistry, 16(5):355–364, October 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007727024877.

McKenzie:1997:OCV

[472] D. J. McKenzie, G. Piraccini, and E. W. Taylor. Oxygen consumption and ventilatory reflex responses are influenced by dietary lipids in sturgeon. Fish Physiology and Biochemistry, 16(5):365–379, October 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007702908947.

Desvilettes:1997:CLC

[473] C. Desvilettes, G. Bourdier, and J. C. Breton. Changes in lipid class and fatty acid composition during development in pike (Esox lucius L.) eggs and larvae. Fish Physiology and Biochemistry, 16(5):381–393, October 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007764130651.

Ricci:1997:MHA

[474] R. Ricci, P. De Vito, and S. Incerpi. Modulation of Na/H antiport in erythrocytes of the teleost fish, Oreochromis mossambicus: effect of vasopressin. Fish Physiology and Biochemistry, 16(5):395–401, October 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007754925785.

Hwang:1997:MCU

[475] P. P. Hwang and C. H. Yang. Modulation of calcium uptake in cadmiumpretreated tilapia (Oreochromis mossambicus) larvae. Fish Physiology and Biochemistry, 16(5):403–410, October 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007707009855.

Holloway:1997:IRS

[476] A. C. Holloway and J. F. Leatherland. The influence of reproductive status on the stimulatory action of N-methyl-D,L-aspartate on growth hormone secretion, in vitro in rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 16(5):411–418, October 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007716215630.

Bornestaf:1997:EAI

[477] C. Bornestaf, E. Antonopoulou, and B. Borg. Effects of aromatase inhibitors on reproduction in male three-spined sticklebacks, Gasterosteus aculeatus, exposed to long and short photoperiods. Fish Physiology and Biochemistry, 16(5):419–423, October 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007776517447.

Ravaglia:1997:CMV

[478] M. A. Ravaglia, F. L. Lo Nostro, and G. M. Somoza. Characterization of molecular variants of GnRH, induction of spermiation and sex reversal using salmon GnRH-a and domperidone in the protogynous diandric fish, Synbranchus marmoratus Bloch, (Teleostei, Synbranchidae). Fish Physiology and Biochemistry, 16(5):425–436, October 1997. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007772416539.

Seubert:1997:TBJ

[479] J. M. Seubert and C. J. Kennedy. The toxicokinetics of benzo[a]pyrene in juvenile Coho salmon, Oncorhynchus kisutch, during smoltification. Fish Physiology and Biochemistry, 16(5):437–447, October 1997. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007704213813.

Otto:1997:REG

[480] D. M. E. Otto, C. K. Sen, and T. W. Moon. Role of exogenous glutathione in teleost fish and its effects on antioxidant defense responses in rainbow trout exposed to 3,3’,4,4’-tetrachlorobiphenyl. Fish Physiology and Biochemistry, 16(6):449–457, November 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007765609695.

Chang:1997:ECK

[481] M.-H. Chang, H.-C. Lin, and P. P. Hwang. Effects of cadmium on the kinetics of calcium uptake in developing tilapia larvae, Oreochromis mossambicus. Fish Physiology and Biochemistry, 16(6):459–470, November 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007780602426.

Aranishi:1997:EPJ

[482] F. Aranishi and M. Nakane. Epidermal proteases of the Japanese eel. Fish Physiology and Biochemistry, 16(6):471–478, November 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007736804243.

Peres:1997:DSS

[483] A. Péres, C. L. Cahu, and J. L. Zambonino Infante. Dietary spermine supplementation induces intestinal maturation in sea bass (Dicentrarchus labrax) larvae. Fish Physiology and Biochemistry, 16(6):479–485, November 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007786128254.

Iijima:1997:PCP

[484] N. Iijima, S. Chosa, and M. Kayama. Purification and characterization of phospholipase A_{2} from the pyloric caeca of red sea bream, Pagrus major. Fish Physiology and Biochemistry, 16(6):487–498, November 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007745021990.

Ghioni:1997:ECM

[485] C. Ghioni, D. R. Tocher, and J. R. Sargent. The effect of culture on morphology, lipid and fatty acid composition, and polyunsaturated fatty acid metabolism of rainbow trout (Oncorhynchus mykiss) skin cells. Fish Physiology and Biochemistry, 16(6):499–513, November 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007709508786.

Ferenczy:1997:CAM

[486] J. Ferenczy, T. Szegletes, and J. Nemcsók. Characterization of acetylcholinesterase and its molecular forms in organs of five freshwater teleosts. Fish Physiology and Biochemistry, 16(6):515–529, November 1997. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007701323808.

Guderley:1997:PPM

[487] H. Guderley, J. St. Pierre, and A. J. Hulbert. Plasticity of the properties of mitochondria from rainbow trout red muscle with seasonal acclimatization. Fish Physiology and Biochemistry, 16(6):531–541, November 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007708826437.

Kobayashi:1997:GRH

[488] M. Kobayashi, M. Amano, and K. Aida. Gonadotropin-releasing hormone and gonadotropin in goldfish and masu salmon. Fish Physiology and Biochemistry, 17(1-6):1–8, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007764430746.

Bjornsson:1997:BSG

[489] B. Th. Björnsson. The biology of salmon growth hormone: from daylight to dominance. Fish Physiology and Biochemistry, 17(1-6):9–24, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007712413908.

Sumpter:1997:ECF

[490] J. P. Sumpter. Environmental control of fish reproduction: a different perspective. Fish Physiology and Biochemistry, 17(1-6):25–31, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007782305962.

Goetz:1997:ORO

[491] F. W. Goetz and M. Garczynski. The ovarian regulation of ovulation in teleost fish. Fish Physiology and Biochemistry, 17(1-6):33–38, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007765902327.

Naito:1997:DAA

[492] N. Naito, Y. Koide, and Y. Nakai. Distinct alpha (α)-subunits of salmon glycoprotein hormones: production sites in the pituitary with sexual maturity. Fish Physiology and Biochemistry, 17(1-6):39–44, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007713918257.

Goos:1997:GRH

[493] H. J. Th. Goos, P. T. Bosma, and R. W. Schulz. Gonadotropin-releasing hormones in the African catfish: molecular forms, localization, potency and receptors. Fish Physiology and Biochemistry, 17(1-6):45–51, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007734422800.

Kah:1997:ERB

[494] O. Kah, I. Anglade, and P. Jégo. Estrogen receptors in the brain-pituitary complex and the neuroendocrine regulation of gonadotropin release in rainbow trout. Fish Physiology and Biochemistry, 17(1-6):53–62, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007749715473.

Amano:1997:MSG

[495] M. Amano, K. Ikuta, and K. Aida. The maturation of the salmon GnRH system and its regulation by gonadal steroids in masu salmon. Fish Physiology and Biochemistry, 17(1-6):63–70, December 1997. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007797631403.

Okuzawa:1997:DEG

[496] K. Okuzawa, J. Granneman, and H. Kagawa. Distinct expression of GnRH genes in the red seabream brain. Fish Physiology and Biochemistry, 17 (1-6):71–79, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007760329837.

Parhar:1997:RFM

[497] I. S. Parhar and Y. Sakuma. Regulation of forebrain and midbrain GnRH neurons in juvenile teleosts. Fish Physiology and Biochemistry, 17(1-6):81–84, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007770020983.

Rosenfeld:1997:GBS

[498] H. Rosenfeld, B. Levavi-Sivan, and A. Elizur. The GTH β subunits of tilapia: gene cloning and expression. Fish Physiology and Biochemistry, 17(1-6):85–92, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007778221892.

Melamed:1997:MLG

[499] P. Melamed, G. Gur, and Z. Yaron. The mRNA levels of GtH Iβ, GtH IIβ and GH in relation to testicular development and testosterone treatment in pituitary cells of male tilapia. Fish Physiology and Biochemistry, 17 (1-6):93–98, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007737414565.

Rebers:1997:MGH

[500] F. E. M. Rebers, C. P. Tensen, and J. Bogerd. Modulation of glycoprotein hormone α- and gonadotropin IIβ-subunit mRNA levels in the pituitary gland of mature male African catfish, Clarias gariepinus. Fish Physiology and Biochemistry, 17(1-6):99–108, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007725321061.

Thomas:1997:BCR

[501] P. Thomas, D. Breckenridge-Miller, and C. Detweiler. Binding characteristics and regulation of the 17,20β, 21-trihydroxy-4-pregnen-3-one (20β-s) receptor on testicular and sperm plasma membranes of spotted seatrout (Cynoscion nebulosus). Fish Physiology and Biochemistry, 17(1-6):109–116, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007781128677.

Kime:1997:SDP

[502] D. E. Kime and M. Ebrahimi. Synthesis of 17,20α- and 17,20β-dihydroxy-4-pregnen-3-ones, 11-ketotestosterone and their conjugates by gills of teleost fish. Fish Physiology and Biochemistry, 17(1-6):117–121, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007777010930.

Pakdel:1997:RGE

[503] F. Pakdel, F. Delaunay, and Y. Valotaire. Regulation of gene expression and biological activity of rainbow trout estrogen receptor. Fish Physiology and Biochemistry, 17(1-6):123–133, December 1997. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007706207857.

Murata:1997:IEG

[504] K. Murata, K. Yamamoto, and K. Yamagami. Intrahepatic expression of genes encoding choriogenins: precursor proteins of the egg envelope of fish, the medaka, oryzias lapites. Fish Physiology and Biochemistry, 17(1-6):135–142, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007702106948.

Ge:1997:ARG

[505] W. Ge, W. Ge, and Y. Nagahama. Activin and its receptors in the goldfish. Fish Physiology and Biochemistry, 17(1-6):143–153, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007718019166.

Cavaco:1997:MBP

[506] J. E. B. Cavaco, H. F. Vischer, and R. W. Schulz. Mismatch between patterns of circulating and testicular androgens in African catfish, Clarias gariepinus. Fish Physiology and Biochemistry, 17(1-6):155–162, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007716514816.

Ohta:1997:AIM

[507] H. Ohta, H. Kagawa, and K. Hirose. Artificial induction of maturation and fertilization in the Japanese eel, Anguilla japonica. Fish Physiology and Biochemistry, 17(1-6):163–169, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007720600588.

Sato:1997:DEP

[508] N. Sato, I. Kawazoe, and K. Aida. Development of an emulsion prepared with lipophilized gelatin and its application for hormone administration in the Japanese eel Anguilla japonica. Fish Physiology and Biochemistry, 17 (1-6):171–178, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007704128929.

Nagae:1997:CTP

[509] M. Nagae, S. Adachi, and K. Yamauchi. Changes in transcription of pituitary glycoprotein hormone α and gonadotropin IIβ subunits during ovarian development induced by repeated injections of salmon pituitary homogenate in the Japanese eel, Anguilla japonica. Fish Physiology and Biochemistry, 17(1-6):179–186, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007708312999.

Yueh:1997:TPO

[510] W. S. Yueh and C. F. Chang. 17α, 20β, 21-trihydroxy-4-pregnen-3-one and 17α, 20β-dihydroxy-4-pregnen-3-one stimulated spermiation in protandrous black porgy, Acanthopagrus schlegeli. Fish Physiology and Biochemistry, 17(1-6):187–193, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007774628330.

Dulka:1997:AIN

[511] J. G. Dulka. Androgen-induced neural plasticity and the regulation of electric-social behavior in the brown ghost knifefish: current status and future directions. Fish Physiology and Biochemistry, 17(1-6):195–202, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007778815126.

Waring:1997:SEC

[512] C. P. Waring and A. Moore. Sublethal effects of a carbamate pesticide on pheromonal mediated endocrine function in mature male Atlantic salmon (Salmo salar L.) parr. Fish Physiology and Biochemistry, 17(1-6):203–211, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007747316943.

Chow:1997:GVI

[513] B. K.-C. Chow. The goldfish vasoactive intestinal polypeptide receptor: functional studies and tissue distribution. Fish Physiology and Biochemistry, 17(1-6):213–222, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007758224695.

Yuen:1997:MCC

[514] T. T. H. Yuen, P. Y. Mok, and B. K. C. Chow. Molecular cloning of a cDNA encoding proglucagon from goldfish, Carassius auratus. Fish Physiology and Biochemistry, 17(1-6):223–230, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007793705131.

Dickhoff:1997:RGE

[515] W. W. Dickhoff, B. R. Beckman, and S. Moriyama. The role of growth in endocrine regulation of salmon smoltification. Fish Physiology and Biochemistry, 17(1-6):231–236, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007710308765.

Farchi-Pisanty:1997:TRF

[516] O. Farchi-Pisanty, H. Sternberg, and B. Moav. Transcriptional regulation of fish growth hormone gene. Fish Physiology and Biochemistry, 17(1-6):237–246, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007762325604.

Funkenstein:1997:OSM

[517] B. Funkenstein, H. Kawauchi, and B. Cavari. Ontogeny of somatolactin mRNA in the gilthead sea bream Sparus aurata. Fish Physiology and Biochemistry, 17(1-6):247–252, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007714409674.

Moriyama:1997:CPI

[518] S. Moriyama, H. Shimma, and H. Kagawa. Changes in plasma insulinlike growth factor-i levels in the precociously maturing amago salmon, Oncorhynchus masou ishikawai. Fish Physiology and Biochemistry, 17(1-6):253–259, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007725010022.

Olivereau:1997:LTS

[519] M. Olivereau and J. M. Olivereau. Long-term starvation in the European eel: general effects and responses of pituitary growth hormone-(GH) and somatolactin-(SL) secreting cells. Fish Physiology and Biochemistry, 17(1-6):261–269, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007766426512.

Zhu:1997:SPS

[520] Y. Zhu and P. Thomas. Studies on the physiology of somatolactin secretion in red drum and Atlantic croaker. Fish Physiology and Biochemistry, 17 (1-6):271–278, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007718510583.

Shepherd:1997:ESD

[521] B. S. Shepherd, B. Ron, and E. G. Grau. Effects of salinity, dietary level of protein and 17α-methyltestosterone on growth hormone (GH) and prolactin (tPRL_{177} and tPRL_{188}) levels in the tilapia, Oreochromis mossambicus. Fish Physiology and Biochemistry, 17(1-6):279–288, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007770511491.

Ogawa:1997:ESP

[522] M. Ogawa, T. Sugai, and T. Watanuki. Effects of salmon prolactin and growth hormone on plasma osmolality, Na^{+} concentration and protein content in the saffron cod. Fish Physiology and Biochemistry, 17(1-6): 289–293, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007797906040.

Xu:1997:OAG

[523] B. Xu, H. Miao, and D. Li. Osmoregulatory actions of growth hormone in juvenile tilapia (Oreochromis niloticus). Fish Physiology and Biochemistry, 17(1-6):295–301, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007750022878.

Madsen:1997:EEN

[524] S. S. Madsen, A. B. Mathiesen, and B. Korsgaard. Effects of 17β-estradiol and 4-nonylphenol on smoltification and vitellogenesis in Atlantic salmon (Salmo salar). Fish Physiology and Biochemistry, 17(1-6):303–312, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007754123787.

Warne:1997:VAN

[525] J. M. Warne and R. J. Balment. Vascular actions of neurohypophysial peptides in the flounder. Fish Physiology and Biochemistry, 17(1-6):313–318, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007745821969.

Loretz:1997:NPI

[526] C. A. Loretz, C. A. Loretz, and Y. Takei. Natriuretic peptide inhibition of intestinal salt absorption in the Japanese eel: physiological significance. Fish Physiology and Biochemistry, 17(1-6):319–324, December 1997. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007772925951.

Acher:1997:MEN

[527] R. Acher, J. Chauvet, and Y. Rouillé. Molecular evolution of neurohypophysial hormones in relation to osmoregulation: the two fish options. Fish Physiology and Biochemistry, 17(1-6):325–332, December 1997. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007720909113.

Tierney:1997:IBR

[528] M. L. Tierney, K. Hamano, and N. Hazon. Interactions between the renin-angiotensin system and catecholamines on the cardiovascular system of elasmobranchs. Fish Physiology and Biochemistry, 17(1-6):333–337, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007795200105.

Brown:1997:AAV

[529] J. A. Brown and S. Amer. Antidiuretic action of vasoactive endothelins in the in situ perfused rainbow trout kidney. Fish Physiology and Biochemistry, 17(1-6):339–345, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007791116034.

Schoenmakers:1997:CCE

[530] Th.J. M. Schoenmakers, G. Flik, and S. E. Wendelaar Bonga. Calcium currents and exocytosis in single isolated pars intermedia cells from tilapia (Oreochromis mossambicus). Fish Physiology and Biochemistry, 17(1-6): 347–355, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007726712400.

Harding:1997:PPA

[531] K. E. Harding, J. M. Warne, and R. J. Balment. Pituitary and plasma AVT content in the flounder (Platichthys flesus). Fish Physiology and Biochemistry, 17(1-6):357–362, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007761801419.

Perry:1997:PPF

[532] S. F. Perry and J. N. Fryer. Proton pumps in the fish gill and kidney. Fish Physiology and Biochemistry, 17(1-6):363–369, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007746217349.

Cutler:1997:EKA

[533] C. P. Cutler, I. L. Sanders, and G. Cramb. Expression of Na^{+}, K^{+}- ATPase β subunit isoforms in the European eel (Anguilla anguilla). Fish Physiology and Biochemistry, 17(1-6):371–376, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007794710505.

Renfro:1997:HRR

[534] J. L. Renfro. Hormonal regulation of renal inorganic phosphate transport in the winter flounder, Pleuronectes americanus. Fish Physiology and Biochemistry, 17(1-6):377–383, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007786424617.

Flik:1997:SDI

[535] G. Flik, T. Kaneko, and J. C. Fenwick. Sodium dependent ion transporters in trout gills. Fish Physiology and Biochemistry, 17(1-6):385–396, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007768825043.

Ura:1997:ISC

[536] K. Ura, S. Mizuno, and K. Yamauchi. Immunohistochemical study on changes in gill Na^{+} /K^{+}-ATPase α-subunit during smoltification in the wild masu salmon, Oncorhynchus masou. Fish Physiology and Biochemistry, 17(1-6):397–403, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007755008204.

Chan:1997:FPC

[537] D. K. O. Chan and C. K. C. Wong. Functional partition of cells in the gill epithelium of the Japanese eel, Anguilla japonica, and the role of hormones. Fish Physiology and Biochemistry, 17(1-6):405–413, December 1997. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007701826203.

Comhaire:1998:BCU

[538] S. Comhaire, R. Blust, and O. L. J. Vanderborght. Branchial cobalt uptake in the carp, Cyprinus carpio: Effect of calcium channelblockers and calcium injection. Fish Physiology and Biochemistry, 18(1):1–13, February 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007746117932.

Devos:1998:ECC

[539] E. Devos, P. Devos, and M. Cornet. Effect of cadmium on the cytoskeleton and morphology of gill chloride cells in parr and smolt Atlantic salmon (Salmo salar). Fish Physiology and Biochemistry, 18(1):15–27, February 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007782027346.

Ohta:1998:COP

[540] H. Ohta and Y. Shinriki. Changes in osmotic pressure that trigger the initiation of spermmotility in the river sculpin cottus hangiongensis. Fish Physiology and Biochemistry, 18(1):29–35, February 1998. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007773429318.

Blazquez:1998:SFE

[541] M. Blázquez, S. Zanuy, and F. Piferrer. Structural and functional effects of early exposure to estradiol-17β and 17α-ethynylestradiol on the gonads of the gonochoristic teleost Dicentrarchus labrax. Fish Physiology and Biochemistry, 18(1):37–47, February 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007736110663.

Wang:1998:DAR

[542] W.-D. Wang, Y.-M. Chen, and C.-H. Hu. Detection of Ah receptor and Ah receptor nuclear translocator mRNAs in the oocytes and developing embryos of zebrafish (Danio rerio). Fish Physiology and Biochemistry, 18(1):49–57, February 1998. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007790214142.

Iijima:1998:PCB

[543] N. Iijima, S. Tanaka, and Y. Ota. Purification and characterization of bile salt-activated lipase from the hepatopancreas of red sea bream, Pagrus major. Fish Physiology and Biochemistry, 18(1):59–69, February 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007725513389.

Hamre:1998:TAG

[544] K. Hamre, R. K. Berge, and ø. Lie. Turnover of α-, γ, and δ tocopherol and distribution insubcellular and lipoprotein fractions indicate presence of an hepatic tocopherolbinding protein in Atlantic salmon (Salmo salar L.). Fish Physiology and Biochemistry, 18(1):71–83, February 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007798102002.

delaHiguera:1998:ITD

[545] M. de la Higuera, A. Garzón, and J. A. Lupiáñez. Influence of temperature and dietary-protein supplementation eitherwith free or coated lysine on the fractional protein-turnover rates in thewhite muscle of carp. Fish Physiology and Biochemistry, 18(1):85–95, February 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007723219594.

Moutou:1998:HPM

[546] K. A. Moutou, M. D. Burke, and D. F. Houlihan. Hepatic P450 monooxygenase response in rainbow trout (Oncorhynchus mykiss (Walbaum)) administered aquaculture antibiotics. Fish Physiology and Biochemistry, 18 (1):97–106, February 1998. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007765227501.

Anonymous:1998:AAB

[547] Anonymous. Application of advanced biochemical and molecular methodologies to seabream aquaculture and its impact on the environment — Eilat, Israel, September 13–17, 1998. Fish Physiology and Biochemistry, 18(1):107, February 1998. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1017150832676.

Matschak:1998:MEA

[548] T. W. Matschak, D. D. Tyler, and N. C. Stickland. Metabolic enzyme activities in Atlantic salmon (Salmo salar L.) embryos respond more to chronic changes in oxygen availability than to environmental temperature. Fish Physiology and Biochemistry, 18(2):115–123, March 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007706108365.

Driedzic:1998:EAL

[549] W. R. Driedzic, J. L. West, and J. A. Raymond. Enzyme activity levels associated with the production of glycerol as an antifreeze in liver of rainbow smelt (Osmerus mordax). Fish Physiology and Biochemistry, 18(2):125–134, March 1998. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007777501093.

Ono:1998:PCP

[550] H. Ono and N. Iijima. Purification and characterization of phospholipase A_{2} isoforms from the hepatopancreas of red sea bream, Pagrus major. Fish Physiology and Biochemistry, 18(2):135–147, March 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007750618685.

Mourente:1998:VIM

[551] G. Mourente and D. R. Tocher. The in vivo incorporation and metabolism of [1-^{14}C] linolenate (18:3n − 3) in liver, brain and eyes of juveniles of rainbow trout Oncorhynchus mykiss L. and gilthead sea bream Sparus aurata L. Fish Physiology and Biochemistry, 18(2):149–165, March 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007717312480.

Koppang:1998:EIL

[552] E. O. Koppang, G. A. Thomas, and C. McL. Press. Expression of insulinlike growth factor-i in the gastrointestinal tract of Atlantic salmon (Salmo salar L.). Fish Physiology and Biochemistry, 18(2):167–175, March 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007771102755.

Rodriguez:1998:IHU

[553] C. Rodríguez, J. R. Cejas, and A. Lorenzo. Influence of n − 3 highly unsaturated fatty acid deficiency on the lipid composition of broodstock gilthead seabream (Sparus aurata L.) and on egg quality. Fish Physiology and Biochemistry, 18(2):177–187, March 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007750218840.

Terjesen:1998:ADR

[554] B. F. Terjesen, A. Mangor-Jensen, and H. J. Fyhn. Ammonia dynamics in relation to hatching in Atlantic halibut (Hippoglossus hippoglossus L.). Fish Physiology and Biochemistry, 18(2):189–201, March 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007748424295.

Noble:1998:EDB

[555] E. Noble, H. Barré, and P. J. Dierickx. Effect of diet and βnaphthoflavone on hepatic and renal glutathione S-transferase isoenzymes in carp (Cyprinus carpio). Fish Physiology and Biochemistry, 18(2):203– 212, March 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007758129746.

Valente:1998:BGC

[556] L. M. P. Valente, E. F. S. Gomes, and B. Fauconneau. Biochemical growth characterization of fast and slow-growing rainbow trout strains: effect of cell proliferation and size. Fish Physiology and Biochemistry, 18(3):213– 224, May 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007774929535.

Jobling:1998:LDA

[557] M. Jobling, S. J. S. Johansen, and E. H. Jørgensen. Lipid dynamics in anadromous Arctic charr, Salvelinus alpinus (L.): seasonal variations in lipid storage depots and lipid class composition. Fish Physiology and Biochemistry, 18(3):225–240, May 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007747201521.

Ghioni:1998:EDF

[558] C. Ghioni, D. R. Tocher, and J. R. Sargent. Effects of dichlorvos and formalin on fatty acid metabolism of rainbow trout (Oncorhynchus mykiss) skin cells in primary culture. Fish Physiology and Biochemistry, 18(3): 241–252, May 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007730730655.

Mol:1998:CSI

[559] K. A. Mol, S. Van der Geyten, and V. M. Darras. Comparative study of iodothyronine outer ring and inner ring deiodinase activities in five teleostean fishes. Fish Physiology and Biochemistry, 18(3):253– 266, May 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007722812697.

Haddy:1998:DVB

[560] J. A. Haddy and N. W. Pankhurst. The dynamics of in vitro 17 β-estradiol secretion by isolated ovarian follicles of the rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 18(3):267–275, May 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007734831563.

Miyazaki:1998:DCD

[561] H. Miyazaki, T. Kaneko, and T. Hirano. Developmental changes in drinking rate and ion and water permeability during early life stages of euryhaline tilapia, Oreochromis mossambicus, reared in fresh water and seawater. Fish Physiology and Biochemistry, 18(3):277–284, May 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007743116542.

Bomgren:1998:SDO

[562] P. Bomgren, S. Einarsson, and A.-C. Jönsson. Similarities and differences in oxynticopeptic cell ultrastructure of one marine teleost, Gadus morhua and one freshwater teleost, Oncorhynchus mykiss, during basal and histamine-stimulated phases of acid secretion. Fish Physiology and Biochemistry, 18(3):285–296, May 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007791032472.

Leguen:1998:PRU

[563] I. Leguen, M. Pisam, and P. Poujeol. pHi regulation and ultrastructural analysis in cultured gill cells from freshwater or seawater-adapted trout. Fish Physiology and Biochemistry, 18(3):297–309, May 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007751319268.

Soengas:1998:DDE

[564] J. L. Soengas, E. F. Strong, and M. Aldegunde. Dose-dependent effects of acute melatonin treatments on brain carbohydrate metabolism of rainbow trout. Fish Physiology and Biochemistry, 18(4):311–319, June 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007753814269.

Stouthart:1998:ESR

[565] X. J. H. X. Stouthart, M. A. J. Huijbregts, and S. E. Wendelaar Bonga. Endocrine stress response and abnormal development in carp (Cyprinus carpio) larvae after exposure of the embryos to PCB 126. Fish Physiology and Biochemistry, 18(4):321–329, June 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1016076415178.

Brokken:1998:IPC

[566] L. J. S. Brokken, P. M. Verbost, and S. E. Wendelaar Bonga. Isolation, partial characterization and localization of integumental peroxidase, a stress-related enzyme in the skin of a teleostean fish (Cyprinus carpio L.). Fish Physiology and Biochemistry, 18(4):331–342, June 1998. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007707520177.

Dugan:1998:CDA

[567] S. G. Dugan and T. W. Moon. Cortisol does not affect hepatic α- and β-adrenoceptor properties in rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 18(4):343–352, June 1998. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007710016995.

Lescroart:1998:AIG

[568] O. Lescroart, I. Roelants, and F. Ollevier. Apomorphine-induced growth hormone-release in African catfish (Clarias gariepinus) is related to the animal’s nutritional status. Fish Physiology and Biochemistry, 18(4):353– 361, June 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007795100613.

Hwang:1998:EMP

[569] P. P. Hwang, M. J. Fang, and S. T. Chen. Expression of mRNA and protein of Na^{+}-K^{+}-ATPase α subunit in gills of tilapia (Oreochromis mossambicus). Fish Physiology and Biochemistry, 18(4):363–373, June 1998. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007711606064.

McDonald:1998:RUK

[570] M. D. McDonald and C. M. Wood. Reabsorption of urea by the kidney of the freshwater rainbow trout. Fish Physiology and Biochemistry, 18(4):375–386, June 1998. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007722318812.

Raymond:1998:ECS

[571] J. A. Raymond and A. L. DeVries. Elevated concentrations and synthetic pathways of trimethylamine oxide and urea in some teleost fishes of Mc-Murdo Sound, Antarctica. Fish Physiology and Biochemistry, 18(4):387– 398, June 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007778728627.

Montero:1998:DSA

[572] D. Montero, L. Tort, and J. M. Vergara. Depletion of serum alternative complement pathway activity in gilthead seabream caused by α-tocopherol and n-3 HUFA dietary deficiencies. Fish Physiology and Biochemistry, 18(4):399–407, June 1998. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007734720630.

Matsuyama:1998:CLV

[573] M. Matsuyama, K. Ohta, and A. Kambegawa. Circulating levels and in vitro production of two maturation-inducing hormones in teleost: 17α, 20β-dihydroxy-4-pregnen-3-one and 17α, 20β, 21-trihydroxy-4-pregnen-3-one, in a daily spawning wrasse, pseudolabrus japonicus. Fish Physiology and Biochemistry, 19(1):1–11, July 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007761729290.

Negatu:1998:EIL

[574] Z. Negatu, S. M. Hsiao, and R. A. Wallace. Effects of insulin-like growth factor-i on final oocyte maturation and steroid production in Fundulus heteroclitus. Fish Physiology and Biochemistry, 19(1):13–21, July 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007777926238.

Wong:1998:DAS

[575] A. O. L. Wong, C. K. Murphy, and R. E. Peter. Direct actions of serotonin on gonadotropin-II and growth hormone release from goldfish pituitary cells: interactions with gonadotropin-releasing hormone and dopamine and further evaluation of serotonin receptor specificity. Fish Physiology and Biochemistry, 19(1):23–34, July 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007713622603.

Payan:1998:RBO

[576] P. Payan, G. Borelli, and N. Mayer-Gostan. Relationship between otolith and somatic growth:consequence of starvation on acid-base balance inplasma and endolymph in the rainbow trout Oncorhynchus mykiss. Fish Physiology and Biochemistry, 19(1):35–41, July 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1023/A:1016064813517.

Conceicao:1998:AAP

[577] L. E. C. Conceição, R. O. A. Ozório, and J. A. J. Verreth. Amino acid profiles and amino acid utilization in larval African catfish (Clarias gariepinus): effects of ontogeny and temperature. Fish Physiology and Biochemistry, 19(1):43–58, July 1998. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007738228055.

Peng:1998:MPS

[578] K. W. Peng, S. F. Chew, and Y. K. Ip. The mudskippers Periophthalmodon schlosseri and Boleophthalmus boddaerti can tolerate environmental NH_{3} concentrations of 446 and 36μM, respectively. Fish Physiology and Biochemistry, 19(1):59–69, July 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007745003948.

Smet:1998:AAP

[579] H. De Smet, R. Blust, and L. Moens. Absence of albumin in the plasma of the common carp Cyprinus carpio: binding of fatty acids to high density lipoprotein. Fish Physiology and Biochemistry, 19(1):71–81, July 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007734127146.

Vielma:1998:CPH

[580] J. Vielma and S. P. Lall. Control of phosphorus homeostasis of Atlantic salmon (Salmo salar) in fresh water. Fish Physiology and Biochemistry, 19(1):83–93, July 1998. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007757321695.

Tsai:1998:WGA

[581] J. C. Tsai and P. P. Hwang. The wheat germ agglutinin binding sites and development of the mitochondria-rich cells in gills of tilapia (Oreochromis mossambicus). Fish Physiology and Biochemistry, 19(1):95– 102, July 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007766531264.

Forster:1998:ESA

[582] M. E. Forster, A. H. Forster, and W. Davison. Effects of serotonin, adrenaline and other vasoactive drugs on the branchial blood vessels of the Antarctic fish Pagothenia borchgrevinki. Fish Physiology and Biochemistry, 19(2):103–109, September 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007739015634.

McKenzie:1998:DFA

[583] D. J. McKenzie, D. A. Higgs, and D. J. Randall. Dietary fatty acid composition influences swimming performance in Atlantic salmon (Salmo salar) in seawater. Fish Physiology and Biochemistry, 19(2):111–122, September 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007779619087.

Dosanjh:1998:IDB

[584] B. S. Dosanjh, D. A. Higgs, and G. Deacon. Influence of dietary blends of menhaden oil and canola oil on growth, muscle lipid composition, and thyroidal status of Atlantic salmon (Salmo salar) in sea water. Fish Physiology and Biochemistry, 19(2):123–134, September 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007727618179.

derGeyten:1998:CPD

[585] S. Van der Geyten, K. A. Mol, and V. M. Darras. Changes in plasma T_{3} during fasting/refeeding in tilapia (Oreochromis niloticus) are mainly regulated through changes in hepatic type II iodothyronine deiodinase. Fish Physiology and Biochemistry, 19(2):135–143, September 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007790527748.

Peres:1998:DRA

[586] A. Péres, J. L. Zambonino Infante, and C. Cahu. Dietary regulation of activities and mRNA levels of trypsin and amylase in sea bass (Dicentrarchus labrax) larvae. Fish Physiology and Biochemistry, 19(2):145–152, September 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007775501340.

Izquierdo:1998:DLP

[587] M. S. Izquierdo and R. J. Henderson. The determination of lipase and phospholipase activities in gut contents of turbot (Scophthalmus maximus) by fluorescence-based assays. Fish Physiology and Biochemistry, 19(2): 153–162, September 1998. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007734425931.

Blanco:1998:AHD

[588] G. Blanco, P. Presa, and J. A. Sánchez. Allozyme heterozygosity and development in Atlantic salmon, Salmo salar. Fish Physiology and Biochemistry, 19(2):163–169, September 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007771417270.

Wong:1998:CNO

[589] H. Y. Wong, L. Y. Fung, and S. C. L. Lo. Constitutive nitric oxide synthase (NOS) activities in big-head carp (Aristichthys nobilis). Fish Physiology and Biochemistry, 19(2):171–179, September 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007755015453.

Plaut:1998:CST

[590] I. Plaut. Comparison of salinity tolerance and osmoregulation in two closely related species of blennies from different habitats. Fish Physiology and Biochemistry, 19(2):181–188, September 1998. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007798712727.

Parsons:1998:PBR

[591] G. R. Parsons and J. K. Carlson. Physiological and behavioral responses to hypoxia in the bonnethead shark, Sphyrna tiburo: routine swimming and respiratory regulation. Fish Physiology and Biochemistry, 19(2):189–196, September 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007730308184.

Aranishi:1998:ERJ

[592] F. Aranishi, N. Mano, and H. Hirose. Epidermal response of the Japanese eel to environmental stress. Fish Physiology and Biochemistry, 19(3):197– 203, October 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007746514851.

Aranishi:1998:FLE

[593] F. Aranishi, N. Mano, and H. Hirose. Fluorescence localization of epidermal cathepsins L and B in the Japanese eel. Fish Physiology and Biochemistry, 19(3):205–209, October 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007779600183.

Kozlova:1998:LCC

[594] T. A. Kozlova. Lipid class composition of benthic-pelagic fishes (Cottocomephorus, Cottoidei) from Lake Baikal. Fish Physiology and Biochemistry, 19(3):211–216, October 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007706831689.

Geurden:1998:RBD

[595] I. Geurden, P. Bergot, and P. Sorgeloos. Relationship between dietary phospholipid classes and neutral lipid absorption in newly-weaned turbot, shape Scophthalmus maximus. Fish Physiology and Biochemistry, 19 (3):217–228, October 1998. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007723515204.

Carpene:1998:BDL

[596] E. Carpene, B. Martin, and L. Dalla Libera. Biochemical differences in lateral muscle of wild and farmed gilthead sea bream (series Sparus aurata L.). Fish Physiology and Biochemistry, 19(3):229–238, October 1998.

CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007742328964.

Kieffer:1998:EFD

[597] J. D. Kieffer and B. L. Tufts. Effects of food deprivation on white muscle energy reserves in rainbow trout (Oncorhynchus mykiss): the relationships with body size and temperature. Fish Physiology and Biochemistry, 19 (3):239–245, October 1998. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007759407275.

Rungruangsak-Torrissen:1998:EVR

[598] K. Rungruangsak-Torrissen, G. M. Pringle, and D. F. Houlihan. Effects of varying rearing temperatures on expression of different trypsin isozymes, feed conversion efficiency and growth in Atlantic salmon (shape Salmo salar L.). Fish Physiology and Biochemistry, 19(3):247–255, October 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007731717021.

Alarcon:1998:CFP

[599] F. J. Alarcón, M. Díaz, and E. Abellń. Characterization and functional properties of digestive proteases in two sparids; gilthead seabream (Sparus aurata) and common dentex (Dentex dentex). Fish Physiology and Biochemistry, 19(3):257–267, October 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007717708491.

Vizziano:1998:EGT

[600] D. Vizziano and F. Le Gac. Effect of gonadotropin type II and 17hydroxy-4-pregnene-3,20-dione on 17,20β-dihydroxy-4-pregnen-3-one production by rainbow trout testes at different developmental stages. Fish Physiology and Biochemistry, 19(3):269–277, October 1998. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007784701222.

Grumiaux:1998:ISC

[601] F. Grumiaux, P. Bulet, and N. Dhainaut-Courtois. Isolation and structural characterization of hepatic metallothionein from the roach (shape Rutilus rutilus L.). Fish Physiology and Biochemistry, 19(3):279–286, October 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007783701091.

Sandbacka:1998:RTG

[602] M. Sandbacka, H. Lilius, and B. Isomaa. Rainbow trout gill epithelial cells in primary culture communicate through gap junctions as demonstrated by dye-coupling. Fish Physiology and Biochemistry, 19(3):287– 292, October 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007732715334.

Nunez:1998:MBE

[603] S. Nunez and J. M. Trant. Molecular biology and enzymology of elasmobranch 3β-hydroxysteroid dehydrogenase. Fish Physiology and Biochemistry, 19(4):293–304, December 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007730619721.

Ebbesson:1998:PIH

[604] L. O. E. Ebbesson, B. Th. Björnsson, and P. Ekström. Propylthiouracilinduced hypothyroidism in Coho salmon, Oncorhynchus kisutch: effects on plasma total thyroxine, total triiodothyronine, free thyroxine, and growth hormone. Fish Physiology and Biochemistry, 19(4):305–314, December 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007775516113.

Pankhurst:1998:FEE

[605] N. W. Pankhurst. Further evidence of the equivocal effects of cortisol on in vitro steroidogenesis by ovarian follicles of rainbow trout Oncorhynchus mykiss. Fish Physiology and Biochemistry, 19(4):315–323, December 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007751403189.

Lokman:1998:GHPa

[606] P. Mark Lokman, Gerard J. Vermeulen, and Graham Young. Gonad histology and plasma steroid profiles in wild New Zealand freshwater eels (Anguilla dieffenbachii and A. australis) before and at the onset of the natural spawning migration. I. Females. Fish Physiology and Biochemistry, 19 (4):325–338, December 1998. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007719414295.

Lokman:1998:GHPb

[607] P. M. Lokman and G. Young. Gonad histology and plasma steroid profiles in wild New Zealand freshwater eels (Anguilla dieffenbachii and A. australis) before and at the onset of the natural spawning migration. II. Males. Fish Physiology and Biochemistry, 19(4):339–347, December 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007767330225.

Clearwater:1998:GRH

[608] S. J. Clearwater and L. W. Crim. Gonadotropin releasing hormoneanalogue treatment increases sperm motility, seminal plasma pH and sperm production in yellowtail flounder Pleuronectes ferrugineus. Fish Physiology and Biochemistry, 19(4):349–357, December 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007759620936.

Senthilkumaran:1998:MII

[609] B. Senthilkumaran and K. P. Joy. Methallibure-induced inhibition of hypothalamo-hypophyseal-ovarian activity in the catfish Heteropneustes fossilis involves changes in hypothalamic monoamine activity. Fish Physiology and Biochemistry, 19(4):359–364, December 1998. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007707604097.

Geurden:1998:IFA

[610] I. Geurden, O. S. Reyes, and P. Sorgeloos. Incorporation of fatty acids from dietary neutral lipid in eye, brain and muscle of postlarval turbot fed diets with different types of phosphatidylcholine. Fish Physiology and Biochemistry, 19(4):365–375, December 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007771431134.

Morgan:1998:EEW

[611] I. J. Morgan, L. M. D-Cruz, and C. M. Wood. The effects of elevated winter temperature and sub-lethal pollutants (low pH, elevated ammonia) on protein turnover in the gill and liver of rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 19(4):377–389, December 1998. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007748213387.

Roy:1999:FAC

[612] R. Roy, E. Fodor, and T. Farkas. Fatty acid composition of the ingested food only slightly affects physicochemical properties of liver total phospholipids and plasma membranes in cold-adapted freshwater fish. Fish Physiology and Biochemistry, 20(1):1–11, January 1999. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007755520027.

Olsen:1999:ITD

[613] R. E. Olsen, E. Løvaas, and Ø. Lie. The influence of temperature, dietary polyunsaturated fatty acids, α-tocopherol and spermine on fatty acid composition and indices of oxidative stress in juvenile Arctic char, Salvelinus alpinus (L.). Fish Physiology and Biochemistry, 20(1):13–29, January 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007767827996.

Lay:1999:WDS

[614] P. A. Lay and J. Baldwin. What determines the size of teleost erythrocytes? Correlations with oxygen transport and nuclear volume. Fish Physiology and Biochemistry, 20(1):31–35, January 1999. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007785202280.

Aranishi:1999:LPB

[615] F. Aranishi. Lysis of pathogenic bacteria by epidermal cathepsins L and B in the Japanese eel. Fish Physiology and Biochemistry, 20(1):37–41, January 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007763711158.

Ruane:1999:MRR

[616] N. M. Ruane, D. T. Nolan, and S. E. Wendelaar Bonga. Modulation of the response of rainbow trout (Oncorhynchus mykiss Walbaum) to confinement, by an ectoparasitic (Argulus foliaceus L.) infestation and cortisol feeding. Fish Physiology and Biochemistry, 20(1):43–51, January 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007744617518.

Montero:1999:HSD

[617] D. Montero, M. S. Izquierdo, and J. M. Vergara. High stocking density produces crowding stress altering some physiological and biochemical parameters in gilthead seabream, Sparus aurata, juveniles. Fish Physiology and Biochemistry, 20(1):53–60, January 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007719928905.

McKendry:1999:NPC

[618] J. E. McKendry, N. J. Bernier, and S. F. Perry. Natriuretic peptides and the control of catecholamine release in two freshwater teleost and a marine elasmobranch fish. Fish Physiology and Biochemistry, 20(1):61–77, January 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007792500680.

Jakobsson:1999:KIK

[619] S. Jakobsson, B. Borg, and S. J. Hyllner. An 11-ketotestosterone induced kidney-secreted protein: the nest building glue from male three-spined stickleback, Gasterosteus aculeatus. Fish Physiology and Biochemistry, 20(1):79–85, January 1999. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007776016610.

Owen:1999:EVR

[620] S. F. Owen, I. D. McCarthy, and M. J. Rennie. In vivo rates of protein synthesis in Atlantic salmon (Salmo salar L.) smolts determined using a stable isotope flooding dose technique. Fish Physiology and Biochemistry, 20(1):87–94, January 1999. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007724012975.

Puente-Novoa:1999:PKC

[621] J. M. Puente-Novoa and P. Barja. Protein kinase C in the spleen of the turbot (Scophthalmus maximus L.). Fish Physiology and Biochemistry, 20 (2):101–114, February 1999. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007759610249.

Kraak:1999:PFA

[622] G. Van Der Kraak and S. Biddiscombe. Polyunsaturated fatty acids modulate the properties of the sex steroid binding protein in goldfish. Fish Physiology and Biochemistry, 20(2):115–123, February 1999. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007707609341.

Perez:1999:UER

[623] J. A. Pérez, C. Rodríguez, and R. J. Henderson. The uptake and esterification of radiolabelled fatty acids by enterocytes isolated from rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 20 (2):125–134, February 1999. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007795516689.

Samuelsen:1999:OBP

[624] E. N. Samuelsen, A. K. Imsland, and O. Brix. Oxygen binding properties of three different hemoglobin genotypes in turbot (Scophthalmus maximus Rafinesque): Effect of temperature and pH. Fish Physiology and Biochemistry, 20(2):135–141, February 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007755525270.

Bon:1999:EAP

[625] E. Bon, B. Breton, and F. Le Menn. Effects of accelerated photoperiod regimes on the reproductive cycle of the female rainbow trout: II seasonal variations of plasma gonadotropins (GTH I and GTH II) levels correlated with ovarian follicle growth and egg size. Fish Physiology and Biochemistry, 20(2):143–154, February 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007783708432.

Hou:1999:ESH

[626] Y. Y. Hou, Y. Suzuki, and K. Aida. Effects of steroid hormones on immunoglobulin M (IgM) in rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 20(2):155–162, February 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007799617597.

Leonard:1999:EMD

[627] J. B. K. Leonard and S. D. McCormick. The effect of migration distance and timing on metabolic enzyme activity in an anadromous clupeid, the American shad (Alosa sapidissima). Fish Physiology and Biochemistry, 20(2):163–179, February 1999. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007751701668.

Martinez:1999:TCT

[628] M. Martínez, P. Couture, and H. Guderley. Temporal changes in tissue metabolic capacities of wild Atlantic cod Gadus morhua (L.), from Newfoundland. Fish Physiology and Biochemistry, 20(2):181–191, February 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007747600759.

Pradet-Balade:1999:THR

[629] B. Pradet-Balade, C. Burel, and G. Boeuf. Thyroid hormones downregulate thyrotropin β mRNA level in vivo in the turbot (Psetta maxima). Fish Physiology and Biochemistry, 20(3):193–199, March 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007791415780.

Mol:1999:EEH

[630] K. A. Mol, S. Van der Geyten, and V. M. Darras. Effects of experimental hypo- and hyperthyroidism on iodothyronine deiodinases in Nile tilapia, Oreochromis niloticus. Fish Physiology and Biochemistry, 20 (3):201–207, March 1999. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007739431710.

Raine:1999:OTT

[631] J. C. Raine and J. F. Leatherland. Ontogeny of thyroid tissue and tissue thyroid hormone clearance in rainbow trout embryos reared at two temperatures. Fish Physiology and Biochemistry, 20(3):209–217, March 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007775807438.

Perrot:1999:CDI

[632] V. Perrot and B. Funkenstein. Cellular distribution of insulin-like growth factor II (IGF-II) mRNA and hormonal regulation of IGF-i and IGF-II mRNA expression in rainbow trout testis (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 20(3):219–229, March 1999. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007735314871.

Afonso:1999:EAI

[633] L. O. B. Afonso, G. K. Iwama, and E. M. Donaldson. Effects of the aromatase inhibitor Fadrozole on plasma sex steroid secretion and oocyte maturation in female Coho salmon (Oncorhynchus kisutch) during vitellogenesis. Fish Physiology and Biochemistry, 20(3):231–241, March 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007771723367.

Srivastava:1999:AEA

[634] A. S. Srivastava, I. Oohara, and S. N. Singh. Activity and expression of aspartate aminotransferase during the reproductive cycle of a fresh water fish, Clarias batrachus. Fish Physiology and Biochemistry, 20(3):243– 250, March 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007783213963.

Ramseyer:1999:EDZ

[635] L. Ramseyer, D. Garling, Jr., and J. Link. Effect of dietary zinc supplementation and phytase pre-treatment of soybean meal or corn gluten meal on growth, zinc status and zinc-related metabolism in rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 20(3):251–261, March 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007719722459.

Bell:1999:EEF

[636] J. G. Bell, D. R. Tocher, and J. R. Sargent. Effects of essential fatty aciddeficient diets on growth, mortality, tissue histopathology and fatty acid compositions in juvenile turbot (Scophthalmus maximus). Fish Physiology and Biochemistry, 20(3):263–277, March 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007743532618.

Smith:1999:CHS

[637] T. R. Smith, G. C. Tremblay, and T. M. Bradley. Characterization of the heat shock protein response of Atlantic salmon (Salmo salar). Fish Physiology and Biochemistry, 20(3):279–292, March 1999. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007743329892.

Lemieux:1999:DDE

[638] H. Lemieux, P. Blier, and J.-D. Dutil. Do digestive enzymes set a physiological limit on growth rate and food conversion efficiency in the Atlantic cod (Gadus morhua)? Fish Physiology and Biochemistry, 20(4):293– 303, May 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007791019523.

Dersjant-Li:1999:IDC

[639] Y. Dersjant-Li, J. A. J. Verreth, and M. W. A. Verstegen. The influence of dietary cation-anion differences on acid-base balance, food intake, growth and nutrient utilisation of juvenile African catfish Clarias gariepinus (Burchell). Fish Physiology and Biochemistry, 20(4):305–311, May 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007707507228.

Sun:1999:AZD

[640] L.-T. Sun and S.-S. Jeng. Accumulation of zinc from diet and its release in common carp. Fish Physiology and Biochemistry, 20(4):313– 324, May 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007770431819.

Aldegunde:1999:ECE

[641] M. Aldegunde, J. L. Soengas, and M. D. Andrés. Effects of chronic exposure to γ-HCH (Lindane) on brain serotonergic and gabaergic systems, and serum cortisol and thyroxine levels of rainbow trout, Oncorhynchus mykiss. Fish Physiology and Biochemistry, 20(4):325–330, May 1999. CO-DEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007795121340.

Sternberg:1999:RGH

[642] H. Sternberg and B. Moav. Regulation of the growth hormone gene by fish thyroid/retinoid receptors. Fish Physiology and Biochemistry, 20(4):331– 339, May 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007778632727.

Maestro:1999:DII

[643] M. A. Maestro, E. Méndez, and J. Gutiérrez. Dynamics of insulin and insulin-like growth factor-i (IGF-I) ovarian receptors during maturation in the brown trout (Salmo trutta). Fish Physiology and Biochemistry, 20(4):341–349, May 1999. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007767605620.

Martin:1999:ASS

[644] S. A. M. Martin, A. F. Youngson, and A. Ferguson. Atlantic salmon (Salmo salar) prolactin cDNA sequence and its mRNA expression after transfer of fish between salinities. Fish Physiology and Biochemistry, 20(4):351–359, May 1999. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007711504712.

Heppell:1999:GMM

[645] S. A. Heppell and C. V. Sullivan. Gag (mycteroperca microlepis) vitellogenin: purification, characterization and use for enzyme-linked immunosorbent assay (ELISA) of female maturity in three species of grouper. Fish Physiology and Biochemistry, 20(4):361–374, May 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007730816797.

Lahnsteiner:1999:PBP

[646] F. Lahnsteiner, T. Weismann, and R. A. Patzner. Physiological and biochemical parameters for egg quality determination in lake trout, Salmo trutta lacustris. Fish Physiology and Biochemistry, 20(4):375– 388, May 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007715621550.

Bencic:1999:ALC

[647] D. C. Bencic, M. Krisfalusi, and R. L. Ingermann. ATP levels of Chinook salmon (Oncorhynchus tshawytscha) sperm following in vitro exposure to various oxygen tensions. Fish Physiology and Biochemistry, 20(4):389– 397, May 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007749703803.

Sandbacka:1999:CRT

[648] M. Sandbacka, H. Lilius, and B. Isomaa. Corrigendum — rainbow trout gill epithelial cells in primary culture communicate through gap junctions as demonstrated by dye-coupling. Fish Physiology and Biochemistry, 20(4):399, May 1999. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007813029384.

Amiri:1999:VST

[649] B. M. Amiri, M. Maebayashi, and K. Yamauchi. In vitro steroidogenesis by testicular fragments and ovarian follicles in a hybrid sturgeon, bester. Fish Physiology and Biochemistry, 21(1):1–14, July 1999. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007706128184.

Moon:1999:FHG

[650] T. W. Moon, E. R. Busby, and T. P. Mommsen. Fish hepatocyte glycogen phosphorylase — a sensitive indicator for hormonal modulation. Fish Physiology and Biochemistry, 21(1):15–24, July 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1023/A:1007762229093.

Mancera:1999:ICG

[651] J. M. Mancera and S. D. McCormick. Influence of cortisol, growth hormone, insulin-like growth factor I and 3,3′, 5-triiodo-l-thyronine on hypoosmoregulatory ability in the euryhaline teleost Fundulus heteroclitus. Fish Physiology and Biochemistry, 21(1):25–33, July 1999. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007737924339.

Olsen:1999:LDU

[652] R. E. Olsen, R. Myklebust, and E. Ringø. Lipid digestibility and ultrastructural changes in the enterocytes of Arctic char (Salvelinus alpinus L.) fed linseed oil and soybean lecithin. Fish Physiology and Biochemistry, 21(1):35–44, July 1999. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007726615889.

Mourente:1999:ULD

[653] G. Mourente, A. Rodríguez, and E. Pastor. Utilization of lipids by Dentex dentex L. (Osteichthyes, Sparidae) larvae during lecitotrophia and subsequent starvation. Fish Physiology and Biochemistry, 21(1):45– 58, July 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007789908409.

Kestemont:1999:CEL

[654] P. Kestemont, J. Cooremans, and C. Mélard. Cathepsin L in eggs and larvae of perch Perca fluviatilis: variations with developmental stage and spawning period. Fish Physiology and Biochemistry, 21(1):59– 64, July 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007714314072.

Liu:1999:CNT

[655] Z. Liu, P. Li, and R. Dunham. Characterization of nonautonomous tc1-like transposable elements of channel catfish (Ictalurus punctatus). Fish Physiology and Biochemistry, 21(1):65–72, July 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007710212254.

Suzuki:1999:SEA

[656] T. Suzuki, T. Kurokawa, and D. C. McVey. Sequence and expression analyses of cholecystokinin (CCK) precursor cDNA in the Japanese flounder (Paralichthys olivaceus). Fish Physiology and Biochemistry, 21(1):73– 80, July 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007703422249.

Peters:1999:CEO

[657] R. C. Peters and R. H. S. Westerink. Catfish electroreceptor organ functioning during five days exposure to different calcium environments. Fish Physiology and Biochemistry, 21(1):81–88, July 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.

springer.com/article/10.1023/A:1007712627275.

Blin:1999:TLH

[658] C. Blin, S. Panserat, and R. Krishnamoorthy. Teleost liver hexokinaseand glucokinase-like enzymes: partial cDNA cloning and phylogenetic studies in rainbow trout (Oncorhynchus mykiss), common carp (Cyprinus carpio) and gilthead seabream (Sparus aurata)+. Fish Physiology and Biochemistry, 21(2):93–102, August 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007748204428.

Kagawa:1999:EES

[659] N. Kagawa, K. Ryo, and Y. Mugiya. Enhanced expression of stress protein 70 in the brains of goldfish, Carassius auratus, reared with bluegills, Lepomis macrochirus. Fish Physiology and Biochemistry, 21(2):103–110, August 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007785707501.

Yao:1999:BCG

[660] K. Yao and P.-Y. Le Bail. Biochemical characterization of growth hormone receptor in rainbow trout (Oncorhynchus mykiss) before and after purification. Fish Physiology and Biochemistry, 21(2):111–120, August 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007733723430.

Vissio:1999:CAG

[661] P. G. Vissio, A. V. Stefano, and D. A. Paz. Close association of gonadotropin-releasing hormone fibers and gonadotropin, growth hormone, somatolactin and prolactin expressing cells in pejerrey, Odontesthes bonariensis. Fish Physiology and Biochemistry, 21(2):121–127, August 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007781606592.

Reddy:1999:ECT

[662] P. K. Reddy, R. Renaud, and J. F. Leatherland. Effects of cortisol and triiodo-L-thyronine on the steroidogenic capacity of rainbow trout ovarian follicles at two stages of oocyte maturation. Fish Physiology and Biochemistry, 21(2):129–140, August 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007800707094.

Kulczykowska:1999:DCP

[663] E. Kulczykowska. Diel changes in plasma arginine vasotocin, isotocin, and melatonin in rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 21(2):141–146, August 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007808924841.

Pickova:1999:FAC

[664] J. Pickova, A. Kiessling, and P. C. Dutta. Fatty acid and carotenoid composition of eggs from two nonanadromous Atlantic salmon stocks of cultured and wild origin. Fish Physiology and Biochemistry, 21(2):147– 156, August 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007860908911.

Krieger:1999:YMP

[665] J. Krieger and R. Fleig. Yolk mobilization in perch, Perca fluviatilis L., embryos. Fish Physiology and Biochemistry, 21(2):157–165, August 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007843306186.

Eddy:1999:ENO

[666] F. B. Eddy, L. McGovern, and J. C. McGeer. Effect of a nitric oxide releasing compound isosorbide dinitrate on development and cardiovascular physiology of rainbow trout (Oncorhynchus mykiss). Fish Physiology and Biochemistry, 21(2):167–171, August 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007712622522.

Schlenk:1999:EEC

[667] D. Schlenk, E. J. Perkins, and W. B. Hawkins. Effect of ethanol, clofibric acid and temperature on the uptake and elimination of 2-methylisoborneol in channel catfish (Ictalurus punctatus). Fish Physiology and Biochemistry, 21(2):173–178, August 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007706505337.

Saucier:1999:CIO

[668] D. Saucier, A. K. Julliard, and L. Astic. CYP1A1 immunolocalization in the olfactory organ of rainbow trout and its possible induction by βnaphthoflavone: analysis in adults and embryos around hatching. Fish Physiology and Biochemistry, 21(2):179–192, August 1999. CODEN FP-BIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007754421267.

Bencic:1999:MST

[669] D. C. Bencic, M. Krisfalusi, and R. L. Ingermann. Maintenance of steelhead trout (Oncorhynchus mykiss) sperm at different in vitro oxygen tensions alters ATP levels and cell functional characteristics. Fish Physiology and Biochemistry, 21(3):193–200, October 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007880426488.

Yoshiura:1999:MCC

[670] Y. Yoshiura, Y. C. Sohn, and K. Aida. Molecular cloning of the cDNA encoding the β subunit of thyrotropin and regulation of its gene expression by thyroid hormones in the goldfish, Carassius auratus. Fish Physiology and Biochemistry, 21(3):201–210, October 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007884527397.

Reddy:1999:MSI

[671] P. K. Reddy, A. C. Holloway, and J. F. Leatherland. Melatonin, but not somatostatin-14 influences in vitro steroidogenesis by ovarian follicles of rainbow trout. Fish Physiology and Biochemistry, 21(3):211–222, October 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007811907833.

Rungruangsak-Torrissen:1999:MRP

[672] K. Rungruangsak-Torrissen, C. G. Carter, and D. F. Houlihan. Maintenance ration, protein synthesis capacity, plasma insulin and growth of Atlantic salmon (Salmo salar L.) with genetically different trypsin isozymes. Fish Physiology and Biochemistry, 21(3):223–233, October 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007804823932.

Davis:1999:RJR

[673] D. A. Davis, J. P. Lazo, and C. R. Arnold. Response of juvenile red drum (Sciaenops ocellatus) to practical diets supplemented with medium chain triglycerides. Fish Physiology and Biochemistry, 21(3):235–248, October 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007836612376.

Craig:1999:EDL

[674] S. R. Craig, B. S. Washburn, and D. M. Gatlin III. Effects of dietary lipids on body composition and liver function in juvenile red drum, Sciaenops ocellatus. Fish Physiology and Biochemistry, 21(3):249–255, October 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007843420128.

Tocher:1999:PFA

[675] D. R. Tocher and J. R. Dick. Polyunsaturated fatty acid metabolism in a cell culture model of essential fatty acid deficiency in a freshwater fish, carp (Cyprinus carpio). Fish Physiology and Biochemistry, 21 (3):257–267, October 1999. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007832510559.

Nikinmaa:1999:IPR

[676] M. Nikinmaa, A. Bogdanova, and L. V. Virkki. Intracellular pH regulation of rainbow trout and carp thrombocytes. Fish Physiology and Biochemistry, 21(3):269–275, October 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007856808003.

Wartman:1999:EAT

[677] C. A. Wartman, J. M. Mokler, and J. F. Burka. Effects of acclimation temperatures of 7 and 14^{o}C on the contractility of Atlantic salmon (Salmo salar) intestinal smooth muscle in vitro. Fish Physiology and Biochemistry, 21(3):277–284, October 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1007803721945.

Anonymous:1999:PA

[678] Anonymous. Publisher’s announcement. Fish Physiology and Biochemistry, 21(4):285, December 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.com/

article/10.1023/A:1017291417129.

Huriaux:1999:MPW

[679] F. Huriaux, P. Vandewalle, and B. Focant. Myofibrillar proteins in white muscle of the developing African catfish Heterobranchus longifilis (Siluriforms, Clariidae). Fish Physiology and Biochemistry, 21(4):287–301, December 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007835101472.

Saha:1999:SUP

[680] N. Saha and L. Das. Stimulation of ureogenesis in the perfused liver of an Indian air-breathing catfish, Clarias batrachus, infused with different concentrations of ammonium chloride. Fish Physiology and Biochemistry, 21 (4):303–311, December 1999. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007856432423.

Ciereszko:1999:PUA

[681] A. Ciereszko, K. Dabrowski, and J. Glogowski. The presence of uric acid, an antioxidantive substance, in fish seminal plasma. Fish Physiology and Biochemistry, 21(4):313–315, December 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://link.springer.

com/article/10.1023/A:1007886121663.

Martinez:1999:DEA

[682] I. Martínez, F. J. Moyano, and M. Yúfera. Digestive enzyme activity during larval development of the Senegal sole (Solea senegalensis). Fish Physiology and Biochemistry, 21(4):317–323, December 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL https://

link.springer.com/article/10.1023/A:1007802708459.

Aas:1999:BAM

[683] G. H. Aas, B. Bjerkeng, and B. Ruyter. Blood appearance, metabolic transformation and plasma transport proteins of ^{14}C-astaxanthin in Atlantic salmon (Salmo salar L.). Fish Physiology and Biochemistry, 21 (4):325–334, December 1999. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007890224389.

Pavlidis:1999:PTA

[684] M. Pavlidis, M. Berry, and M. Kentouri. Prothrombin time, activated partial thromboplastin time and fibrinogen values in Mediterranean marine teleosts. Fish Physiology and Biochemistry, 21(4):335–343, December 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168

(electronic). URL https://link.springer.com/article/10.1023/A:
1007882020754.

Lee:1999:ETS

[685] Y.-H. Lee, J.-L. Du, and C.-F. Chang. 17β-estradiol, but not testosterone stimulates gonadotropin II concentrations in the protandrous black porgy, Acanthopagrus schlegeli Bleeker. Fish Physiology and Biochemistry, 21 (4):345–351, December 1999. CODEN FPBIEP. ISSN 0920-1742 (print),

1573-5168 (electronic). URL https://link.springer.com/article/10.
1023/A:1007830019846.

Bjerkeng:1999:IKE

[686] B. Bjerkeng, K. Johnsen, and K. J. Nilssen. Influence of 11ketotestosterone, 17β-estradiol, and 3,5,3′-triiodo-L-thyronine on distribution and metabolism of carotenoids in Arctic charr, Salvelinus alpinus L. Fish Physiology and Biochemistry, 21(4):353–364, December 1999. CODEN FPBIEP. ISSN 0920-1742 (print), 1573-5168 (electronic). URL

https://link.springer.com/article/10.1023/A:1007802803008.